### v1

# Segmenting Action-Value Functions Over Time-Scales in SARSA using TD($\Delta$)

Abstract
In numerous episodic reinforcement learning (RL) settings, SARSA-based methodologies are employed to enhance policies aimed at maximizing returns over long horizons. Conventional SARSA algorithms, however, have difficulties in balancing bias and variation due to the reliance on a singular, fixed discount factor. This study expands the temporal difference decomposition approach, TD($\triangle$), to the SARSA algorithm. SARSA, a widely utilised on-policy RL method, enhances action-value functions via temporal difference updates. TD($\triangle$) facilitates learning over several time-scales by breaking the action-value function into components associated with distinct discount factors. This decomposition improves learning efficiency and stability, particularly in problems necessitating long-horizon optimization. We illustrate that our methodology mitigates bias in SARSA’s updates while facilitating accelerated convergence in contexts characterized by dense rewards. Experimental findings across many benchmark tasks indicate that the proposed SARSA($\triangle$) surpasses conventional TD learning methods in both tabular and deep RL contexts.
keywords:
SARSA, Temporal Difference, Action-Value Functions, On-policy, Finite Horizon
1 Introduction
In reinforcement learning (RL), the objective is to enhance an agent’s actions to maximize cumulative rewards through interaction with an environment. Temporal difference (TD) learning methods, like Q-learning and SARSA, have demonstrated efficacy across several tasks by allowing agents to estimate action-value functions that forecast anticipated future rewards. Traditionally, these strategies utilise a discount factor $10\leq\gamma<$ 1, with values approaching $\gamma=0$ to emphasise short-term rewards over long-term ones, therefore diminishing the planning horizon and enhancing learning stability and efficiency[1]. Discount factors $\gamma<1$ frequently yield enhanced results during the early phases of learning, as demonstrated by Prokhorov and Wunsch[2]. In specific contexts, such as long-horizon tasks[3, 4] where long-term planning is essential—like navigation or decision-making tasks with deferred rewards—this method can create bias, constraining the agent’s capacity to acquire optimal long-horizon policies.
For instance: SARSA[1] (State-Action-Reward-State-Action) is an on-policy RL algorithm that modifies an action-value function Q(s,a) according to the state-action transitions encountered by the agent. The conventional SARSA algorithm performs well in contexts with shorter time horizons; however, in long-horizon tasks, the singular discount factor $\gamma$ employed in SARSA results in a compromise between bias and variance. Small values of $\gamma$ yield myopic policies, whereas larger values induce substantial variance in the estimated action-value functions, rendering the learning process unstable and inefficient. In response, other solutions have been suggested, including[4, 2, 5, 6, 7, 8].
Recent research in TD learning has investigated the segmentation of value functions over many time-scales, enabling agents to concurrently enhance value estimates at numerous horizons[8]. The TD($\Delta$) architecture enhances scalability and performance in situations with long-term dependencies by decomposing value functions into components that encompass various discount factors, referred to as TD($\Delta$). The TD($\Delta$) framework’s adaptability demonstrates significant potential in mitigating the bias-variance trade-offs inherent in TD learning.
We suggest SARSA($\Delta$), an extension of the TD($\Delta$) method applied to the SARSA algorithm, building on this concept. TD($\Delta$) emphasises the actor-critic paradigm and conventional TD learning, whereas SARSA($\Delta$) disaggregates the state-action value function into several partial estimators across various discount factors, termed delta estimators. These estimators approximate the difference $W_{z}(s,a)=Q_{\gamma_{z}}(s,a)\mathit{-}Q_{\gamma_{z-1}}(s,a)$ between action-value functions, facilitating efficient learning in scenarios where actions affect long-horizon outcomes. SARSA($\Delta$) optimises a sequence of delta estimators, each associated with a distinct discount factor, similar to TD($\Delta$). The segregation enables the autonomous learning of each component, with lower discount factors achieving faster convergence, while bigger discount factors enhance this foundation for improved long-horizon planning.
This study introduces a multi-step SARSA($\Delta$) update rule that disaggregates state-action value functions, facilitating accelerated convergence and enhanced policy efficacy, especially in intricate environments characterised by dense rewards. Furthermore, we illustrate that SARSA($\Delta$) may seamlessly adapt to any multi-step reinforcement learning techniques, including n-step SARSA and eligibility traces. Through theoretical analysis and empirical findings, we emphasise the benefits of SARSA($\Delta$) in diverse tasks, including a basic ring MDP utilised by Kearns & Singh[9], demonstrating its superiority over conventional SARSA in scenarios necessitating a balance between short-term action consequences and long-term reward maximisation. SARSA($\Delta$) possesses the capability to adapt discount factors across varying time scales, serving as an effective instrument in environments where time scale is critical.
The impetus for extending TD($\Delta$) to SARSA($\Delta$) lies in the fact that temporal difference (TD($\Delta$)) learning offers a framework for deconstructing value functions across many time scales, hence overcoming the constraints of single-scale learning in long-horizon problems. By extending TD($\Delta$) to SARSA($\Delta$), we seek to decompose the action-value function Q(s,a) into elements that emphasise distinct discount factors. This decomposition enables SARSA to utilise the advantages of learning across several time scales, hence stabilising the learning process and mitigating the bias-variance tradeoff. This addon provides a more detailed and scalable method for action-value function learning in intricate environments.
The primary contributions of this paper are as follows: (i) We expand TD($\Delta$) to SARSA($\Delta$) by partitioning the action-value function Q(s,a) into delta components over several time scales. (ii) The suggested SARSA($\Delta$) algorithm enhances learning efficiency and stability by independently learning components for various discount factors and integrating them to represent the complete action-value function. (iii) Both theoretical and empirical evaluations have shown that this multi-scale breakdown has a number of benefits, particularly in environments where there are long-horizon rewards.
The subsequent sections of the paper are structured as follows: Section 2 contains a discussion of related works. Section 3 presents the necessary background information. Sections 4 and 5 illustrate the principles of SARSA($\Delta$) and theoretical analysis, respectively. Section 6 presents a comprehensive overview of the experiments conducted. Ultimately, we provide a conclusion in section 7.
2 Related Work
In reinforcement learning (RL), optimising for long-horizon rewards is tough due to the complexity of learning with an undiscounted return. Temporal discounting is frequently employed to streamline this process; nevertheless, it may induce bias. Researchers[8] have investigated techniques to separate value functions across various time scales by analysing disparities among value functions with lower discount factors, hence enhancing scalability and performance. The study introduces $\gamma$-nets[10], a technique for generalising value function estimate across timescales in reinforcement learning. The authors of the paper[11] present an approach for integrating hyperbolic discounting into RL. Traditional RL employs exponential discounting, which does not align with the hyperbolic discounting observed in human and animal behaviors. The authors present an agent designed to approximate hyperbolic discount functions by utilizing temporal-difference learning methods. Furthermore, the study reveals that learning value functions across multiple time horizons improves performance, especially when implemented as an auxiliary task in value-based reinforcement learning algorithms such as Rainbow. This method demonstrates potential in contexts characterized by uncertainty and risk. Expected SARSA[7], a variant of SARSA, utilizes information regarding the stochastic nature of the behavior policy to execute updates with reduced variance, which speeds up learning. This study[12] presents temporal-difference (TD) learning methods, which underpin numerous RL algorithms, notably SARSA, by forecasting future outcomes based on the discrepancies between consecutive forecasts rather than the ultimate result. The research demonstrates the convergence of TD techniques under specific situations and emphasises their benefits in predictive tasks, particularly when incremental learning is crucial. This study[13] presents theoretical convergence results for temporal-difference learning methods utilising function approximation. Extending this analysis to SARSA($\Delta$) enables us to assert comparable convergence properties, provided that the delta components are updated in accordance with the SARSA update rule. It is noteworthy that none of the previously mentioned works employ short-term estimates for training long-term action-value functions. Our method can also be utilised as a generalised action-value function, and its capacity to query smaller timescales represents an additional advantage. Performance improvements primarily arise from the separation of action-value functions across varying time scales through the use of TD($\Delta$) in SARSA.
Recently, there has been some research conducted on the precise selection of the discount factor[2, 5, 6, 4]. The authors[6] of this study introduce meta-gradient methods for the adaptive selection of discount factors in RL, emphasising the necessity of balancing short- and long-term rewards, which corresponds with the rationale for dissecting action-value functions across several discount factors in SARSA($\Delta$).
Lastly, another extensive category of study pertinent to our research indirectly encompasses hierarchical RL and the bias-variance tradeoff. These studies[14, 15, 16, 17, 18, 19, 20] present the concept of deconstructing value functions into subtasks by hierarchical RL, analogous to the manner in which TD($\Delta$) disaggregates value functions across temporal scales. This substantiates the assertion that value function decomposition can enhance learning efficacy and stability. This paper[9] functions as a theoretical foundation for comprehending the bias-variance tradeoff that arises when decomposing Q-values across various time scales, and it also establishes constraints on the bias and variance of TD updates. This study[21] presents empirical findings demonstrating the efficacy of deep RL techniques utilizing value function approximation in Atari-like environments. It also endorses the necessity for techniques that diminish variance in long-horizon tasks.
3 Background and notation
Examine a completely observable Markov Decision Process (MDP) [22], characterized by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ signifies the action space, and $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}\rightarrow[0,1]$ represents the transition probabilities that associate state-action pairings with distributions across subsequent states, whereas $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denotes the reward function. At each timestep $t$, the agent occupies state $s_{t}$, chooses an action $a_{t}$, obtains a reward $r_{t}=r(s_{t},a_{t})$, and transitions to the subsequent state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$.
In a conventional Markov Decision Process (MDP) framework, an agent seeks to maximize the discounted return defined as $Q_{\gamma}^{\pi}(s,a)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t+1}\mid s_{t}=s,a_{t}=a\right]$, where $\gamma$ denotes the discount factor and $\pi:\mathcal{S}\rightarrow\mathcal{A}\rightarrow[0,1]$ represents the policy adhered to by the agent. The action-value function $Q_{\gamma}^{\pi}(s,a)$ is determined as the fixed point of the Bellman operator $\mathcal{T}Q=r^{\pi}+\gamma\mathcal{P}^{\pi}Q$, where $r^{\pi}$ denotes the expected immediate reward and $\mathcal{P}^{\pi}$ represents the transition probability operator associated with the policy $\pi$. For convenience, we remove the superscript $\pi$ for the rest of the paper.
Using temporal difference (TD)[23] learning, the action-value estimate $\hat{Q}_{\gamma}$ can approximate the true action-value function $Q_{\gamma}$. The one-step TD error $\delta^{\gamma}_{t}=r_{t+1}+\gamma\hat{Q}_{\gamma}(s_{t+1},a_{t+1})\mathit{-}\hat{Q}_{\gamma}(s_{t},a_{t})$ is used to update the action-value function given a transition $(s_{t},a_{t},r_{t},s_{t+1})$.
SARSA is an on-policy RL approach that acquires the action-value function Q(s,a) for a specified policy $\pi$. The action-value function denotes the anticipated cumulative reward originating from state s, executing action a, and following policy $\pi$. Subsequently, the SARSA update rule is articulated as follows:
$Q(s_{t},a_{t})\leftarrow Q(s_{t},a_{t})+\alpha\left[r_{t}+\gamma Q(s_{t+1},a_{t+1})-Q(s_{t},a_{t})\right]$ (1)
where $\alpha$ denotes the learning rate, $\gamma$ represents the discount factor, and $r_{t}$ signifies the immediate reward. Long-horizon tasks present challenges for SARSA since the choice of $\gamma$ can compromise learning efficiency with appropriate long-term reward maximizing.
On the other hand, for a complete trajectory, we can employ the discounted sum of one-step TD errors, generally known as the $\lambda$-return[23] or, equivalently, the Generalized Advantage Estimator (GAE)[24]. The GAE enhances advantage estimates by reconciling the trade-off between variance and bias via parameters $\lambda$ and $\gamma$. The equation for the Generalized Advantage Estimator is generally expressed as:
$A(s_{t},a_{t})=\sum_{k=0}^{\infty}(\lambda\gamma)^{k}\delta_{t+k}^{\gamma}$ (2)
Where $\delta_{t+k}$ is the TD error at time t, computed as follows: $\delta_{t+k}=r_{t+k}+\gamma Q(s_{t+k+1},a_{t+k+1})-Q(s_{t+k},a_{t+k})$
Loss function for Q-Value estimation utilizing GAE. The loss function $\mathcal{L}(\theta)$ for approximating the Q-value function is defined as the mean squared error between the current Q-value estimate $Q(s_{t},a_{t};\theta)$ and the target Q-value adjusted by the advantage estimator. Thus, it is possible to concisely write the loss function using $A(s_{t},a_{t})$ and $Q(s_{t},a_{t})$ as follows:
$\mathcal{L}(\theta)=\mathbb{E}\bigg{[}\bigg{(}Q(s_{t},a_{t};\theta)-\bigg{(}Q(s_{t},a_{t})+A(s_{t},a_{t})\bigg{)}\bigg{)}^{2}\bigg{]}$ (3)
Despite SARSA being on-policy and lacking an explicit policy update, policy selection in SARSA can still be influenced by Eq.4. For actor-critic architectures[25, 26, 27], the action-value function is updated according to Eq.3, and a stochastic parameterized policy (actor, $\pi_{\omega}(a|s)$) is learned from this value estimator through the advantage function, with the loss being.
$\mathcal{L}(\omega)=\mathbb{E}\bigg{[}-log\pi(a|s;\omega)A(s,a)\bigg{]}$ (4)
In on-policy methods like SARSA, the likelihood ratio $\rho$ between the current policy and a prior policy is typically not employed. SARSA evaluates actions in accordance with the current policy, concurrently updating the action-value function as actions are performed. Thus, Proximal Policy Optimization (PPO)[28] does not implement a clipping objective function in SARSA as used in this study[8]. Eq.4 is utilized as the objective function.
4 Methodology
The TD($\Delta$) framework enhances conventional temporal difference learning by partitioning action-value functions across several discount factors $\gamma_{0},\gamma_{1},...,\gamma_{z}$. This decomposition enables the action-value function to be acquired as a summation of delta estimators, each representing the disparity between action-value functions at varying discount factors. The principal advantage of this method is the capacity to regulate variance and bias in the learning process by concentrating on smaller, more manageable time-scales.
4.1 TD($\Delta$) Framework
To calculate $W_{z}$ for SARSA as outlined in the study, it is essential to comprehend how delta estimators ($W_{z}$) approximate the discrepancies between action-value functions at consecutive discount factors. Let us delineate this approach for SARSA in a stepwise manner.
Delta Estimators ($W_{z}$) and Action-Value Functions ($Q(s,a)$): Delta estimators, $W_{z}$, represent the difference between value functions with successive discount factors:
$W_{z}=Q_{\gamma_{z}}\mathit{-}Q_{\gamma_{z-1}}$ (5)
In this context, $Q_{\gamma_{z}}$ denotes the action-value function with discount factor $\gamma_{z}$, but $Q_{\gamma_{z-1}}$ signifies the action-value function with the prior discount factor, $\gamma_{z-1}$.
4.2 Single-Step TD ( SARSA($\Delta$) )
This section begins with a description of the delta estimators ($W_{z}$) for the action-value functions $Q$ utilized in SARSA. The extension of TD($\Delta$) to SARSA is denoted as SARSA($\Delta$). SARSA($\Delta$) aims to obtain the action-value function $Q(s,a)$, representing the expected return of adhering to policy $\pi$ from state $s$, performing action $a$, and subsequently continuing to follow $\pi$. The aim is to improve the overall reward through the optimization of the action-value function across different time-scales. The core principle in SARSA($\Delta$) involves dividing the action-value function $Q(s,a)$ into a series of delta components $W_{z}(s,a)$, with each component associated with a specific discount factor $\gamma_{z}$. The interrelation among these delta components (i.e., delta function) is articulated as follows:
$W_{z}(s,a):=Q_{\gamma_{z}}(s,a)\mathit{-}Q_{\gamma_{z-1}}(s,a)$ (6)
where $\gamma_{0},\gamma_{1},...,\gamma_{z}$ represent the discount factors across various time-scales, and define $W_{0}(s,a):=Q_{\gamma_{0}}(s,a)$. The action-value function $Q_{\gamma_{z}}(s,a)$ is defined as the cumulative sum of all W-components up to z:
$Q_{\gamma_{z}}(s,a)=\sum_{j=0}^{z}W_{j}(s,a)$ (7)
The update rule in standard SARSA is determined by the state-action pair in accordance with the policy. The action-value function is updated through the TD error, with the action in the subsequent state selected in accordance with the policy:
$Q(s_{t},a_{t})\leftarrow Q(s_{t},a_{t})+\alpha\left[r_{t}+\gamma Q(s_{t+1},a_{t+1})-Q(s_{t},a_{t})\right]$ (8)
Here, $r_{t}$ denotes the reward obtained from executing action $a_{t}$ in state $s_{t}$, while $(s_{t+1},a_{t+1})$ represents the subsequent state-action pair. The SARSA update rule is applied individually to each delta component $W_{z}(s,a)$. The update for each component employs the same framework as the traditional SARSA method, yet it is adjusted to incorporate the delta function at each specific time-scale. We can express the update Eq.8 for single-step TD SARSA ($\Delta$) with multiple time-scales by applying the concept of separating the update into multiple discount factors[8] as follows:
$W_{z}(s_{t},a_{t})=\mathbb{E}\left[(\gamma_{z}-\gamma_{z-1})Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})+\gamma_{z}W_{z}(s_{t+1},a_{t+1})\right]$ (9)
where $Q_{\gamma_{z}}(s_{t},a_{t})$ represents the action-value function with a discount factor $\gamma_{z}$. $W_{z}(s_{t},a_{t})$ denotes the delta function for SARSA ($\Delta$), while $\gamma_{z}$ represents the discount factor for time-scale z, while $Q_{\gamma_{z}}(s_{t+1},a_{t+1})$ is the SARSA update for action-value associated with the subsequent state-action pair. The single-step TD SARSA ($\Delta$) mirrors the original TD update, but it is applied to action-value functions Q rather than state-value functions V. The single-step Bellman Eq. for the Q-value function in standard SARSA is as follows:
$Q_{\gamma_{z}}(s_{t},a_{t})=\mathbb{E}\left[r_{t}+\gamma_{z}Q_{\gamma_{z}}(s_{t+1},a_{t+1})\right]$ (10)
This denotes the expected value of the reward acquired at time t, in addition to the discounted value of the action-value function in the subsequent state-action pair $(s_{t+1},a_{t+1})$. Currently, rather than utilizing a singular discount factor $\gamma$, we are employing a sequence of discount factors $\gamma_{0},\gamma_{1},.....,\gamma_{z}$. The delta function $W_{z}(s_{t},a_{t})$ is defined as the difference between action-value functions associated with successive discount factors from Eq. 6:
$W_{z}(s_{t},a_{t})=Q_{\gamma_{z}}(s_{t},a_{t})-Q_{\gamma_{z-1}}(s_{t},a_{t})$ (11)
To expand the delta function $W_{z}$ utilizing its definition, we replace the Bellman Eq. for both $Q_{\gamma_{z}}$ and $Q_{\gamma_{z-1}}$, thereby expressing $W_{z}$ in relation to these action-value functions. Commence by composing:
[…]
Since the reward terms $r_{t}$ are included in both formulations, simplify by eliminating them.
$W_{z}(s_{t},a_{t})=\mathbb{E}\left[\gamma_{z}Q_{\gamma_{z}}(s_{t+1},a_{t+1})-\gamma_{z-1}Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})\right]$
Decomposing the terms allows for a more detailed analysis of the expression. Utilizing the recursive relationship from Eq. 11
$Q_{\gamma_{z}}(s_{t+1},a_{t+1})=W_{z}(s_{t+1},a_{t+1})+Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})$
Substitute the definition of $W_{z}$ into the Eq. below.
$W_{z}(s_{t},a_{t})=\mathbb{E}\left[\gamma_{z}(W_{z}(s_{t+1},a_{t+1})+Q_{\gamma_{z-1}}(s_{t+1},a_{t+1}))-\gamma_{z-1}Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})\right]$
Make the terms simpler:
$W_{z}(s_{t},a_{t})=\mathbb{E}\left[\gamma_{z}W_{z}(s_{t+1},a_{t+1})+(\gamma_{z}-\gamma_{z-1})Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})\right]$
Integrate the terms to derive the final update Eq. for $W_{z}(s_{t},a_{t})$.
$W_{z}(s_{t},a_{t})=\mathbb{E}\left[(\gamma_{z}-\gamma_{z-1})Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})+\gamma_{z}W_{z}(s_{t+1},a_{t+1})\right]$ (13)
This Eq. demonstrates that the update for $W_{z}$ integrates the disparity in discount factors $\gamma_{z}-\gamma_{z-1}$ with respect to the action-value function $Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})$ and the bootstrapping from the subsequent step $W_{z}(s_{t+1},a_{t+1})$. This denotes a Bellman Eq. for $W_{z}$, combining a decay factor $\gamma_{z}$ and the reward $Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})$. Consequently, it can be utilized to define the expected TD update for $W_{z}$. In this expression, $Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})$ can be expressed as the summation of $W_{z}(s_{t+1},a_{t+1})$ for $z\leq Z-1$, indicating that the Bellman Eq. for $W_{z}$ is contingent upon the values of all delta functions $W_{z}$ for $z\leq Z-1$.
This approach treats the delta value function at each time-scale as an autonomous RL issue, with rewards obtained from the action-value function of the immediately preceding time-scale. Thus, for a target discounted action-value function $Q_{\gamma_{z}}(s,a)$, all delta components can be trained simultaneously through a TD update, utilizing the prior values of each estimator for bootstrapping. This process necessitates the assumption of a sequence of discount factors, denoted as $\gamma_{z}$, which includes both the minimum and maximum values, $\gamma_{0}$ and $\gamma_{z}$[8].
4.3 Multi-Step TD ( SARSA($\Delta$) )
Numerous studies indicate that multi-step TD methods typically exhibit greater efficiency compared to single-step TD methods [29]. In multi-step TD SARSA($\Delta$), rewards are aggregated over multiple steps instead of depending on a single future reward. This approach takes into account the differences between consecutive discount factors, $\gamma_{z}$ and $\gamma_{z-1}$, while utilizing the value function at state $s_{t+k_{z}}$ for bootstrapping. In SARSA($\Delta$), the update Eqs. are modified to aggregate rewards across a series of transitions prior to bootstrapping from either the Q-values or delta estimators. In multi-step SARSA($\Delta$), the agent aggregates rewards over K steps, leading to a corresponding modification in the update rule for each $W_{z}$.
$W_{z}(s_{t},a_{t})=\mathbb{E}\left[\sum_{j=1}^{k_{z}-1}(\gamma_{z}^{j}-\gamma_{z-1}^{j})r_{t+j}+(\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}})Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})+\gamma_{z}^{k_{z}}W_{z}(s_{t+k_{z}},a_{t+k_{z}})\right]$ (14)
In this case, $r_{t+j}$ represents the reward obtained at time step $t+j$. $a_{t+k_{z}}$ and the action is chosen greedily as $a=\arg\max_{a}Q(s_{t+k_{z}},a)$, where $k_{z}$ is the number of steps linked to the discount factor $\gamma_{z}$. This method generalizes Q-learning to multi-step TD learning by aggregating rewards across several stages and utilizing bootstrapping from both the current and prior time-scale action-value functions. In normal multi-step TD learning, the TD error is computed by aggregating rewards over several stages rather than relying on a single future step, followed by bootstrapping from the action-value at the last step. For example, utilizing a singular discount factor $\gamma$, the multi-step TD update equation is expressed as:
$Q(s_{t},a_{t})=\mathbb{E}\left[\sum_{j=0}^{k-1}\gamma^{j}r_{t+j}+\gamma^{k}Q(s_{t+k},a_{t+k})\right]$ (15)
Herein, The initial segment of the Eq. aggregates the rewards across k steps, each diminished by $\gamma$. The second component derives from the action-value function at time step t+k. We expand this methodology to include various discount factors and delta decomposition, concentrating on representing the delta function $W_{z}(s_{t},a_{t})$ as the difference between action-value functions with discount factors $\gamma_{z}$ and $\gamma_{z-1}$. According to the definition of $W_{z}$, we obtain:
$W_{z}(s_{t},a_{t})=Q_{\gamma_{z}}(s_{t},a_{t})-Q_{\gamma_{z-1}}(s_{t},a_{t})$
We now implement the multi-step variant of the Bellman Eq. for both $Q_{\gamma_{z}}(s_{t},a_{t})$ and $Q_{\gamma_{z-1}}(s_{t},a_{t})$. The multi-step Bellman Eq. for $Q_{\gamma_{z}}(s_{t},a_{t})$ is expressed as follows:
$Q_{\gamma_{z}}(s_{t},a_{t})=\mathbb{E}\left[\sum_{j=0}^{k_{z}-1}\gamma_{z}^{j}r_{t+j}+\gamma_{z}^{k_{z}}Q_{\gamma_{z}}(s_{t+k_{z}},a_{t+k_{z}})\right]$
For $Q_{\gamma_{z-1}}(s_{t},a_{t})$, the multi-step Bellman Eq. is expressed as follows:
$Q_{\gamma_{z-1}}(s_{t},a_{t})=\mathbb{E}\left[\sum_{j=0}^{k_{z}-1}\gamma_{z-1}^{j}r_{t+j}+\gamma_{z-1}^{k_{z}}Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})\right]$
We will now utilize the delta estimator formulas to subtract these two Eqs.
$W_{z}(s_{t},a_{t})=\left[\sum_{j=0}^{k_{z}-1}\gamma_{z}^{j}r_{t+j}+\gamma_{z}^{k_{z}}Q_{\gamma_{z}}(s_{t+k_{z}},a_{t+k_{z}})\right]-\left[\sum_{j=0}^{k_{z}-1}\gamma_{z-1}^{j}r_{t+j}+\gamma_{z-1}^{k_{z}}Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})\right]$
We will now broaden and simplify the expressions. The immediate reward terms $r_{t+j}$ are present in both Eqs. but are adjusted by distinct discount factors, $\gamma_{z}$ and $\gamma_{z-1}$. This enables us to express the disparity in rewards as:
$\sum_{j=0}^{k_{z}-1}\gamma_{z}^{j}r_{t+j}-\sum_{j=0}^{k_{z}-1}\gamma_{z-1}^{j}r_{t+j}=\sum_{j=0}^{k_{z}-1}(\gamma_{z}^{j}-\gamma_{z-1}^{j})r_{t+j}$
Subsequently, for the bootstrapping terms, we utilize the recursive relationship as outlined:
$Q_{\gamma_{z}}(s_{t+k_{z}},a_{t+k_{z}})=W_{z}(s_{t+k_{z}},a_{t+k_{z}})+Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})$
As a result, we get:
$\gamma_{z}^{k_{z}}Q_{\gamma_{z}}(s_{t+k_{z}},a_{t+k_{z}})-\gamma_{z-1}^{k_{z}}Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})=(\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}})Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})+\gamma_{z}^{k_{z}}W_{z}(s_{t+k_{z}},a_{t+k_{z}})$
We now amalgamate the terms for both the rewards and bootstrapping to produce the final expression:
$W_{z}(s_{t},a_{t})=\mathbb{E}\left[\sum_{j=0}^{k_{z}-1}(\gamma_{z}^{j}-\gamma_{z-1}^{j})r_{t+j}+(\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}})Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})+\gamma_{z}^{k_{z}}W_{z}(s_{t+k_{z}},a_{t+k_{z}})\right]$ (16)
This Eq. extends the conventional multi-step TD update to include various multiple discount factors and action-value functions in RL. As a result, each $W_{z}$ obtains a share of the rewards from the environment up to time-step $k_{z}-1$. Furthermore, each $W_{z}$ utilizes its distinct action-value function in conjunction with the value from the prior time-scale. A variation of this algorithm, utilizing k-step bootstrapping as described in [29], is presented in Algorithm 1. Despite Algorithm 1 exhibiting quadratic complexity in relation to Z, it can be executed with linear complexity for substantial Z by preserving $\hat{Q}$ action-values at each time-scale $\gamma_{z}$.
Algorithm 1 Multi-Step TD ( SARSA($\Delta$) )
Inputs: Pick out the discount factors ($\gamma_{0},\gamma_{1},...,\gamma_{Z}$), bootstrapping steps ($k_{0},k_{1},...,k_{Z}$), and learning rates ($\alpha_{0},\alpha_{1},...,\alpha_{Z}$).
Set the initial value of $W_{z}(s,a)=0$ for all states, actions, and scales z.
for episode = 0, 1, 2,… $\triangleright$ Loop for each episode
     Initialize state $s_{0}$ and choose an initial action $a_{0}$ in accordance with a policy.
     for t = 0, 1, 2,… do $\triangleright$ Loop for each time step
         Take action $a_{t}$, observe reward $r_{t}$ and next state $s_{t+1}$.
         Select action $a_{t+1}$ based on a policy.
         for z = 0, 1, …, Z do
              if z = 0 then
                  $G^{0}=\sum_{j=0}^{k_{0}-1}\gamma_{0}^{j}r_{t+j}+\gamma_{0}^{k_{0}}W_{0}(s_{t+k_{0}},a_{t+k_{0}})$
              $\triangleright$ Utilizing Eq.7, we substitute $Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})$ by summing the W-components up to $W_{z-1}$ in Eq.4.3.
                  $G^{z}=\sum_{j=0}^{k_{z}-1}(\gamma_{z}^{j}-\gamma_{z-1}^{j})r_{t+j}+(\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}})\sum_{z=0}^{z-1}W_{z-1}(s_{t+k_{z}},a_{t+k_{z}})+\gamma_{z}^{k_{z}}W_{z}(s_{t+k_{z}},a_{t+k_{z}})$
              end if
         end for
         for z = 0, 1, 2,…, Z do
              $W_{z}(s_{t},a_{t})\leftarrow W_{z}(s_{t},a_{t})+\alpha_{z}(G^{z}-W_{z}(s_{t},a_{t}))$
         end for
     end for
end for
4.4 SARSA TD($\lambda,\Delta$)
Eq.17 introduces the $\lambda$-return[23, 1], which combines rewards over multiple steps to create a target for TD($\lambda$) updates. The $\lambda$-return $G_{t}^{\gamma,\lambda}$ is defined as:
$G_{t}^{\gamma,\lambda}(s_{t},a_{t})=\hat{Q}_{\gamma}(s_{t},a_{t})+\sum_{k=0}^{\infty}(\lambda\gamma)^{k}\delta_{t+k}^{\gamma}$ (17)
Eq.18 defines the TD($\lambda$) operator, which is used to iteratively apply $\lambda$-returns in updating the value functions. For SARSA, the TD($\lambda$) operator updates the action-value function by summing the $\lambda$-discounted TD errors as follows:
$\mathrm{T}_{\lambda}Q(s_{t},a_{t})=Q(s_{t},a_{t})+(I-\lambda\gamma P)^{-1}(\mathrm{T}Q(s_{t},a_{t})-Q(s_{t},a_{t}))$ (18)
Here,P is the transition matrix for state-action pairs under the policy $\pi$, and Q is the action-value function. Similarly, Eq.19 defines the $\lambda$-return specific to delta estimators in TD($\lambda,\Delta$), denoted as $G_{t}^{z,\lambda_{z}}$ for each delta estimator $W_{z}$:
$G_{t}^{z,\lambda_{z}}:=\hat{W}_{z}(s_{t},a_{t})+\sum_{k=0}^{\infty}(\lambda_{z}\gamma_{z})^{k}\delta_{t+k}^{z}$ (19)
where $\delta_{t}^{0}:=\delta_{t}^{\gamma_{0}}$ and $\delta_{t}^{z}:=(\gamma_{z}-\gamma_{z-1})\hat{Q}_{\gamma_{z-1}}(s_{t+1},a_{t+1})+\gamma_{z}\hat{W}_{z}(s_{t+1},a_{t+1})-\hat{W}_{z}(s_{t+1},a_{t+1})$ are the TD-errors.
4.5 SARSA TD($\lambda,\Delta$) with Generalized Advantage Estimation (GAE)
Generalized Advantage Estimation (GAE) [28] seeks to compute the advantage function by aggregating multi-step TD errors. To apply this approach to delta estimators $W_{z}(s_{t},a_{t})$, we compute advantage estimates $A^{\Delta}(s_{t},a_{t})$ for each time-scale. Each advantage estimate $A^{\Delta}(s_{t},a_{t})$ employs a multi-step TD error pertinent to the delta estimator $W_{z}$, expressed as follows:
$A^{\Delta}(s_{t},a_{t})=\sum_{k=0}^{T-1}(\lambda_{z}\gamma_{z})^{k}\delta_{t+k}^{\Delta}$ (20)
where $\delta_{t+k}^{\Delta}:=r_{t}+\gamma_{z}\sum_{z=0}^{Z}\hat{W}_{z}(s_{t+1},a_{t+1})-\sum_{z=0}^{Z}\hat{W}_{z}(s_{t},a_{t})$.
We utilize $\gamma_{Z}$ as the discount factor and the sum of all W estimators as a surrogate for $Q_{\gamma z}$. This objective is applicable to PPO by implementing the policy update from Eq.4 and replacing A with $A^{\Delta}$. Additionally, to train each $W_{z}$, we utilize a truncated form of their corresponding $\lambda$-return, as outlined in Eq.19. For more information, see Algorithm 2.
Algorithm 2 PPO-TD($\lambda$, SARSA($\Delta$))
Inputs: Pick out the discount factors ($\gamma_{0},\gamma_{1},...,\gamma_{Z}$), bootstrapping steps ($k_{0},k_{1},...,k_{Z}$), and learning rates ($\alpha_{0},\alpha_{1},...,\alpha_{Z}$).
Set the initial value of $W_{z}(s,a)=0$ for all states s, actions a, and scales z.
Initialize policy $\omega$, and values $\theta^{z}\forall z$
for episode = 0, 1, 2,… $\triangleright$ Loop for each episode
     Initialize state $s_{0}$ and choose an initial action $a_{0}$ in accordance with a policy.
     for t = 0, 1, 2,… do $\triangleright$ Loop for each time step
         Take action $a_{t}$, observe reward $r_{t}$ and next state $s_{t+1}$.
         Select next action $a_{t+1}$ based on a policy.
         for z = 0, 1, …, Z do
              if $t\geq T$ then
                  $G^{z,\lambda_{z}}\leftarrow\hat{W}_{z}(s_{t-T},a_{t-T})+\sum_{k=0}^{T-1}(\lambda_{z}\gamma_{z})^{k}\delta_{t-T+k}^{z}\forall z$triangleright Computing multi-step return $G^{z,\lambda_{z}}$ and TD-error $\delta_{t-T+k}^{z}$ using Eq.19.
              end if
         end for
         for z = 0, 1, 2,…, Z do
              $\hat{W}_{z}(s_{t-T},a_{t-T})\leftarrow\hat{W}_{z}(s_{t-T},a_{t-T})+\alpha_{z}\left(G^{z,\lambda_{z}}-\hat{W}_{z}(s_{t-T},a_{t-T})\right)$
              $A^{\Delta}=\sum_{k=0}^{T-1}(\lambda_{z}\gamma_{z})^{k}\delta_{t-T+k}^{\Delta}$triangleright Where $A^{\Delta}$ and $\delta_{t-T+k}^{\Delta}$ are computed using Eq.20.
              $\theta^{z}\leftarrow\theta^{z}+\alpha_{z}\left(G^{z,\lambda_{z}}-\hat{W}_{z}(s_{t-T},a_{t-T})\right)\nabla\hat{W}_{z}(s_{t-T},a_{t-T})$triangleright Update $\theta^{z}$ with TD (Eq.3) using $G^{z,\lambda_{z}}\forall z$.
              $\omega\leftarrow\omega+\alpha_{\omega}A^{\Delta}\nabla\log\pi_{\omega}(a_{t}|s_{t})$triangleright Update the policy parameters $\omega$ with PPO (Eq.4) for SARSA using $A^{\Delta}$
         end for
     end for
end for
5 Analysis
Next, We examine our delta estimators through the lens of the bias-variance trade-off. In SARSA, bias emerges from dependence on current estimates, affecting updates over short time-scales, particularly with smaller discount factors, such as lower values of $\gamma$ in $W_{z}$. Updates utilizing smaller discount factors demonstrate reduced variance due to their reliance on current estimates; however, they exhibit increased bias as they fail to integrate substantial information regarding future rewards. In contrast, variance escalates for updates associated with extended time-scales, as future reward information adds further variability. In TD($\Delta$), action-value components $W_{z}$ with higher $\gamma_{z}$ are associated with greater long-term rewards; however, this results in increased variance stemming from the stochastic characteristics of reward sequences. We begin by demonstrating that our estimator is equivalent to the standard estimator $\hat{Q}_{\gamma}$ under specific conditions, as outlined in Theorem 1. This comparison elucidates the essential metrics of our estimator that may indicate potential advantages over the standard $\hat{Q}_{\gamma}$ estimator. Building on this result and previous research by Kearns and Singh[9], we modify these analyses for SARSA to investigate the effects of bias and variance, along with the TD($\Delta$) decomposition framework, in relation to action-value functions. Our objective is to extend the bias-variance error bound framework to action-value settings using TD($\Delta$)(i.e., Theorem 4), thereby offering a better understanding of how these quantities can be balanced to obtain optimal results[8].
5.1 SARSA Equivalency Configurations and Enhancements
In specific situations, we can show that our delta estimators for SARSA align with the conventional action-value estimator when restructured into an action-value function. This discussion centers on linear function approximation of the specified form:
$\hat{Q}(s,a)_{\gamma}:=\langle\theta^{\gamma},\phi(s,a)\rangle\quad\text{and}\quad\hat{W}_{z}(s,a):=\langle\theta^{z},\phi(s,a)\rangle,\forall z$
where $\theta$ and $\theta^{z}$ represent weight vectors in $\mathbb{R}^{d}$ and the function $\phi:S\times A\rightarrow\mathbb{R}^{d}$ represents a mapping from a state-action pair to a specified d-dimensional feature space. The weight vector $\theta$ for SARSA is updated according to the TD($\lambda$) learning rule in the following manner:
$\theta_{t+1}^{\gamma}=\theta_{t}^{\gamma}+\alpha\bigg{(}G_{t}^{\gamma,\lambda}-\hat{Q}_{\gamma}(s,a)\bigg{)}\phi(s_{t},a_{t}),$ (21)
where $G_{t}^{\gamma,\lambda}$ represents the TD($\lambda$) return defined in Eq.17. Likewise, the delta estimator approach TD($\lambda_{z},\Delta$) is used to update each $\hat{W}_{z}$:
$\theta_{t+1}^{z}=\theta_{t}^{z}+\alpha\bigg{(}G_{t}^{Z,\lambda_{z}}-\hat{W}_{z}(s_{t},a_{t})\bigg{)}\phi(s_{t},a_{t}),$ (22)
where $G_{t}^{Z,\lambda_{z}}$ is specified identically to the TD($\Delta$) return defined in Eq.19, modified for the particular action-value function. In these equations, $\alpha$ and $\{\alpha_{z}\}_{z}$ represent positive learning rates. The following theorem demonstrates the equivalence of two SARSA version algorithms in this context. The subsequent theorems resemble those presented by Romoff et al. [8]; however, we offer a proof concerning the action-value function in relation to SARSA, whereas they established the proof for the value function.
Theorem 1.
If $\alpha_{z}=\alpha$, $\lambda_{z}\gamma_{z}=\lambda_{\gamma}$, $\forall z$, and if we choose the initial conditions in such a way that $\sum_{z=0}^{Z}\theta_{0}^{\gamma}=\theta_{0}^{\gamma}$, then the iterates produced by TD($\lambda$) (Eq. 21) and TD($\lambda,\Delta$) (Eq. 22) with linear function approximation satisfy the following conditions:
$\Sigma_{z=0}^{Z}\theta_{t}^{z}=\theta_{t}^{\gamma}$ (23)
The proof is presented in appendix PROOFS.
Equivalence is attained when $\lambda_{z}\gamma_{z}=\lambda_{\gamma},\forall z$. When $\lambda$ approaches 1 and $\gamma_{z}$ is less than $\gamma$, this condition suggests that $\lambda_{z}=\lambda_{\gamma}/\gamma_{z}$ could potentially surpass one, resulting in a risk of divergence in TD($\lambda_{z}$). The subsequent theorem demonstrates that the TD($\lambda$) operator, as defined in Eq.18, qualifies as a contraction mapping for the range $1\leq\lambda<\frac{1+\gamma}{2\gamma}$, thereby confirming that $\lambda_{\gamma}<1$[8].
Theorem 2.
$\forall\lambda\in\left[0,\frac{1+\gamma}{2\gamma}\right]$, the operator $T_{\lambda}$ is defined by the Eq. $T_{\lambda}Q=Q+(I-\lambda\gamma P)^{-1}(TQ-Q)$ $,\quad\forall Q\in\mathbb{R}^{|S|\times|A|}$, is well-defined. Furthermore, $T_{\lambda}Q$ constitutes a contraction in relation to the max norm, with its contraction coefficient expressed as $\frac{\gamma}{|1-\lambda\gamma|}$.
The proof is presented in appendix PROOFS.
The analysis presented in Theorem 1 can be adapted to a new context in which $k_{z}$-step TD ($\Delta$) is employed for each $W_{z}$ component, rather than TD($\lambda$, $\Delta$), a variant of TD learning. Theorem 1 demonstrates that with linear function approximation, standard multi-step TD and multi-step TD($\Delta$) can be equivalent when the number of steps ($k_{z}$) is consistent across all time scales (i.e., $k_{z}=k,\forall z$).
Although these methods could theoretically be equivalent, such equivalence is not common. Generally, for equivalence to be maintained, the learning rate must remain consistent across all time scales. This represents a significant limitation, as shorter time scales, which involve fewer future steps, can be learned more rapidly than longer time scales that require consideration of additional steps ahead. Furthermore, in practical applications, particularly involving nonlinear function approximation such as deep neural networks, adaptive optimizers are frequently employed[30, 28]. These optimizers modify the learning rate in accordance with the complexity of the learning task, which is contingent upon the characteristics of the delta estimator and its target. The effective learning rate varies across different time scales, making it challenging to apply a uniform learning rate in this context.
In addition to the learning rate, the approach of decomposing the action-value function into various components (specifically, $W_{z}$ components) presents advantages that traditional TD learning, including non-delta estimators or higher $W_{z}$ components, lacks. The ability to utilize various k-step returns (or $\lambda$-return) across different time scales provides enhanced control and adaptability in the learning process. If $k_{z}<k_{z+1},\forall z\quad(or\quad\gamma_{z}\lambda_{z}<\gamma_{z+1}\lambda_{z+1},\forall z)$, this approach may decrease variance while introducing bias, as shorter time scales, which can be learned more rapidly, may not align perfectly with longer time scales in terms of the values they estimate.
5.2 Evaluation for Minimizing $k_{z}$ Values in SARSA via Phased Updates
To show that how our approach differ from the single estimator case, let’s follow the tabular phased version of k-step TD presented by kearns & Singh[9]. In SARSA, we aim to estimate the value of a state-action pair (s, a) start from each state $s\in\mathcal{S}$ and with action $a\in\mathcal{A}$, then n trajectories generated $\bigg{\{}S_{0}^{(j)}=s,a_{0},\gamma_{0},...,S_{k}^{(j)},a_{k}^{(j)},\gamma_{k}^{(j)},S_{k+1}^{(j)},...\bigg{\}}_{1\leq j\leq n}$ by following the policy $\pi$ and averaging over trajectories. Let $Q_{\gamma,t}(s,a)$ denote the Q-value estimate at phase t for the discount factor $\gamma$. For each iteration t, also called phase t, to get the phase-based estimate $Q_{\gamma,t}(s,a)$ for (s, a), we average over the trajectories, resulting in:
$\hat{Q}_{\gamma,t}(s,a)=\frac{1}{n}\sum_{j=1}^{n}\bigg{(}\sum_{i=0}^{k-1}\gamma^{i}r_{i}^{(j)}+\gamma^{k}\hat{Q}_{\gamma,t-1}(s_{k}^{(j)},a_{k}^{(j)})\bigg{)}$ (24)
Theorem 3 below from the theorems of kearns & Singh[9] but they show it for state value function while we show an upper bound on the error for action-value function estimation defined by $\Delta_{t}^{\hat{Q}_{\gamma}}=max_{s,a}\bigg{\{}\bigg{|}\hat{Q}_{\gamma,t}(s,a)-Q_{\gamma}(s,a)\bigg{|}\bigg{\}}$
Theorem 3.
Inspired by Kearns & Singh[9], for any $10<\delta<$ 1, let $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$. with probability $1-\delta$,
$\Delta_{t}^{\hat{Q}_{\gamma}}\leq\epsilon\bigg{(}\underbrace{\frac{1-\gamma^{k}}{1-\gamma}\bigg{)}}_{\text{variance term}}+\underbrace{\gamma^{k}\Delta_{t-1}^{\hat{Q}_{\gamma}}}_{\text{bias term}}$ (25)
The proof is presented in appendix PROOFS.
In Eq. 25, A variance term ($\frac{1-\gamma^{k}}{1-\gamma}$) is arising due to sampling error from rewards collected along trajectories. In particular, $\epsilon$ bounds the deviation of the empirical average of rewards from the true expected reward. Second term ($\gamma^{k}\Delta_{t-1}^{\hat{Q}_{\gamma}}$) in Eq. 25 is a bias term that arises from bootstrapping off the previous phase’s estimates. Similarly, we consider a phased SARSA version of multi-step $TD(\Delta)$. For each phase t, we update each W as follows:
$\hat{W}_{z,t}(s,a)=\frac{1}{n}\sum_{j=1}^{n}\bigg{(}\sum_{i=1}^{k-1}(\gamma_{z}^{i}-\gamma_{z-1}^{i})r_{i}^{(j)}+(\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}})Q_{\gamma_{z-1}}(s_{t+k}^{(j)},a_{t+k}^{(j)})+\gamma_{z}^{k_{z}}\hat{W}_{z}(s_{t+k}^{(j)},a_{t+k}^{(j)})$ (26)
We now determine the phased error’s upper bound. The sum of the errors made by each W component $\sum_{z=0}^{Z}\Delta_{t}^{z}$ is known as $TD(\Delta)$, where $\Delta_{t}^{z}=max_{s,a}\bigg{\{}\bigg{|}\hat{W}_{z}(s,a)-W_{z}(s,a)\bigg{|}\bigg{\}}$. Key distinctions between SARSA and Q-learning in terms of the error bound: In SARSA, the error is determined by the agent’s actual actions, which means that the Q-values are updated according to the state-action pairs that the agent visits during exploration. The error in Q-learning is derived from the greedy action, which is the action that maximizes the Q-value for the next state, under the assumption that the agent always acts optimally.
Theorem 4.
Assume that $\gamma_{0}\leq\gamma_{1}\leq\gamma_{2}\leq.....\leq\gamma_{z}=\gamma$ and $K_{0}\leq K_{1}\leq...\leq K_{z}=K$, for any $10<\delta<$ 1, let $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$, with probability $1-\delta$,
$\sum_{z=0}^{Z}\Delta_{t}^{z}\leq\epsilon\bigg{(}\frac{1-\gamma^{k}}{1-\gamma}\bigg{)}+\epsilon\bigg{(}\underbrace{\sum_{z=0}^{Z-1}\frac{\gamma_{z}^{k_{z+1}}-\gamma^{k_{z}}_{z}}{1-\gamma_{z}}\bigg{)}}_{\text{variance reduction}}+\underbrace{\sum_{z=0}^{Z-1}\bigg{(}\gamma^{k_{z}}_{z}-\gamma^{k_{z+1}}_{z}\bigg{)}\sum_{u=0}^{z}\Delta_{t-1}^{u}}_{\text{bias introduction}}+\gamma^{k}\sum_{z=0}^{Z}\Delta_{t-1}^{z}$ (27)
The proof is presented in appendix PROOFS.
The author examines the comparison of the limits for phased TD($\lambda$) in theorem 3 with those for phased TD($\Delta$) in theorem 4. This comparison demonstrates that phased TD($\Delta$) facilitates variance reduction equivalent to $\epsilon\sum_{z=0}^{Z-1}\frac{\gamma_{z}^{k_{z+1}}-\gamma^{k_{z}}_{z}}{1-\gamma_{z}}\leq 0$ but introduces a potential bias quantified as $\sum_{z=0}^{Z-1}\bigg{(}\gamma^{k_{z}}_{z}-\gamma^{k_{z+1}}_{z}\bigg{)}\sum_{u=0}^{z}\Delta_{t-1}^{u}\geq 0$. Utilizing phased TD($\Delta$) to decrease $k_{z}$ synchronizes updates more effectively with recent actions, hence lowering variance from high-discounted return components. Shortened $k_{z}$ values accrue bias more swiftly, as SARSA bootstraps from current policy estimations that may vary over phases. It is observed that when all $k_{z}$ values are identical, both algorithms yield the same upper bound.
In SARSA, the return estimate encompasses subsequent states and actions dictated by the policy. The expected discounted return over T steps closely approximates the infinite-horizon discounted expected return after $\mathrm{T}$, where $\mathrm{T}\approx\frac{1}{1-\gamma}$[31]. Consequently, we can effectively simplify $k_{z}$ for any $\gamma_{z}$ such that $k_{z}\approx\frac{1}{1-\gamma_{z}}$, thereby adhering to this principle. Therefore, with $\mathrm{T}$ samples, similar to TD($\Delta$)[8], SARSA employing TD($\Delta$) utilizes $k_{z}\approx\frac{1}{1-\gamma_{z}}$ to establish a suitable balance between bias and variance over all time scales significantly less than $\mathrm{T}$. By establishing $k_{z}$ for each $\gamma_{z}$, where $\gamma_{z}^{\frac{1}{1-\gamma_{z}}}\leq\frac{1}{e}$, the approach guarantees that the effective horizon and variance are maintained within limits without requiring extensive parameter optimization. Each increase in $W_{z}$ facilitates the doubling of the effective horizon, hence ensuring logarithmic scaling with the quantity of action-value functions. SARSA can utilize this to modify additional parameters in alignment with the specified time scales, enhancing the algorithm’s efficiency in calculating long-term rewards without added complexity.
6 Experiments
We illustrate the efficacy of SARSA($\Delta$) via tests on multiple RL benchmarks, including the Ring MDP[9] and dense reward Atari environments. The findings indicate that SARSA($\Delta$) attains more rapid convergence and superior long-term efficacy relative to conventional SARSA and TD learning. All hyperparameter settings, extended details, and the reproducibility checklist for machine learning research[32] are included in the supplemental materials of this paper[8].
7 Conclusion
SARSA($\Delta$) expands SARSA by including multiple discount factors through TD($\Delta$), resulting in improved performance and stability. This approach for a more precise breakdown of future rewards while maintaining the benefits of on-policy learning. The Q-Learning version of TD($\Delta$) is a simple extension of this. We put it off for later.
Acknowledgments
Author contributions statement
Data availability
The data sets produced in this work can be obtained from the corresponding author upon an appropriate request.
Competing interests
The authors disclose no conflicting interests.
Appendix
PROOFS
Proof: Theorem 1
The proof utilizes induction. Initially, it is necessary to establish that the base case is valid at t=0. This statement is evidently valid given our initial assumption, particularly in the context of zero initialization. At a specific time-step t, we assume the validity of the statement, namely, $\sum_{z=0}^{Z}\theta_{t}^{z}=\theta_{t}^{\gamma}$. We will now demonstrate that it also holds at the subsequent time step, t+1.
$\sum_{z=0}^{Z}\theta_{t+1}^{z}$ $=\sum_{z=0}^{Z}\bigg{(}\theta_{t}^{z}+\alpha_{z}\bigg{(}G_{t}^{Z,\lambda_{z}}-\hat{W}_{z}(s_{t},a_{t})\bigg{)}\phi(s_{t},a_{t})\bigg{)}$
[…]
$=\theta_{t}^{\gamma}+\sum_{z=0}^{Z}\alpha_{z}\bigg{(}\sum_{k=t}^{\infty}(\lambda_{z}\gamma_{z})^{k-t}\delta_{k}^{z}\bigg{)}\phi(s_{t},a_{t})\text{ , Assumption based on induction.}$.
$=\theta_{t}^{\gamma}+\alpha\sum_{k=t}^{\infty}(\lambda\gamma)^{k-t}\underbrace{\bigg{(}\sum_{z=0}^{Z}\delta_{k}^{z}\bigg{)}}_{*}\phi(s_{t},a_{t})\text{ , Because of $\alpha_{z}=\alpha$, $\lambda_{z}\gamma_{z}=\lambda_{\gamma}$, $\forall z$}$
To show that $\sum_{z=0}^{Z}\theta_{t+1}^{z}=\theta_{t+1}^{\gamma}$, we must establish that the term (*) = $\sum_{z=0}^{Z}\delta_{k}^{z}$ is equivalent to the conventional TD error $\delta_{k}^{\gamma}$.
$\sum_{z=0}^{Z}\delta_{k}^{z}$ $=r_{k}+\gamma_{0}\hat{Q}_{\gamma_{0}}(s_{k+1},a_{k+1})-\hat{Q}_{\gamma_{0}}(s_{k},a_{k})+\sum_{z=1}^{Z}\bigg{(}(\gamma_{z}-\gamma_{z-1})\sum_{u=0}^{z-1}\langle\theta_{t}^{u},\phi(s_{k+1},a_{k+1})\rangle+\gamma_{z}\langle\theta_{t}^{z},\phi(s_{k+1},a_{k+1})\rangle-\langle\theta_{t}^{z},\phi(s_{k},a_{k})\rangle\bigg{)}$
$=r_{k}+\gamma_{0}\hat{Q}_{\gamma_{0}}(s_{k+1},a_{k+1})+\bigg{\langle}\sum_{z=1}^{Z}\bigg{(}\gamma_{z}\sum_{u=0}^{z}\theta_{t}^{u}-\gamma_{z-1}\sum_{u=0}^{z-1}\theta_{t}^{u}\bigg{)},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-\bigg{\langle}\sum_{z=0}^{Z}\theta_{t}^{z},\phi(s_{k},a_{k})\bigg{\rangle}$
$=r_{k}+\gamma_{0}\hat{Q}_{\gamma_{0}}(s_{k+1},a_{k+1})+\gamma_{z}\bigg{\langle}\sum_{z=0}^{Z}\theta_{t}^{z},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-\gamma_{0}\bigg{\langle}\theta_{t}^{0},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-Q_{\gamma}(s_{k},a_{k})$
$=r_{k}+\gamma_{0}\hat{Q}_{\gamma_{0}}(s_{k+1},a_{k+1})+\gamma\hat{Q}_{\gamma}(s_{k+1},a_{k+1})-\gamma_{0}\hat{Q}_{\gamma_{0}}(s_{k+1},a_{k+1})-\hat{Q}_{\gamma}(s_{k},a_{k})$
$=r_{k}+\gamma\hat{Q}_{\gamma}(s_{k+1},a_{k+1})-\hat{Q}_{\gamma}(s_{k},a_{k})$
$=\delta_{k}^{\gamma}$ (29)
Proof: Theorem 2
The following is the definition of the Bellman operator:
$\mathrm{T}=r+\gamma P$
The discount factor, the expected reward function, and the transition probability operator that is induced by the policy $\pi$ are denoted by the symbols $\gamma$, r, and P, respectively. A geometric weighted sum of $\mathrm{T}$ is the definition of the TD($\lambda$) operator, which can be expressed as follows:
$\mathrm{T}=(1-\lambda)\sum_{k=0}^{\infty}\lambda^{k}(\mathrm{T})^{k+1}$ (31)
It may be inferred that $\lambda\in[0,1]$ is necessary for the aforementioned sum to be finite. A corresponding definition of $\mathrm{T}_{\lambda}$ is as follows for every function W:
$\mathrm{T}_{\lambda}W=W+(I-\lambda\gamma P)^{-1}(\mathrm{T}W-W)$ (32)
The separated action-value function is denoted by W in this form, and the update is given in terms of the operator $\mathrm{T}_{\lambda}$, which is responsible for applying time-scale separation. In the event that $10\leq\lambda\gamma<$ 1, the formula is considered to be well-defined. This is because the spectral norm of the operator $\lambda\gamma P$ is less than 1, and thus, the Eq. $I-\lambda\gamma P$ is invertible. On the other hand, when the value of $\lambda$ is bigger than one, the equivalence between Eqs. 31 and 32 is lost. This is not a problem because, in fact, we use the TD error for training, which, according to our expectations, corresponds to the definition of $\mathrm{T}_{\lambda}$ that is provided in Eq. 32. Now, let’s have a look at the contraction property that the operator $\mathrm{T}_{\lambda}$ possesses. To begin, the Eq. designated as 32 can be rewritten as follows:
$\mathrm{T}_{\lambda}$ $=(I-\lambda\gamma P)^{-1}(\mathrm{T}W-\lambda\gamma PW)$ (33)
The contraction property is demonstrated by examining two action-value functions, $W_{1}$ and $W_{2}$. The distinction between the updates of $\mathrm{T}_{\lambda}W_{1}$ and $\mathrm{T}_{\lambda}W_{2}$ is:
$\mathrm{T}_{\lambda}W_{1}-\mathrm{T}_{\lambda}W_{2}$ $=(I-\lambda\gamma P)^{-1}\bigg{(}\mathrm{T}W_{1}-\mathrm{T}W_{2}-\lambda\gamma P(W_{1}-W_{2})\bigg{)}$
$=(I-\lambda\gamma P)^{-1}\bigg{(}\gamma P(W_{1}-W_{2})-\lambda\gamma P(W_{1}-W_{2})\bigg{)}$
$=(I-\lambda\gamma P)^{-1}\bigg{(}\gamma(1-\lambda)P(W_{1}-W_{2})\bigg{)}$
We derive the following:
$\|\mathrm{T}_{\lambda}W_{1}-\mathrm{T}_{\lambda}W_{2}\|\leq\frac{\gamma|1-\lambda|}{1-\lambda\gamma}\|W_{1}-W_{2}\|\textit{ where P = 1}$
It is established that $0\leq\lambda\leq 0$ constitutes a contraction. For $\lambda>1$, the following condition must be satisfied:
$\frac{\gamma(\lambda-1)}{1-\lambda\gamma}<1\Rightarrow\gamma\lambda-\gamma<1-\lambda\gamma$
$\Rightarrow 2\lambda\gamma<1+\gamma\Rightarrow\lambda<\frac{1+\gamma}{2\gamma}$ (34)
Consequently, $\mathrm{T}_{\lambda}$ constitutes a contraction if $0\leq\lambda<\frac{1+\gamma}{2\gamma}$. This directly indicates that for $\gamma<1$, $\gamma\lambda<1$.
Proof: Theorem 3
[…]
$Q^{\pi}(s,a)=\mathbf{E}\bigg{[}\gamma_{0}+\gamma r_{1}+...+\gamma^{k-1}r_{k-1}+\gamma^{k}Q^{\pi}(s_{k},a_{k})\bigg{]}$
Here the expectations are over a random trajectory under $\pi$; thus $\mathbf{E}[r_{\iota}](\iota\leq k\mathit{-}1)$ denotes the expected value of the $\iota$th reward received, while $\mathbf{E}[Q^{\pi}(s_{k},a_{k})]$ is the expected value of the true value function at the $k$th state-action reached. The phased TD(k) update sums the terms $\gamma^{\iota}\bigg{(}\frac{1}{n}\bigg{)}\sum_{i=1}^{n}r_{\iota}^{i}$, whose expectations are exactly the $\gamma^{\iota}\mathbf{E}[r_{\iota}]$ appearing above [9].
[…]
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{j=1}^{n}r_{i}^{(j)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\geq\epsilon\bigg{)}\leq 2e^{\bigg{(}\mathit{-}\frac{2n^{{}^{2}}\epsilon^{2}}{n2^{2}}\bigg{)}}$ (35)
If we assume n and the probability of exceeding an $\epsilon$ value is fixed to be atmost $\delta$, we can solve for the resulting value of $\epsilon$:
$2e^{\bigg{(}\mathit{-}\frac{2n^{{}^{2}}\epsilon^{2}}{n2^{2}}\bigg{)}}=\delta$
$e^{\bigg{(}\mathit{-}\frac{2n^{{}^{2}}\epsilon^{2}}{n2^{2}}\bigg{)}}=\frac{\delta}{2}$
Taking the Natural Logarithm
$ln\bigg{(}e^{\bigg{(}\mathit{-}\frac{2n^{{}^{2}}\epsilon^{2}}{n2^{2}}\bigg{)}}\bigg{)}=ln\bigg{(}\frac{\delta}{2}\bigg{)}$
$\mathit{-}\frac{2n^{{}^{2}}\epsilon^{2}}{n2^{2}}=ln\frac{\delta}{2}$
$\mathit{-}\frac{n\epsilon^{2}}{2}=ln\frac{\delta}{2}$
$\mathit{-}\frac{n\epsilon^{2}}{2}=ln\frac{\delta}{2}$
$\frac{n\epsilon^{2}}{2}=\mathit{-}ln\frac{\delta}{2}$
$\frac{n\epsilon^{2}}{2}=\mathit{-}\bigg{(}ln(\delta)-ln(2)\bigg{)}$
$\frac{n\epsilon^{2}}{2}=\mathit{-}\bigg{(}ln(\delta)+ln(2)\bigg{)}$
$\frac{n\epsilon^{2}}{2}=\bigg{(}ln(2)\mathit{-}ln(\delta)\bigg{)}$
$\frac{n\epsilon^{2}}{2}=ln\frac{2}{\delta}$
$n\epsilon^{2}=2log\frac{2}{\delta}$
$\epsilon^{2}=\frac{2log\frac{2}{\delta}}{n}$
$\epsilon=\sqrt{\frac{2log\frac{2}{\delta}}{n}}$ (36)
So by Hoeffding, if we have n samples, with probability at least $1\mathit{-}\delta$,
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{j=1}^{n}r_{i}^{(j)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\bigg{)}\leq\epsilon=\sqrt{\frac{2log\frac{2}{\delta}}{n}}$ (37)
Since we are dealing with k different state-action pairs, we use a union bound. To ensure the probability holds for all k hypotheses, adjust $\delta$ to $\delta/k$:
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{j=1}^{n}r_{i}^{(j)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\bigg{)}\leq\epsilon=\sqrt{\frac{2log\frac{2k}{\delta}}{n}}$
We can now assume that all $\mathbb{E}[r_{i}]$ terms are estimated to at least $\epsilon$ accuracy. Substituting this back into the definition of the k-step TD update we get
$\hat{Q}_{t+1}(s,a)\mathit{-}Q(s,a)=\frac{1}{n}\sum_{j=1}^{n}\bigg{(}r_{0}+\gamma r_{1}+...+\gamma^{k-1}r_{k-1}+\gamma^{k}Q_{t}(s_{k},a_{k})\bigg{)}\mathit{-}Q(s,a)$
$=\sum_{i=0}^{k-1}\gamma^{i}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}r_{i}^{(j)}-\mathbb{E}[r_{i}]\bigg{)}+\gamma^{k}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}Q_{t}(s_{k}^{i},a_{k}^{i})-\mathbb{E}[Q(s_{k},a_{k})]\bigg{)}$ (38)
where in the second line we re-expressed the value in terms of a sum of k rewards. We now upper bounded the difference from $E[r_{i}]$ by $\epsilon$ to get
$\hat{Q}_{t+1}(s,a)\mathit{-}Q(s,a)\leq\sum_{i=0}^{k-1}\gamma^{l}\epsilon+\gamma^{k}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}Q_{t}(s_{k}^{i},a_{k}^{i})-\mathbb{E}[Q(s_{k},a_{k})]\bigg{)}$
The variance term arises from the deviation of the empirical average of rewards from the true expected reward:
$\epsilon\bigg{(}\frac{1-\gamma^{k}}{1-\gamma}\bigg{)}$
$\leq\epsilon\bigg{(}\frac{1-\gamma^{k}}{1-\gamma}\bigg{)}+\gamma^{k}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}Q_{t}(s_{k}^{i},a_{k}^{i})-\mathbb{E}[Q(s_{k},a_{k})]\bigg{)}$ (39)
The bias term is propagated through bootstrapping:
$\gamma^{k}\Delta^{\hat{Q}_{\gamma}}_{t-1}$
and then the second term is bounded by $\Delta^{\hat{Q}_{\gamma}}_{t-1}$ by assumption. Hence, combining these terms, we get the overall bound for SARSA:
$\Delta^{\hat{Q}_{\gamma}}_{t}\leq\epsilon\bigg{(}\frac{1-\gamma^{k}}{1-\gamma}\bigg{)}+\gamma^{k}\Delta^{\hat{Q}_{\gamma}}_{t-1}$ (40)
Proof: Theorem 4
Phased $TD(\Delta)$ update rules for $z\geq 1$:
$\hat{W}_{z,t}(s,a)=\frac{1}{n}\sum_{j-1}^{n}\bigg{(}\sum_{i=1}^{k_{z}-1}(\gamma_{z}^{i}-\gamma_{z-1}^{i})r_{i}^{(j)}+(\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}})\hat{Q}_{\gamma_{z-1}}(s_{k}^{(j)},a_{k}^{(j)})+\gamma_{z}^{k_{z}}\hat{W}_{z}(s_{k}^{(j)},a_{k}^{(j)})\bigg{)}$ (41)
We know that according to the multi-step update rule 14 for $z\geq 1$:
$W_{z}(s_{t},a_{t})=\mathbb{E}\bigg{[}\sum_{i=1}^{k_{z}-1}\bigg{(}\gamma_{z}^{i}-\gamma_{z-1}^{i}\bigg{)}r_{i}+\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}Q_{\gamma_{z-1}}(s_{k},a_{k})+\gamma_{z}^{k_{z}}W_{z}(s_{k},a_{k})\bigg{]}$ (42)
Then, subtracting the two expressions gives for $z\geq 1$:
$\hat{W}_{z,t}(s,a)-W_{z}(s_{t},a_{t})=\sum_{i=1}^{k_{z}-1}\bigg{(}\gamma_{z}^{i}-\gamma_{z-1}^{i}\bigg{)}\bigg{(}\frac{1}{n}\sum_{j-1}^{n}r_{i}^{(j)}-\mathbb{E}[r_{i}]\bigg{)}+\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\bigg{(}\sum_{u=0}^{z-1}\hat{W}_{u}(s_{k}^{(j)},a_{k}^{(j)})-\mathbb{E}[W_{z}(s_{k},a_{k})]\bigg{)}$
$+\gamma_{z}^{k_{z}}\bigg{(}W_{z}(s_{k}^{(j)},a_{k}^{(j)})-\mathbb{E}[W_{z}(s_{k},a_{k})]\bigg{)}$ (43)
Assume that $k_{0}\leq k_{1}\leq...k_{z}=k$, the W estimates share at most $k_{z}=k$ reward terms $\frac{1}{n}\sum_{j-1}^{n}r_{i}^{(j)}$, Using Hoeffding inequality and union bound, we obtain that with probability $1-\delta$, each k empirical average reward $\frac{1}{n}\sum_{j-1}^{n}r_{i}^{(j)}$ deviates from the true expected reward $\mathbb{E}[r_{i}]$ by at most $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$. Hence, with probability $1-\delta$, $\forall z\geq 1$, we have:
$\Delta_{t}^{z}\leq\epsilon\sum_{i=1}^{k_{z}-1}\bigg{(}\gamma_{z}^{i}-\gamma_{z-1}^{i}\bigg{)}+\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\sum_{u=0}^{z-1}\Delta_{t-1}^{u}+\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}$
$=\epsilon\bigg{(}\frac{1-\gamma_{z}^{k_{z}}}{1-\gamma_{z}}-\frac{1-\gamma_{z-1}^{k_{z}}}{1-\gamma_{z-1}}\bigg{)}+\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\sum_{u=0}^{z-1}\Delta_{t-1}^{u}+\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}$ (44)
and $\Delta_{t}^{0}\leq\epsilon\frac{1-\gamma_{0}^{k_{0}}}{1-\gamma_{0}}+\gamma_{0}^{k_{0}}\Delta_{t-1}^{0}$
Summing the two previous inequalities gives:
$\sum_{z=0}^{Z}\Delta_{t}^{z}\leq\epsilon\frac{1-\gamma_{0}^{k_{0}}}{1-\gamma_{0}}+\epsilon\sum_{z=1}^{Z}\bigg{(}\frac{1-\gamma_{z}^{k_{z}}}{1-\gamma_{z}}-\frac{1-\gamma_{z-1}^{k_{z}}}{1-\gamma_{z-1}}\bigg{)}+\sum_{z=1}^{Z}\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\sum_{u=0}^{z-1}\Delta_{t-1}^{u}+\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}$
$=\underbrace{\epsilon\frac{1-\gamma_{Z}^{k_{Z}}}{1-\gamma_{Z}}+\epsilon\sum_{z=0}^{Z-1}\frac{\gamma_{z}^{k_{z+1}}-\gamma_{z}^{k_{z}}}{1-\gamma_{z}}}_{\text{(*)variance term}}+\underbrace{\sum_{z=1}^{Z}\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\sum_{u=0}^{z-1}\Delta_{t-1}^{u}+\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}}_{\text{(**)bias term}}$ (45)
Let’s focus now further on the bias term (**)
$\sum_{z=1}^{Z}\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\sum_{u=0}^{z-1}\Delta_{t-1}^{u}+\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}=\sum_{u=0}^{Z-1}\sum_{z=u+1}^{Z}\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\Delta_{t-1}^{u}+\sum_{z=1}^{Z}\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}$
$=\sum_{u=0}^{Z-1}\Delta_{t-1}^{u}\bigg{(}\sum_{z=u+1}^{Z}\gamma_{z}^{k_{z}}-\sum_{z=u}^{Z-1}\gamma_{z}^{k_{z+1}}\bigg{)}+\sum_{z=1}^{Z}\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}$
$=\sum_{u=0}^{Z-1}\Delta_{t-1}^{u}\bigg{(}\sum_{z=u+1}^{Z-1}(\gamma_{z}^{k_{z}}-\gamma_{z}^{k_{z+1}})+(\gamma_{Z}^{k_{z}}-\gamma_{u}^{k_{u+1}})\bigg{)}+\sum_{z=1}^{Z}\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}$
$=\sum_{u=0}^{Z-1}\sum_{z=u+1}^{Z-1}(\gamma_{z}^{k_{z}}-\gamma_{z}^{k_{z+1}})\Delta_{t-1}^{u}+\gamma_{Z}^{k_{z}}\sum_{z=0}^{Z}\Delta_{t-1}^{z}+\sum_{z=0}^{Z-1}(\gamma_{z}^{k_{z}}-\gamma_{z}^{k_{z+1}})\Delta_{t-1}^{z}$
$=\sum_{u=0}^{Z-1}\sum_{z=u}^{Z-1}(\gamma_{z}^{k_{z}}-\gamma_{z}^{k_{z+1}})\Delta_{t-1}^{u}+\gamma_{Z}^{k_{z}}\sum_{z=0}^{Z}\Delta_{t-1}^{z}$
$=\sum_{z=0}^{Z-1}(\gamma_{z}^{k_{z}}-\gamma_{z}^{k_{z+1}})\sum_{u=0}^{z}\Delta_{t-1}^{u}+\gamma_{Z}^{k_{z}}\sum_{z=0}^{Z}\Delta_{t-1}^{z}$ (46)
Finally, we obtain:
$\sum_{z=0}^{Z}\Delta_{t}^{z}\leq\underbrace{\epsilon\bigg{(}\frac{1-\gamma^{k}}{1-\gamma}\bigg{)}+\epsilon\bigg{(}\sum_{z=0}^{Z-1}\frac{\gamma_{z}^{k_{z+1}}-\gamma^{k_{z}}_{z}}{1-\gamma_{z}}\bigg{)}}_{\text{variance reduction}}+\underbrace{\sum_{z=0}^{Z-1}\bigg{(}\gamma^{k_{z}}_{z}-\gamma^{k_{z+1}}_{z}\bigg{)}\sum_{u=0}^{z}\Delta_{t-1}^{u}+\gamma^{k}\sum_{z=0}^{Z}\Delta_{t-1}^{z}}_{\text{bias introduction}}$ (47)

### v2

@@
-Segmenting Action-Value Functions Over Time-Scales in SARSA using TD($\Delta$)
+Segmenting Action-Value Functions Over Time-Scales in SARSA via TD($\Delta$)
 Abstract
@@
 Conventional SARSA algorithms, however, have difficulties in balancing bias and variation due to the reliance on a singular, fixed discount factor.
-This study expands the temporal difference decomposition approach, TD($\triangle$), to the SARSA algorithm.
+This study expands the temporal difference decomposition approach, TD($\Delta$), to the SARSA algorithm, which we designate as SARSA($\Delta$).
 SARSA, a widely utilised on-policy RL method, enhances action-value functions via temporal difference updates.
-TD($\triangle$) facilitates learning over several time-scales by breaking the action-value function into components associated with distinct discount factors.
+TD($\Delta$) facilitates learning over several time-scales by breaking the action-value function into components associated with distinct discount factors.
 This decomposition improves learning efficiency and stability, particularly in problems necessitating long-horizon optimization.
-We illustrate that our methodology mitigates bias in SARSA’s updates while facilitating accelerated convergence in contexts characterized by dense rewards.
-Experimental findings across many benchmark tasks indicate that the proposed SARSA($\triangle$) surpasses conventional TD learning methods in both tabular and deep RL contexts.
+We illustrate that our methodology mitigates bias in SARSA’s updates while facilitating accelerated convergence in both deterministic and stochastic environments.
+Experimental findings across many benchmark tasks indicate that the proposed SARSA($\Delta$) surpasses conventional TD learning methods in both tabular and deep RL environments.
 keywords:
@@
 Temporal difference (TD) learning methods, like Q-learning and SARSA, have demonstrated efficacy across several tasks by allowing agents to estimate action-value functions that forecast anticipated future rewards.
-Traditionally, these strategies utilise a discount factor $10\leq\gamma<$ 1, with values approaching $\gamma=0$ to emphasise short-term rewards over long-term ones, therefore diminishing the planning horizon and enhancing learning stability and efficiency[1].
-Discount factors $\gamma<1$ frequently yield enhanced results during the early phases of learning, as demonstrated by Prokhorov and Wunsch[2].
+Traditionally, these strategies utilise a discount factor $10\leq\eta<$ 1, with values approaching $\eta=0$ to emphasise short-term rewards over long-term ones, therefore diminishing the planning horizon and enhancing learning stability and efficiency[1].
+Discount factors $\eta<1$ frequently yield enhanced results during the early phases of learning, as demonstrated by Prokhorov and Wunsch[2].
 In specific contexts, such as long-horizon tasks[3, 4] where long-term planning is essential—like navigation or decision-making tasks with deferred rewards—this method can create bias, constraining the agent’s capacity to acquire optimal long-horizon policies.
 For instance: SARSA[1] (State-Action-Reward-State-Action) is an on-policy RL algorithm that modifies an action-value function Q(s,a) according to the state-action transitions encountered by the agent.
-The conventional SARSA algorithm performs well in contexts with shorter time horizons; however, in long-horizon tasks, the singular discount factor $\gamma$ employed in SARSA results in a compromise between bias and variance.
-Small values of $\gamma$ yield myopic policies, whereas larger values induce substantial variance in the estimated action-value functions, rendering the learning process unstable and inefficient.
+The conventional SARSA algorithm performs well in contexts with shorter time horizons; however, in long-horizon tasks, the singular discount factor $\eta$ employed in SARSA results in a compromise between bias and variance.
+Small values of $\eta$ yield myopic policies, whereas larger values induce substantial variance in the estimated action-value functions, rendering the learning process unstable and inefficient.
 In response, other solutions have been suggested, including[4, 2, 5, 6, 7, 8].
-Recent research in TD learning has investigated the segmentation of value functions over many time-scales, enabling agents to concurrently enhance value estimates at numerous horizons[8].
-The TD($\Delta$) architecture enhances scalability and performance in situations with long-term dependencies by decomposing value functions into components that encompass various discount factors, referred to as TD($\Delta$).
-The TD($\Delta$) framework’s adaptability demonstrates significant potential in mitigating the bias-variance trade-offs inherent in TD learning.
-We suggest SARSA($\Delta$), an extension of the TD($\Delta$) method applied to the SARSA algorithm, building on this concept.
-TD($\Delta$) emphasises the actor-critic paradigm and conventional TD learning, whereas SARSA($\Delta$) disaggregates the state-action value function into several partial estimators across various discount factors, termed delta estimators.
-These estimators approximate the difference $W_{z}(s,a)=Q_{\gamma_{z}}(s,a)\mathit{-}Q_{\gamma_{z-1}}(s,a)$ between action-value functions, facilitating efficient learning in scenarios where actions affect long-horizon outcomes.
-SARSA($\Delta$) optimises a sequence of delta estimators, each associated with a distinct discount factor, similar to TD($\Delta$).
-The segregation enables the autonomous learning of each component, with lower discount factors achieving faster convergence, while bigger discount factors enhance this foundation for improved long-horizon planning.
-This study introduces a multi-step SARSA($\Delta$) update rule that disaggregates state-action value functions, facilitating accelerated convergence and enhanced policy efficacy, especially in intricate environments characterised by dense rewards.
+Recent research in TD learning has investigated the segmentation of value functions across multiple time scales, enabling agents to enhance value estimates concurrently at different horizons[8].
+The TD($\Delta$) architecture improves scalability and performance in scenarios involving long-term dependencies by breaking down value functions into components that incorporate different discount factors, known as TD($\Delta$).
+The adaptability of the TD($\Delta$) framework shows considerable promise in addressing the bias-variance trade-offs that are characteristic of TD learning.
+We propose SARSA($\Delta$), an enhancement of the TD($\Delta$) approach integrated into the SARSA algorithm, based on this principle.
+TD($\Delta$) emphasizes the actor-critic framework and conventional TD learning, whereas SARSA($\Delta$) divides the state-action value function into multiple partial estimators corresponding to various discount factors, known as delta estimators.
+These estimators approximate the difference $D_{m}(s,a)=Q_{\eta_{m}}(s,a)-Q_{\eta_{m-1}}(s,a)$ between action-value functions, thereby improving learning efficiency in scenarios where actions affect long-horizon outcomes.
+Similar to TD($\Delta$), SARSA($\Delta$) seeks to optimize a set of delta estimators, where each estimator is associated with a distinct discount factor.
+With smaller discount factors, convergence happens faster because to segmentation, while greater discount factors strengthen this foundation for better long-horizon planning.
+This study introduces a multi-step SARSA($\Delta$) update rule that disaggregates action-value functions, facilitating accelerated convergence and enhanced policy efficacy, especially in intricate deterministic and stochastic environments.
 Furthermore, we illustrate that SARSA($\Delta$) may seamlessly adapt to any multi-step reinforcement learning techniques, including n-step SARSA and eligibility traces.
@@
 This addon provides a more detailed and scalable method for action-value function learning in intricate environments.
-The primary contributions of this paper are as follows: (i) We expand TD($\Delta$) to SARSA($\Delta$) by partitioning the action-value function Q(s,a) into delta components over several time scales.
+The primary contributions of this paper are as follows: (i) We expand TD($\Delta$) to SARSA($\Delta$) by partitioning the action-value function Q(s,a) into delta components over several time-scales.
 (ii) The suggested SARSA($\Delta$) algorithm enhances learning efficiency and stability by independently learning components for various discount factors and integrating them to represent the complete action-value function.
 (iii) Both theoretical and empirical evaluations have shown that this multi-scale breakdown has a number of benefits, particularly in environments where there are long-horizon rewards.
-The subsequent sections of the paper are structured as follows: Section 2 contains a discussion of related works.
-Section 3 presents the necessary background information.
-Sections 4 and 5 illustrate the principles of SARSA($\Delta$) and theoretical analysis, respectively.
-Section 6 presents a comprehensive overview of the experiments conducted.
-Ultimately, we provide a conclusion in section 7.
+The ensuing portions of the paper are organized as follows: Section 2 presents a discourse on pertinent literature.
+Section 3 provides the essential background information.
+The principles of SARSA($\Delta$) and theoretical analysis are shown in Sections 4 and 5, respectively.
+Section 6 provides a detailed summary of the experiments performed.
+Ultimately, we present a conclusion in section 7.
 2 Related Work
@@
 Researchers[8] have investigated techniques to separate value functions across various time scales by analysing disparities among value functions with lower discount factors, hence enhancing scalability and performance.
-The study introduces $\gamma$-nets[10], a technique for generalising value function estimate across timescales in reinforcement learning.
+The study introduces $\gamma$-nets[10], a technique for generalising value function estimate across timescales in RL.
 The authors of the paper[11] present an approach for integrating hyperbolic discounting into RL.
@@
 The authors present an agent designed to approximate hyperbolic discount functions by utilizing temporal-difference learning methods.
-Furthermore, the study reveals that learning value functions across multiple time horizons improves performance, especially when implemented as an auxiliary task in value-based reinforcement learning algorithms such as Rainbow.
+Furthermore, the study reveals that learning value functions across multiple time horizons improves performance, especially when implemented as an auxiliary task in value-based RL algorithms such as Rainbow.
 This method demonstrates potential in contexts characterized by uncertainty and risk.
@@
 It is noteworthy that none of the previously mentioned works employ short-term estimates for training long-term action-value functions.
-Our method can also be utilised as a generalised action-value function, and its capacity to query smaller timescales represents an additional advantage.
+Our method can also be utilised as a generalised action-value function, and its capacity to query smaller time-scales represents an additional advantage.
 Performance improvements primarily arise from the separation of action-value functions across varying time scales through the use of TD($\Delta$) in SARSA.
@@
 The authors[6] of this study introduce meta-gradient methods for the adaptive selection of discount factors in RL, emphasising the necessity of balancing short- and long-term rewards, which corresponds with the rationale for dissecting action-value functions across several discount factors in SARSA($\Delta$).
-Lastly, another extensive category of study pertinent to our research indirectly encompasses hierarchical RL and the bias-variance tradeoff.
+Lastly, another extensive category of study pertinent to our research indirectly encompasses hierarchical RL and the bias-variance trade-off.
 These studies[14, 15, 16, 17, 18, 19, 20] present the concept of deconstructing value functions into subtasks by hierarchical RL, analogous to the manner in which TD($\Delta$) disaggregates value functions across temporal scales.
 This substantiates the assertion that value function decomposition can enhance learning efficacy and stability.
-This paper[9] functions as a theoretical foundation for comprehending the bias-variance tradeoff that arises when decomposing Q-values across various time scales, and it also establishes constraints on the bias and variance of TD updates.
+This paper[9] functions as a theoretical foundation for comprehending the bias-variance trade-off that arises when decomposing Q-values across various time-scales, and it also establishes constraints on the bias and variance of TD updates.
 This study[21] presents empirical findings demonstrating the efficacy of deep RL techniques utilizing value function approximation in Atari-like environments.
@@
 Examine a completely observable Markov Decision Process (MDP) [22], characterized by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ signifies the action space, and $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}\rightarrow[0,1]$ represents the transition probabilities that associate state-action pairings with distributions across subsequent states, whereas $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denotes the reward function.
-At each timestep $t$, the agent occupies state $s_{t}$, chooses an action $a_{t}$, obtains a reward $r_{t}=r(s_{t},a_{t})$, and transitions to the subsequent state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$.
-In a conventional Markov Decision Process (MDP) framework, an agent seeks to maximize the discounted return defined as $Q_{\gamma}^{\pi}(s,a)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t+1}\mid s_{t}=s,a_{t}=a\right]$, where $\gamma$ denotes the discount factor and $\pi:\mathcal{S}\rightarrow\mathcal{A}\rightarrow[0,1]$ represents the policy adhered to by the agent.
-The action-value function $Q_{\gamma}^{\pi}(s,a)$ is determined as the fixed point of the Bellman operator $\mathcal{T}Q=r^{\pi}+\gamma\mathcal{P}^{\pi}Q$, where $r^{\pi}$ denotes the expected immediate reward and $\mathcal{P}^{\pi}$ represents the transition probability operator associated with the policy $\pi$.
+At each timestep $n$, the agent occupies state $s_{n}$, chooses an action $a_{n}$, obtains a reward $r_{n}=r(s_{n},a_{n})$, and transitions to the subsequent state $s_{n+1}\sim\mathcal{P}(\cdot|s_{n},a_{n})$.
+In a conventional Markov Decision Process (MDP) framework, an agent seeks to maximize the discounted return defined as $Q_{\eta}^{\pi}(s,a)=\mathbb{E}\left[\sum_{n=0}^{\infty}\eta^{n}r_{n+1}\mid s_{n}=s,a_{n}=a\right]$, where $\eta$ denotes the discount factor and $\pi:\mathcal{S}\rightarrow\mathcal{A}\rightarrow[0,1]$ represents the policy adhered to by the agent.
+The action-value function $Q_{\eta}^{\pi}(s,a)$ is determined as the fixed point of the Bellman operator $\mathcal{T}Q=r^{\pi}+\eta\mathcal{P}^{\pi}Q$, where $r^{\pi}$ denotes the expected immediate reward and $\mathcal{P}^{\pi}$ represents the transition probability operator associated with the policy $\pi$.
 For convenience, we remove the superscript $\pi$ for the rest of the paper.
-Using temporal difference (TD)[23] learning, the action-value estimate $\hat{Q}_{\gamma}$ can approximate the true action-value function $Q_{\gamma}$.
-The one-step TD error $\delta^{\gamma}_{t}=r_{t+1}+\gamma\hat{Q}_{\gamma}(s_{t+1},a_{t+1})\mathit{-}\hat{Q}_{\gamma}(s_{t},a_{t})$ is used to update the action-value function given a transition $(s_{t},a_{t},r_{t},s_{t+1})$.
-SARSA is an on-policy RL approach that acquires the action-value function Q(s,a) for a specified policy $\pi$.
-The action-value function denotes the anticipated cumulative reward originating from state s, executing action a, and following policy $\pi$.
-Subsequently, the SARSA update rule is articulated as follows:
-$Q(s_{t},a_{t})\leftarrow Q(s_{t},a_{t})+\alpha\left[r_{t}+\gamma Q(s_{t+1},a_{t+1})-Q(s_{t},a_{t})\right]$ (1)
-where $\alpha$ denotes the learning rate, $\gamma$ represents the discount factor, and $r_{t}$ signifies the immediate reward.
-Long-horizon tasks present challenges for SARSA since the choice of $\gamma$ can compromise learning efficiency with appropriate long-term reward maximizing.
+Using temporal difference (TD)[23] learning, the action-value estimate $\hat{Q}_{\eta}$ can approximate the true action-value function $Q_{\eta}$.
+The one-step TD error $\delta^{\eta}_{n}=r_{n+1}+\eta\hat{Q}_{\eta}(s_{n+1},a_{n+1})\mathit{-}\hat{Q}_{\eta}(s_{n},a_{n})$ is used to update the action-value function given a transition $(s_{n},a_{n},r_{n},s_{n+1})$.
+An on-policy RL technique called SARSA learns the action-value function Q(s,a) for a given policy $\pi$.
+The action-value function represents the expected total reward that comes from state s, taking action a, and following policy $\pi$.
+Then the SARSA update rule is clarified as follows:
+$Q(s_{n},a_{n})\leftarrow Q(s_{n},a_{n})+\alpha\left[r_{n}+\eta Q(s_{n+1},a_{n+1})-Q(s_{n},a_{n})\right]$ (1)
+Here, $\alpha$ is the learning rate, $\eta$ is the discount factor, and $r_{n}$ is the immediate reward.
+Long-horizon tasks present challenges for SARSA since the choice of $\eta$ can compromise learning efficiency with appropriate long-term reward maximizing.
 On the other hand, for a complete trajectory, we can employ the discounted sum of one-step TD errors, generally known as the $\lambda$-return[23] or, equivalently, the Generalized Advantage Estimator (GAE)[24].
-The GAE enhances advantage estimates by reconciling the trade-off between variance and bias via parameters $\lambda$ and $\gamma$.
+The GAE enhances advantage estimates by reconciling the trade-off between variance and bias via parameters $\lambda$ and $\eta$.
 The equation for the Generalized Advantage Estimator is generally expressed as:
-$A(s_{t},a_{t})=\sum_{k=0}^{\infty}(\lambda\gamma)^{k}\delta_{t+k}^{\gamma}$ (2)
-Where $\delta_{t+k}$ is the TD error at time t, computed as follows: $\delta_{t+k}=r_{t+k}+\gamma Q(s_{t+k+1},a_{t+k+1})-Q(s_{t+k},a_{t+k})$
+$A(s_{n},a_{n})=\sum_{k=0}^{\infty}(\lambda\eta)^{k}\delta_{t+k}^{\eta}$ (2)
+Where $\delta_{n+k}$ is the TD error at time n, computed as follows: $\delta_{n+k}=r_{n+k}+\eta Q(s_{n+k+1},a_{n+k+1})-Q(s_{n+k},a_{n+k})$
 Loss function for Q-Value estimation utilizing GAE.
-The loss function $\mathcal{L}(\theta)$ for approximating the Q-value function is defined as the mean squared error between the current Q-value estimate $Q(s_{t},a_{t};\theta)$ and the target Q-value adjusted by the advantage estimator.
-Thus, it is possible to concisely write the loss function using $A(s_{t},a_{t})$ and $Q(s_{t},a_{t})$ as follows:
-$\mathcal{L}(\theta)=\mathbb{E}\bigg{[}\bigg{(}Q(s_{t},a_{t};\theta)-\bigg{(}Q(s_{t},a_{t})+A(s_{t},a_{t})\bigg{)}\bigg{)}^{2}\bigg{]}$ (3)
+The loss function $\mathcal{L}(\theta)$ for approximating the Q-value function is defined as the mean squared error between the current Q-value estimate $Q(s_{n},a_{n};\theta)$ and the target Q-value adjusted by the advantage estimator.
+Thus, it is possible to concisely write the loss function using $A(s_{n},a_{n})$ and $Q(s_{n},a_{n})$ as follows:
+$\mathcal{L}(\theta)=\mathbb{E}\bigg{[}\bigg{(}Q(s_{n},a_{n};\theta)-\bigg{(}Q(s_{n},a_{n})+A(s_{n},a_{n})\bigg{)}\bigg{)}^{2}\bigg{]}$ (3)
 Despite SARSA being on-policy and lacking an explicit policy update, policy selection in SARSA can still be influenced by Eq.4.
-For actor-critic architectures[25, 26, 27], the action-value function is updated according to Eq.3, and a stochastic parameterized policy (actor, $\pi_{\omega}(a|s)$) is learned from this value estimator through the advantage function, with the loss being.
-$\mathcal{L}(\omega)=\mathbb{E}\bigg{[}-log\pi(a|s;\omega)A(s,a)\bigg{]}$ (4)
-In on-policy methods like SARSA, the likelihood ratio $\rho$ between the current policy and a prior policy is typically not employed.
+For actor-critic architectures[25, 26, 27], the action-value function is updated according to Eq.3, and a stochastic parameterized policy (actor, $\pi_{\nu}(a|s)$) is learned from this value estimator through the advantage function, with the loss being.
+$\mathcal{L}(\nu)=\mathbb{E}\bigg{[}-log\pi(a|s;\nu)A(s,a)\bigg{]}$ (4)
+Proximal Policy Optimization (PPO) [28], an enhancement of actor-critic methodologies, confines policy updates to a designated optimization area, referred to as a trust region, utilizing a clipping objective that evaluates the current parameters, $\nu$, against the prior parameters, $\nu_{old}$:
+$\mathcal{L}(\nu)=\mathbb{E}\bigg{[}min\bigg{(}\rho(\nu)A(s,a),\psi(\nu)A(s,a)\bigg{)}\bigg{]}$ (5)
+In on-policy methods like SARSA, the likelihood ratio $\rho(\nu)$ between the current policy and a prior policy is typically employed.
 SARSA evaluates actions in accordance with the current policy, concurrently updating the action-value function as actions are performed.
-Thus, Proximal Policy Optimization (PPO)[28] does not implement a clipping objective function in SARSA as used in this study[8].
-Eq.4 is utilized as the objective function.
+However, when utilized within an actor-critic framework, the SARSA-learning agent may adopt a similar policy ratio:
+$\rho(\nu)=\frac{Q(s,a;\nu)}{Q(s,a;\nu_{old})}$ (6)
+where $\psi(\nu)=clip(\rho,1-\epsilon,1+\epsilon)$ represents the clipped likelihood ratio, and $\epsilon<1$ is a negligible parameter employed to constrain the update.
+The loss function for the policy update can be articulated as follows:
+$\mathcal{L}(\nu)=\mathbb{E}\bigg{[}min\bigg{(}\rho(\nu)A(s,a),clip(\rho(\nu),1-\epsilon,1+\epsilon)A(s,a)\bigg{)}\bigg{]}$ (7)
 4 Methodology
-The TD($\Delta$) framework enhances conventional temporal difference learning by partitioning action-value functions across several discount factors $\gamma_{0},\gamma_{1},...,\gamma_{z}$.
-This decomposition enables the action-value function to be acquired as a summation of delta estimators, each representing the disparity between action-value functions at varying discount factors.
-The principal advantage of this method is the capacity to regulate variance and bias in the learning process by concentrating on smaller, more manageable time-scales.
+The TD($\Delta$) paradigm boosts regular TD learning by splitting action-value functions over several discount factors $\eta_{0},\eta_{1},...,\eta_{m}$.
+This decomposition facilitates the derivation of the action-value function as a sum of delta estimators, where each estimator indicates the distinction between action-value functions at various discount factors.
+Concentrating on smaller, more manageable time scales facilitates the regulation of learning variance and bias, which constitutes the primary benefit of this approach.
 4.1 TD($\Delta$) Framework
-To calculate $W_{z}$ for SARSA as outlined in the study, it is essential to comprehend how delta estimators ($W_{z}$) approximate the discrepancies between action-value functions at consecutive discount factors.
-Let us delineate this approach for SARSA in a stepwise manner.
-Delta Estimators ($W_{z}$) and Action-Value Functions ($Q(s,a)$): Delta estimators, $W_{z}$, represent the difference between value functions with successive discount factors:
-$W_{z}=Q_{\gamma_{z}}\mathit{-}Q_{\gamma_{z-1}}$ (5)
-In this context, $Q_{\gamma_{z}}$ denotes the action-value function with discount factor $\gamma_{z}$, but $Q_{\gamma_{z-1}}$ signifies the action-value function with the prior discount factor, $\gamma_{z-1}$.
+To compute $D_{m}$ for SARSA as outlined in the study, it is essential to comprehend the mechanism by which delta estimators ($D_{m}$) approximate the discrepancies between action-value functions across successive discount factors.
+This approach for SARSA will be outlined in a stepwise manner.
+Delta Estimators ($D_{m}$) and Action-Value Functions ($Q(s,a)$): Delta estimators, denoted as $D_{m}$, quantify the variation between action-value functions associated with consecutive discount factors:
+$D_{m}=Q_{\eta_{m}}\mathit{-}Q_{\eta_{m-1}}$ (8)
+In this context, $Q_{\eta_{m}}$ denotes the action-value function with discount factor $\eta_{m}$, but $Q_{\eta_{m-1}}$ signifies the action-value function with the prior discount factor, $\eta_{m-1}$.
 4.2 Single-Step TD ( SARSA($\Delta$) )
-This section begins with a description of the delta estimators ($W_{z}$) for the action-value functions $Q$ utilized in SARSA.
-The extension of TD($\Delta$) to SARSA is denoted as SARSA($\Delta$).
-SARSA($\Delta$) aims to obtain the action-value function $Q(s,a)$, representing the expected return of adhering to policy $\pi$ from state $s$, performing action $a$, and subsequently continuing to follow $\pi$.
-The aim is to improve the overall reward through the optimization of the action-value function across different time-scales.
-The core principle in SARSA($\Delta$) involves dividing the action-value function $Q(s,a)$ into a series of delta components $W_{z}(s,a)$, with each component associated with a specific discount factor $\gamma_{z}$.
-The interrelation among these delta components (i.e., delta function) is articulated as follows:
-$W_{z}(s,a):=Q_{\gamma_{z}}(s,a)\mathit{-}Q_{\gamma_{z-1}}(s,a)$ (6)
-where $\gamma_{0},\gamma_{1},...,\gamma_{z}$ represent the discount factors across various time-scales, and define $W_{0}(s,a):=Q_{\gamma_{0}}(s,a)$.
-The action-value function $Q_{\gamma_{z}}(s,a)$ is defined as the cumulative sum of all W-components up to z:
-$Q_{\gamma_{z}}(s,a)=\sum_{j=0}^{z}W_{j}(s,a)$ (7)
+This section initiates with an overview of the delta estimators ($D_{m}$) pertaining to the action-value functions $Q$ employed in SARSA.
+SARSA is extended to TD($\Delta$), which is referred to as SARSA($\Delta$).
+SARSA($\Delta$) seeks to derive the action-value function $Q(s,a)$, which indicates the anticipated return of following policy $\pi$ from state $s$, executing action $a$, and then persistently adhering to $\pi$ thereafter.
+The objective is to enhance the total reward by optimizing the action-value function over various time-scales.
+The fundamental concept in SARSA($\Delta$) is the segmentation of the action-value function $Q(s,a)$ into multiple delta components $D_{m}(s,a)$, where each component is linked to a distinct discount factor $\eta_{m}$.
+The relationship between these delta components (i.e., delta function) is expressed in the following manner:
+$D_{m}(s,a):=Q_{\eta_{m}}(s,a)\mathit{-}Q_{\eta_{m-1}}(s,a)$ (9)
+where $\eta_{0},\eta_{1},...,\eta_{m}$ represent the discount factors across various time-scales, and define $D_{0}(s,a):=Q_{\eta_{0}}(s,a)$.
+The action-value function $Q_{\eta_{m}}(s,a)$ is defined as the cumulative sum of all D-components up to m:
+$Q_{\eta_{m}}(s,a)=\sum_{x=0}^{m}D_{x}(s,a)$ (10)
 The update rule in standard SARSA is determined by the state-action pair in accordance with the policy.
 The action-value function is updated through the TD error, with the action in the subsequent state selected in accordance with the policy:
-$Q(s_{t},a_{t})\leftarrow Q(s_{t},a_{t})+\alpha\left[r_{t}+\gamma Q(s_{t+1},a_{t+1})-Q(s_{t},a_{t})\right]$ (8)
-Here, $r_{t}$ denotes the reward obtained from executing action $a_{t}$ in state $s_{t}$, while $(s_{t+1},a_{t+1})$ represents the subsequent state-action pair.
-The SARSA update rule is applied individually to each delta component $W_{z}(s,a)$.
-The update for each component employs the same framework as the traditional SARSA method, yet it is adjusted to incorporate the delta function at each specific time-scale.
-We can express the update Eq.8 for single-step TD SARSA ($\Delta$) with multiple time-scales by applying the concept of separating the update into multiple discount factors[8] as follows:
-$W_{z}(s_{t},a_{t})=\mathbb{E}\left[(\gamma_{z}-\gamma_{z-1})Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})+\gamma_{z}W_{z}(s_{t+1},a_{t+1})\right]$ (9)
-where $Q_{\gamma_{z}}(s_{t},a_{t})$ represents the action-value function with a discount factor $\gamma_{z}$.
-$W_{z}(s_{t},a_{t})$ denotes the delta function for SARSA ($\Delta$), while $\gamma_{z}$ represents the discount factor for time-scale z, while $Q_{\gamma_{z}}(s_{t+1},a_{t+1})$ is the SARSA update for action-value associated with the subsequent state-action pair.
+$Q(s_{n},a_{n})\leftarrow Q(s_{n},a_{n})+\alpha\left[r_{n}+\eta Q(s_{n+1},a_{n+1})-Q(s_{n},a_{n})\right]$ (11)
+In this context, $r_{n}$ signifies the reward acquired from performing action $a_{n}$ in state $s_{n}$, whereas $(s_{n+1},a_{n+1})$ denotes the ensuing state-action pair.
+The SARSA update rule is implemented independently to each delta component $D_{n}(s,a)$.
+The modifications for each component utilize the same architecture as the conventional SARSA approach, but it is tailored to include the delta function at each distinct time-scale.
+The update in Eq.11 for single-step TD SARSA ($\Delta$) can be articulated with numerous time-scales by employing the principle of splitting the update into several discount factors[8] as demonstrated below:
+$D_{m}(s_{n},a_{n})=\mathbb{E}\left[(\eta_{m}-\eta_{m-1})Q_{\eta_{m-1}}(s_{n+1},a_{n+1})+\eta_{m}D_{m}(s_{n+1},a_{n+1})\right]$ (12)
+where $Q_{\eta_{m}}(s_{n},a_{n})$ represents the action-value function with a discount factor $\eta_{m}$.
+$D_{m}(s_{n},a_{n})$ denotes the delta function for SARSA ($\Delta$), while $\eta_{m}$ represents the discount factor for time-scale m, while $Q_{\eta_{m}}(s_{n+1},a_{n+1})$ is the SARSA update for action-value associated with the subsequent state-action pair.
 The single-step TD SARSA ($\Delta$) mirrors the original TD update, but it is applied to action-value functions Q rather than state-value functions V.
@@
 for the Q-value function in standard SARSA is as follows:
-$Q_{\gamma_{z}}(s_{t},a_{t})=\mathbb{E}\left[r_{t}+\gamma_{z}Q_{\gamma_{z}}(s_{t+1},a_{t+1})\right]$ (10)
-This denotes the expected value of the reward acquired at time t, in addition to the discounted value of the action-value function in the subsequent state-action pair $(s_{t+1},a_{t+1})$.
-Currently, rather than utilizing a singular discount factor $\gamma$, we are employing a sequence of discount factors $\gamma_{0},\gamma_{1},.....,\gamma_{z}$.
-The delta function $W_{z}(s_{t},a_{t})$ is defined as the difference between action-value functions associated with successive discount factors from Eq.
-6:
-$W_{z}(s_{t},a_{t})=Q_{\gamma_{z}}(s_{t},a_{t})-Q_{\gamma_{z-1}}(s_{t},a_{t})$ (11)
-To expand the delta function $W_{z}$ utilizing its definition, we replace the Bellman Eq.
-for both $Q_{\gamma_{z}}$ and $Q_{\gamma_{z-1}}$, thereby expressing $W_{z}$ in relation to these action-value functions.
+$Q_{\eta_{m}}(s_{n},a_{n})=\mathbb{E}\left[r_{n}+\eta_{n}Q_{\eta_{n}}(s_{n+1},a_{n+1})\right]$ (13)
+This denotes the expected value of the reward acquired at time n, in addition to the discounted value of the action-value function in the subsequent state-action pair $(s_{n+1},a_{n+1})$.
+Currently, rather than utilizing a singular discount factor $\eta$, we are employing a sequence of discount factors $\eta_{0},\eta_{1},.....,\eta_{m}$.
+The delta function $D_{m}(s_{n},a_{n})$ is defined as the difference between action-value functions associated with successive discount factors from Eq.
+9:
+$D_{m}(s_{n},a_{n})=Q_{\eta_{m}}(s_{n},a_{n})-Q_{\eta_{m-1}}(s_{n},a_{n})$ (14)
+To expand the delta function $D_{z}$ utilizing its definition, we replace the Bellman Eq.
+for both $Q_{\eta_{m}}$ and $Q_{\eta_{m-1}}$, thereby expressing $D_{m}$ in relation to these action-value functions.
 Commence by composing:
@@
 […]
-Since the reward terms $r_{t}$ are included in both formulations, simplify by eliminating them.
-$W_{z}(s_{t},a_{t})=\mathbb{E}\left[\gamma_{z}Q_{\gamma_{z}}(s_{t+1},a_{t+1})-\gamma_{z-1}Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})\right]$
+Since the reward terms $r_{n}$ are included in both formulations, simplify by eliminating them.
+$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\eta_{m}Q_{\eta_{m}}(s_{n+1},a_{n+1})-\eta_{m-1}Q_{\eta_{m-1}}(s_{n+1},a_{n+1})\right]$
 Decomposing the terms allows for a more detailed analysis of the expression. Utilizing the recursive relationship from Eq. 14
-$Q_{\gamma_{z}}(s_{t+1},a_{t+1})=W_{z}(s_{t+1},a_{t+1})+Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})$
-Substitute the definition of $W_{z}$ into the Eq. below.
-$W_{z}(s_{t},a_{t})=\mathbb{E}\left[\gamma_{z}(W_{z}(s_{t+1},a_{t+1})+Q_{\gamma_{z-1}}(s_{t+1},a_{t+1}))-\gamma_{z-1}Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})\right]$
+$Q_{\eta_{m}}(s_{n+1},a_{n+1})=D_{m}(s_{n+1},a_{n+1})+Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$
+Substitute the definition of $D_{m}$ into the Eq. below.
+$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\eta_{m}(D_{m}(s_{n+1},a_{n+1})+Q_{\eta_{m-1}}(s_{n+1},a_{n+1}))-\eta_{n-1}Q_{\eta_{n-1}}(s_{n+1},a_{n+1})\right]$
 Make the terms simpler:
-$W_{z}(s_{t},a_{t})=\mathbb{E}\left[\gamma_{z}W_{z}(s_{t+1},a_{t+1})+(\gamma_{z}-\gamma_{z-1})Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})\right]$
-Integrate the terms to derive the final update Eq. for $W_{z}(s_{t},a_{t})$.
-$W_{z}(s_{t},a_{t})=\mathbb{E}\left[(\gamma_{z}-\gamma_{z-1})Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})+\gamma_{z}W_{z}(s_{t+1},a_{t+1})\right]$ (13)
+$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\eta_{m}D_{m}(s_{n+1},a_{n+1})+(\eta_{m}-\eta_{m-1})Q_{\eta_{m-1}}(s_{n+1},a_{n+1})\right]$
+Integrate the terms to derive the final update Eq. for $D_{m}(s_{n},a_{n})$.
+$D_{m}(s_{n},a_{n})=\mathbb{E}\left[(\eta_{m}-\eta_{m-1})Q_{\eta_{m-1}}(s_{n+1},a_{n+1})+\eta_{m}D_{m}(s_{n+1},a_{n+1})\right]$ (16)
 This Eq.
-demonstrates that the update for $W_{z}$ integrates the disparity in discount factors $\gamma_{z}-\gamma_{z-1}$ with respect to the action-value function $Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})$ and the bootstrapping from the subsequent step $W_{z}(s_{t+1},a_{t+1})$.
+demonstrates that the update for $D_{m}$ integrates the disparity in discount factors $\eta_{m}-\eta_{m-1}$ with respect to the action-value function $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$ and the bootstrapping from the subsequent step $D_{m}(s_{n+1},a_{n+1})$.
 This denotes a Bellman Eq.
-for $W_{z}$, combining a decay factor $\gamma_{z}$ and the reward $Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})$.
-Consequently, it can be utilized to define the expected TD update for $W_{z}$.
-In this expression, $Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})$ can be expressed as the summation of $W_{z}(s_{t+1},a_{t+1})$ for $z\leq Z-1$, indicating that the Bellman Eq.
-for $W_{z}$ is contingent upon the values of all delta functions $W_{z}$ for $z\leq Z-1$.
+for $D_{m}$, combining a decay factor $\eta_{m}$ and the reward $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$.
+Consequently, it can be utilized to define the expected TD update for $D_{m}$.
+In this expression, $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$ can be expressed as the summation of $D_{m}(s_{n+1},a_{n+1})$ for $m\leq M-1$, indicating that the Bellman Eq.
+for $D_{m}$ is contingent upon the values of all delta functions $D_{m}$ for $m\leq M-1$.
 This approach treats the delta value function at each time-scale as an autonomous RL issue, with rewards obtained from the action-value function of the immediately preceding time-scale.
-Thus, for a target discounted action-value function $Q_{\gamma_{z}}(s,a)$, all delta components can be trained simultaneously through a TD update, utilizing the prior values of each estimator for bootstrapping.
-This process necessitates the assumption of a sequence of discount factors, denoted as $\gamma_{z}$, which includes both the minimum and maximum values, $\gamma_{0}$ and $\gamma_{z}$[8].
+Thus, for a target discounted action-value function $Q_{\eta_{m}}(s,a)$, all delta components can be trained simultaneously through a TD update, utilizing the prior values of each estimator for bootstrapping.
+This process necessitates the assumption of a sequence of discount factors, denoted as $\eta_{m}$, which includes both the minimum and maximum values, $\eta_{0}$ and $\eta_{m}$[8].
 4.3 Multi-Step TD ( SARSA($\Delta$) )
@@
 In multi-step TD SARSA($\Delta$), rewards are aggregated over multiple steps instead of depending on a single future reward.
-This approach takes into account the differences between consecutive discount factors, $\gamma_{z}$ and $\gamma_{z-1}$, while utilizing the value function at state $s_{t+k_{z}}$ for bootstrapping.
+This approach takes into account the differences between consecutive discount factors, $\eta_{m}$ and $\eta_{m-1}$, while utilizing the value function at state $s_{n+k_{m}}$ for bootstrapping.
 In SARSA($\Delta$), the update Eqs.
 are modified to aggregate rewards across a series of transitions prior to bootstrapping from either the Q-values or delta estimators.
-In multi-step SARSA($\Delta$), the agent aggregates rewards over K steps, leading to a corresponding modification in the update rule for each $W_{z}$.
-$W_{z}(s_{t},a_{t})=\mathbb{E}\left[\sum_{j=1}^{k_{z}-1}(\gamma_{z}^{j}-\gamma_{z-1}^{j})r_{t+j}+(\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}})Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})+\gamma_{z}^{k_{z}}W_{z}(s_{t+k_{z}},a_{t+k_{z}})\right]$ (14)
-In this case, $r_{t+j}$ represents the reward obtained at time step $t+j$.
-$a_{t+k_{z}}$ and the action is chosen greedily as $a=\arg\max_{a}Q(s_{t+k_{z}},a)$, where $k_{z}$ is the number of steps linked to the discount factor $\gamma_{z}$.
+In multi-step SARSA($\Delta$), the agent aggregates rewards over K steps, leading to a corresponding modification in the update rule for each $D_{m}$.
+$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=1}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{x}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})\right]$ (17)
+In this case, $r_{m+x}$ represents the reward obtained at time step $n+x$.
+$a_{n+k_{m}}$ and the action is chosen greedily as $a=\arg\max_{a}Q(s_{n+k_{m}},a)$, where $k_{m}$ is the number of steps linked to the discount factor $\eta_{m}$.
 This method generalizes Q-learning to multi-step TD learning by aggregating rewards across several stages and utilizing bootstrapping from both the current and prior time-scale action-value functions.
 In normal multi-step TD learning, the TD error is computed by aggregating rewards over several stages rather than relying on a single future step, followed by bootstrapping from the action-value at the last step.
-For example, utilizing a singular discount factor $\gamma$, the multi-step TD update equation is expressed as:
-$Q(s_{t},a_{t})=\mathbb{E}\left[\sum_{j=0}^{k-1}\gamma^{j}r_{t+j}+\gamma^{k}Q(s_{t+k},a_{t+k})\right]$ (15)
+For example, utilizing a singular discount factor $\eta$, the multi-step TD update equation is expressed as:
+$Q(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k-1}\eta^{x}r_{n+x}+\eta^{k}Q(s_{n+k},a_{n+k})\right]$ (18)
 Herein, The initial segment of the Eq.
-aggregates the rewards across k steps, each diminished by $\gamma$.
-The second component derives from the action-value function at time step t+k.
-We expand this methodology to include various discount factors and delta decomposition, concentrating on representing the delta function $W_{z}(s_{t},a_{t})$ as the difference between action-value functions with discount factors $\gamma_{z}$ and $\gamma_{z-1}$.
-According to the definition of $W_{z}$, we obtain:
-$W_{z}(s_{t},a_{t})=Q_{\gamma_{z}}(s_{t},a_{t})-Q_{\gamma_{z-1}}(s_{t},a_{t})$
+aggregates the rewards across k steps, each diminished by $\eta$.
+The second component derives from the action-value function at time step n+k.
+We expand this methodology to include various discount factors and delta decomposition, concentrating on representing the delta function $D_{m}(s_{n},a_{n})$ as the difference between action-value functions with discount factors $\eta_{m}$ and $\eta_{m-1}$.
+According to the definition of $D_{m}$, we obtain:
+$D_{m}(s_{n},a_{n})=Q_{\eta_{m}}(s_{n},a_{n})-Q_{\eta_{m-1}}(s_{n},a_{n})$
 We now implement the multi-step variant of the Bellman Eq.
-for both $Q_{\gamma_{z}}(s_{t},a_{t})$ and $Q_{\gamma_{z-1}}(s_{t},a_{t})$.
+for both $Q_{\eta_{m}}(s_{n},a_{n})$ and $Q_{\eta_{m-1}}(s_{n},a_{n})$.
 The multi-step Bellman Eq.
-for $Q_{\gamma_{z}}(s_{t},a_{t})$ is expressed as follows:
-$Q_{\gamma_{z}}(s_{t},a_{t})=\mathbb{E}\left[\sum_{j=0}^{k_{z}-1}\gamma_{z}^{j}r_{t+j}+\gamma_{z}^{k_{z}}Q_{\gamma_{z}}(s_{t+k_{z}},a_{t+k_{z}})\right]$
-For $Q_{\gamma_{z-1}}(s_{t},a_{t})$, the multi-step Bellman Eq. is expressed as follows:
-$Q_{\gamma_{z-1}}(s_{t},a_{t})=\mathbb{E}\left[\sum_{j=0}^{k_{z}-1}\gamma_{z-1}^{j}r_{t+j}+\gamma_{z-1}^{k_{z}}Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})\right]$
+for $Q_{\eta_{m}}(s_{n},a_{n})$ is expressed as follows:
+$Q_{\eta_{m}}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k_{m}-1}\eta_{m}^{x}r_{n+x}+\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})\right]$
+For $Q_{\eta_{m-1}}(s_{n},a_{n})$, the multi-step Bellman Eq. is expressed as follows:
+$Q_{\eta_{m-1}}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k_{m}-1}\eta_{m-1}^{x}r_{n+x}+\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})\right]$
 We will now utilize the delta estimator formulas to subtract these two Eqs.
-$W_{z}(s_{t},a_{t})=\left[\sum_{j=0}^{k_{z}-1}\gamma_{z}^{j}r_{t+j}+\gamma_{z}^{k_{z}}Q_{\gamma_{z}}(s_{t+k_{z}},a_{t+k_{z}})\right]-\left[\sum_{j=0}^{k_{z}-1}\gamma_{z-1}^{j}r_{t+j}+\gamma_{z-1}^{k_{z}}Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})\right]$
+$D_{m}(s_{n},a_{n})=\left[\sum_{x=0}^{k_{m}-1}\eta_{m}^{x}r_{n+x}+\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})\right]-\left[\sum_{x=0}^{k_{m}-1}\eta_{m-1}^{x}r_{n+x}+\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})\right]$
 We will now broaden and simplify the expressions.
-The immediate reward terms $r_{t+j}$ are present in both Eqs.
-but are adjusted by distinct discount factors, $\gamma_{z}$ and $\gamma_{z-1}$.
+The immediate reward terms $r_{n+x}$ are present in both Eqs.
+but are adjusted by distinct discount factors, $\eta_{m}$ and $\eta_{m-1}$.
 This enables us to express the disparity in rewards as:
-$\sum_{j=0}^{k_{z}-1}\gamma_{z}^{j}r_{t+j}-\sum_{j=0}^{k_{z}-1}\gamma_{z-1}^{j}r_{t+j}=\sum_{j=0}^{k_{z}-1}(\gamma_{z}^{j}-\gamma_{z-1}^{j})r_{t+j}$
+$\sum_{x=0}^{k_{m}-1}\eta_{m}^{x}r_{n+x}-\sum_{x=0}^{k_{m}-1}\eta_{m-1}^{x}r_{n+x}=\sum_{x=0}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}$
 Subsequently, for the bootstrapping terms, we utilize the recursive relationship as outlined:
-$Q_{\gamma_{z}}(s_{t+k_{z}},a_{t+k_{z}})=W_{z}(s_{t+k_{z}},a_{t+k_{z}})+Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})$
+$Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})=D_{m}(s_{n+k_{m}},a_{n+k_{m}})+Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})$
 As a result, we get:
-$\gamma_{z}^{k_{z}}Q_{\gamma_{z}}(s_{t+k_{z}},a_{t+k_{z}})-\gamma_{z-1}^{k_{z}}Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})=(\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}})Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})+\gamma_{z}^{k_{z}}W_{z}(s_{t+k_{z}},a_{t+k_{z}})$
+$\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})-\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})=(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})$
 We now amalgamate the terms for both the rewards and bootstrapping to produce the final expression:
-$W_{z}(s_{t},a_{t})=\mathbb{E}\left[\sum_{j=0}^{k_{z}-1}(\gamma_{z}^{j}-\gamma_{z-1}^{j})r_{t+j}+(\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}})Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})+\gamma_{z}^{k_{z}}W_{z}(s_{t+k_{z}},a_{t+k_{z}})\right]$ (16)
+$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})\right]$ (19)
 This Eq.
 extends the conventional multi-step TD update to include various multiple discount factors and action-value functions in RL.
-As a result, each $W_{z}$ obtains a share of the rewards from the environment up to time-step $k_{z}-1$.
-Furthermore, each $W_{z}$ utilizes its distinct action-value function in conjunction with the value from the prior time-scale.
+As a result, each $D_{m}$ obtains a share of the rewards from the environment up to time-step $k_{m}-1$.
+Furthermore, each $D_{m}$ utilizes its distinct action-value function in conjunction with the value from the prior time-scale.
 A variation of this algorithm, utilizing k-step bootstrapping as described in [29], is presented in Algorithm 1.
-Despite Algorithm 1 exhibiting quadratic complexity in relation to Z, it can be executed with linear complexity for substantial Z by preserving $\hat{Q}$ action-values at each time-scale $\gamma_{z}$.
+Despite Algorithm 1 exhibiting quadratic complexity in relation to M, it can be executed with linear complexity for substantial M by preserving $\hat{Q}$ action-values at each time-scale $\eta_{m}$.
 Algorithm 1 Multi-Step TD ( SARSA($\Delta$) )
-Inputs: Pick out the discount factors ($\gamma_{0},\gamma_{1},...,\gamma_{Z}$), bootstrapping steps ($k_{0},k_{1},...,k_{Z}$), and learning rates ($\alpha_{0},\alpha_{1},...,\alpha_{Z}$).
-Set the initial value of $W_{z}(s,a)=0$ for all states, actions, and scales z.
+Inputs: Pick out the discount factors ($\eta_{0},\eta_{1},...,\eta_{M}$), bootstrapping steps ($k_{0},k_{1},...,k_{M}$), and learning rates ($\alpha_{0},\alpha_{1},...,\alpha_{M}$).
+Set the initial value of $D_{m}(s,a)=0$ for all states, actions, and scales m.
 for episode = 0, 1, 2,… $\triangleright$ Loop for each episode
      Initialize state $s_{0}$ and choose an initial action $a_{0}$ in accordance with a policy.
-     for t = 0, 1, 2,… do $\triangleright$ Loop for each time step
-         Take action $a_{t}$, observe reward $r_{t}$ and next state $s_{t+1}$.
-         Select action $a_{t+1}$ based on a policy.
-         for z = 0, 1, …, Z do
-              if z = 0 then
-                  $G^{0}=\sum_{j=0}^{k_{0}-1}\gamma_{0}^{j}r_{t+j}+\gamma_{0}^{k_{0}}W_{0}(s_{t+k_{0}},a_{t+k_{0}})$
-              $\triangleright$ Utilizing Eq.7, we substitute $Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})$ by summing the W-components up to $W_{z-1}$ in Eq.4.3.
-                  $G^{z}=\sum_{j=0}^{k_{z}-1}(\gamma_{z}^{j}-\gamma_{z-1}^{j})r_{t+j}+(\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}})\sum_{z=0}^{z-1}W_{z-1}(s_{t+k_{z}},a_{t+k_{z}})+\gamma_{z}^{k_{z}}W_{z}(s_{t+k_{z}},a_{t+k_{z}})$
+     for n = 0, 1, 2,… do $\triangleright$ Loop for each time step
+         Take action $a_{n}$, observe reward $r_{n}$ and next state $s_{n+1}$.
+         Select action $a_{n+1}$ based on a policy.
+         for m = 0, 1, …, M do
+              if m = 0 then
+                  $G^{0}=\sum_{x=0}^{k_{0}-1}\eta_{0}^{x}r_{n+x}+\eta_{0}^{k_{0}}D_{0}(s_{n+k_{0}},a_{n+k_{0}})$
+              $\triangleright$ Utilizing Eq.10, we substitute $Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})$ by summing the D-components up to $D_{m-1}$ in Eq.4.3.
+                  $G^{m}=\sum_{x=0}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})\sum_{m=0}^{m-1}D_{m-1}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})$
               end if
          end for
-         for z = 0, 1, 2,…, Z do
-              $W_{z}(s_{t},a_{t})\leftarrow W_{z}(s_{t},a_{t})+\alpha_{z}(G^{z}-W_{z}(s_{t},a_{t}))$
+         for m = 0, 1, 2,…, M do
+              $D_{m}(s_{n},a_{n})\leftarrow D_{m}(s_{n},a_{n})+\alpha_{m}(G^{m}-D_{m}(s_{n},a_{n}))$
          end for
@@
 4.4 SARSA TD($\lambda,\Delta$)
-Eq.17 introduces the $\lambda$-return[23, 1], which combines rewards over multiple steps to create a target for TD($\lambda$) updates. The $\lambda$-return $G_{t}^{\gamma,\lambda}$ is defined as:
-$G_{t}^{\gamma,\lambda}(s_{t},a_{t})=\hat{Q}_{\gamma}(s_{t},a_{t})+\sum_{k=0}^{\infty}(\lambda\gamma)^{k}\delta_{t+k}^{\gamma}$ (17)
+Eq.20 introduces the $\lambda$-return[23, 1], which combines rewards over multiple steps to create a target for TD($\lambda$) updates. The $\lambda$-return $G_{n}^{\eta,\lambda}$ is defined as:
+$G_{n}^{\eta,\lambda}(s_{n},a_{n})=\hat{Q}_{\eta}(s_{n},a_{n})+\sum_{k=0}^{\infty}(\lambda\eta)^{k}\delta_{n+k}^{\eta}$ (20)
 Eq.21 defines the TD($\lambda$) operator, which is used to iteratively apply $\lambda$-returns in updating the value functions.
 For SARSA, the TD($\lambda$) operator updates the action-value function by summing the $\lambda$-discounted TD errors as follows:
-$\mathrm{T}_{\lambda}Q(s_{t},a_{t})=Q(s_{t},a_{t})+(I-\lambda\gamma P)^{-1}(\mathrm{T}Q(s_{t},a_{t})-Q(s_{t},a_{t}))$ (18)
+$\mathrm{T}_{\lambda}Q(s_{n},a_{n})=Q(s_{n},a_{n})+(I-\lambda\eta P)^{-1}(\mathrm{T}Q(s_{n},a_{n})-Q(s_{n},a_{n}))$ (21)
 Here,P is the transition matrix for state-action pairs under the policy $\pi$, and Q is the action-value function.
-Similarly, Eq.19 defines the $\lambda$-return specific to delta estimators in TD($\lambda,\Delta$), denoted as $G_{t}^{z,\lambda_{z}}$ for each delta estimator $W_{z}$:
-$G_{t}^{z,\lambda_{z}}:=\hat{W}_{z}(s_{t},a_{t})+\sum_{k=0}^{\infty}(\lambda_{z}\gamma_{z})^{k}\delta_{t+k}^{z}$ (19)
-where $\delta_{t}^{0}:=\delta_{t}^{\gamma_{0}}$ and $\delta_{t}^{z}:=(\gamma_{z}-\gamma_{z-1})\hat{Q}_{\gamma_{z-1}}(s_{t+1},a_{t+1})+\gamma_{z}\hat{W}_{z}(s_{t+1},a_{t+1})-\hat{W}_{z}(s_{t+1},a_{t+1})$ are the TD-errors.
+Similarly, Eq.22 defines the $\lambda$-return specific to delta estimators in TD($\lambda,\Delta$), denoted as $G_{n}^{m,\lambda_{m}}$ for each delta estimator $D_{m}$:
+$G_{n}^{m,\lambda_{m}}:=\hat{D}_{m}(s_{n},a_{n})+\sum_{k=0}^{\infty}(\lambda_{m}\eta_{m})^{k}\delta_{n+k}^{m}$ (22)
+where $\delta_{n}^{0}:=\delta_{n}^{\eta_{0}}$ and $\delta_{n}^{m}:=(\eta_{m}-\eta_{m-1})\hat{Q}_{\eta_{m-1}}(s_{n+1},a_{n+1})+\eta_{m}\hat{D}_{m}(s_{n+1},a_{n+1})-\hat{D}_{m}(s_{n+1},a_{n+1})$ are the TD-errors.
 4.5 SARSA TD($\lambda,\Delta$) with Generalized Advantage Estimation (GAE)
 Generalized Advantage Estimation (GAE) [28] seeks to compute the advantage function by aggregating multi-step TD errors.
-To apply this approach to delta estimators $W_{z}(s_{t},a_{t})$, we compute advantage estimates $A^{\Delta}(s_{t},a_{t})$ for each time-scale.
-Each advantage estimate $A^{\Delta}(s_{t},a_{t})$ employs a multi-step TD error pertinent to the delta estimator $W_{z}$, expressed as follows:
-$A^{\Delta}(s_{t},a_{t})=\sum_{k=0}^{T-1}(\lambda_{z}\gamma_{z})^{k}\delta_{t+k}^{\Delta}$ (20)
-where $\delta_{t+k}^{\Delta}:=r_{t}+\gamma_{z}\sum_{z=0}^{Z}\hat{W}_{z}(s_{t+1},a_{t+1})-\sum_{z=0}^{Z}\hat{W}_{z}(s_{t},a_{t})$.
-We utilize $\gamma_{Z}$ as the discount factor and the sum of all W estimators as a surrogate for $Q_{\gamma z}$.
+To apply this approach to delta estimators $D_{m}(s_{n},a_{n})$, we compute advantage estimates $A^{\Delta}(s_{n},a_{n})$ for each time-scale.
+Each advantage estimate $A^{\Delta}(s_{n},a_{n})$ employs a multi-step TD error pertinent to the delta estimator $D_{n}$, expressed as follows:
+$A^{\Delta}(s_{n},a_{n})=\sum_{k=0}^{T-1}(\lambda_{m}\eta_{m})^{k}\delta_{n+k}^{\Delta}$ (23)
+where $\delta_{n+k}^{\Delta}:=r_{n}+\eta_{m}\sum_{m=0}^{M}\hat{D}_{m}(s_{n+1},a_{n+1})-\sum_{m=0}^{M}\hat{D}_{m}(s_{n},a_{n})$.
+We utilize $\eta_{m}$ as the discount factor and the sum of all D estimators as a surrogate for $Q_{\eta m}$.
 This objective is applicable to PPO by implementing the policy update from Eq.7 and replacing A with $A^{\Delta}$.
-Additionally, to train each $W_{z}$, we utilize a truncated form of their corresponding $\lambda$-return, as outlined in Eq.19.
+Additionally, to train each $D_{m}$, we utilize a truncated form of their corresponding $\lambda$-return, as outlined in Eq.22.
 For more information, see Algorithm 2.
 Algorithm 2 PPO-TD($\lambda$, SARSA($\Delta$))
-Inputs: Pick out the discount factors ($\gamma_{0},\gamma_{1},...,\gamma_{Z}$), bootstrapping steps ($k_{0},k_{1},...,k_{Z}$), and learning rates ($\alpha_{0},\alpha_{1},...,\alpha_{Z}$).
-Set the initial value of $W_{z}(s,a)=0$ for all states s, actions a, and scales z.
-Initialize policy $\omega$, and values $\theta^{z}\forall z$
+Inputs: Pick out the discount factors ($\eta_{0},\eta_{1},...,\eta_{M}$), bootstrapping steps ($k_{0},k_{1},...,k_{M}$), and learning rates ($\alpha_{0},\alpha_{1},...,\alpha_{M}$).
+Set the initial value of $D_{m}(s,a)=0$ for all states s, actions a, and scales m.
+Initialize policy $\nu$, and values $\theta^{m}\forall m$
 for episode = 0, 1, 2,… $\triangleright$ Loop for each episode
      Initialize state $s_{0}$ and choose an initial action $a_{0}$ in accordance with a policy.
-     for t = 0, 1, 2,… do $\triangleright$ Loop for each time step
-         Take action $a_{t}$, observe reward $r_{t}$ and next state $s_{t+1}$.
-         Select next action $a_{t+1}$ based on a policy.
-         for z = 0, 1, …, Z do
-              if $t\geq T$ then
-                  $G^{z,\lambda_{z}}\leftarrow\hat{W}_{z}(s_{t-T},a_{t-T})+\sum_{k=0}^{T-1}(\lambda_{z}\gamma_{z})^{k}\delta_{t-T+k}^{z}\forall z$triangleright Computing multi-step return $G^{z,\lambda_{z}}$ and TD-error $\delta_{t-T+k}^{z}$ using Eq.19.
+     for n = 0, 1, 2,… do $\triangleright$ Loop for each time step
+         Take action $a_{n}$, observe reward $r_{n}$ and next state $s_{n+1}$.
+         Select next action $a_{n+1}$ based on a policy.
+         for m = 0, 1, …, M do
+              if $n\geq T$ then
+                  $G^{m,\lambda_{m}}\leftarrow\hat{D}_{m}(s_{n-T},a_{n-T})+\sum_{k=0}^{T-1}(\lambda_{m}\eta_{m})^{k}\delta_{n-T+k}^{m}\forall m$triangleright Computing multi-step return $G^{m,\lambda_{m}}$ and TD-error $\delta_{n-T+k}^{m}$ using Eq.22.
               end if
          end for
-         for z = 0, 1, 2,…, Z do
-              $\hat{W}_{z}(s_{t-T},a_{t-T})\leftarrow\hat{W}_{z}(s_{t-T},a_{t-T})+\alpha_{z}\left(G^{z,\lambda_{z}}-\hat{W}_{z}(s_{t-T},a_{t-T})\right)$
-              $A^{\Delta}=\sum_{k=0}^{T-1}(\lambda_{z}\gamma_{z})^{k}\delta_{t-T+k}^{\Delta}$triangleright Where $A^{\Delta}$ and $\delta_{t-T+k}^{\Delta}$ are computed using Eq.20.
-              $\theta^{z}\leftarrow\theta^{z}+\alpha_{z}\left(G^{z,\lambda_{z}}-\hat{W}_{z}(s_{t-T},a_{t-T})\right)\nabla\hat{W}_{z}(s_{t-T},a_{t-T})$triangleright Update $\theta^{z}$ with TD (Eq.3) using $G^{z,\lambda_{z}}\forall z$.
-              $\omega\leftarrow\omega+\alpha_{\omega}A^{\Delta}\nabla\log\pi_{\omega}(a_{t}|s_{t})$triangleright Update the policy parameters $\omega$ with PPO (Eq.4) for SARSA using $A^{\Delta}$
+         for m = 0, 1, 2,…, M do
+              $\hat{D}_{m}(s_{n-T},a_{n-T})\leftarrow\hat{D}_{m}(s_{n-T},a_{n-T})+\alpha_{m}\left(G^{m,\lambda_{m}}-\hat{D}_{m}(s_{n-T},a_{n-T})\right)$
+              $A^{\Delta}=\sum_{k=0}^{T-1}(\lambda_{m}\eta_{m})^{k}\delta_{n-T+k}^{\Delta}$triangleright Where $A^{\Delta}$ and $\delta_{n-T+k}^{\Delta}$ are computed using Eq.23.
+              $\theta^{m}\leftarrow\theta^{m}+\alpha_{m}\left(G^{m,\lambda_{m}}-\hat{D}_{m}(s_{n-T},a_{n-T})\right)\nabla\hat{D}_{m}(s_{n-T},a_{n-T})$triangleright Update $\theta^{m}$ with TD (Eq.3) using $G^{m,\lambda_{m}}\forall m$.
+              $\mathcal{L}(\nu)=\mathbb{E}\bigg{[}min\bigg{(}\rho(\nu)A^{\Delta}(s,a),clip(\rho(\nu),1-\epsilon,1+\epsilon)A^{\Delta}(s,a)\bigg{)}\bigg{]}$triangleright from Eq.7
+              $\nu\leftarrow\nu+\alpha_{\nu}\nabla_{\nu}\mathcal{L}(\nu)$triangleright Update the policy parameters $\nu$ with PPO (Eq.7) for SARSA using $A^{\Delta}$
          end for
@@
 Next, We examine our delta estimators through the lens of the bias-variance trade-off.
-In SARSA, bias emerges from dependence on current estimates, affecting updates over short time-scales, particularly with smaller discount factors, such as lower values of $\gamma$ in $W_{z}$.
+In SARSA, bias emerges from dependence on current estimates, affecting updates over short time-scales, particularly with smaller discount factors, such as lower values of $\eta$ in $D_{m}$.
 Updates utilizing smaller discount factors demonstrate reduced variance due to their reliance on current estimates; however, they exhibit increased bias as they fail to integrate substantial information regarding future rewards.
 In contrast, variance escalates for updates associated with extended time-scales, as future reward information adds further variability.
-In TD($\Delta$), action-value components $W_{z}$ with higher $\gamma_{z}$ are associated with greater long-term rewards; however, this results in increased variance stemming from the stochastic characteristics of reward sequences.
-We begin by demonstrating that our estimator is equivalent to the standard estimator $\hat{Q}_{\gamma}$ under specific conditions, as outlined in Theorem 1.
-This comparison elucidates the essential metrics of our estimator that may indicate potential advantages over the standard $\hat{Q}_{\gamma}$ estimator.
+In TD($\Delta$), action-value components $D_{m}$ with higher $\eta_{m}$ are associated with greater long-term rewards; however, this results in increased variance stemming from the stochastic characteristics of reward sequences.
+We begin by demonstrating that our estimator is equivalent to the standard estimator $\hat{Q}_{\eta}$ under specific conditions, as outlined in Theorem 1.
+This comparison elucidates the essential metrics of our estimator that may indicate potential advantages over the standard $\hat{Q}_{\eta}$ estimator.
 Building on this result and previous research by Kearns and Singh[9], we modify these analyses for SARSA to investigate the effects of bias and variance, along with the TD($\Delta$) decomposition framework, in relation to action-value functions.
@@
 This discussion centers on linear function approximation of the specified form:
-$\hat{Q}(s,a)_{\gamma}:=\langle\theta^{\gamma},\phi(s,a)\rangle\quad\text{and}\quad\hat{W}_{z}(s,a):=\langle\theta^{z},\phi(s,a)\rangle,\forall z$
-where $\theta$ and $\theta^{z}$ represent weight vectors in $\mathbb{R}^{d}$ and the function $\phi:S\times A\rightarrow\mathbb{R}^{d}$ represents a mapping from a state-action pair to a specified d-dimensional feature space.
+$\hat{Q}(s,a)_{\eta}:=\langle\theta^{\eta},\phi(s,a)\rangle\quad\text{and}\quad\hat{D}_{m}(s,a):=\langle\theta^{m},\phi(s,a)\rangle,\forall m$
+where $\theta$ and $\theta^{m}$ represent weight vectors in $\mathbb{R}^{d}$ and the function $\phi:S\times A\rightarrow\mathbb{R}^{d}$ represents a mapping from a state-action pair to a specified d-dimensional feature space.
 The weight vector $\theta$ for SARSA is updated according to the TD($\lambda$) learning rule in the following manner:
-$\theta_{t+1}^{\gamma}=\theta_{t}^{\gamma}+\alpha\bigg{(}G_{t}^{\gamma,\lambda}-\hat{Q}_{\gamma}(s,a)\bigg{)}\phi(s_{t},a_{t}),$ (21)
-where $G_{t}^{\gamma,\lambda}$ represents the TD($\lambda$) return defined in Eq.17. Likewise, the delta estimator approach TD($\lambda_{z},\Delta$) is used to update each $\hat{W}_{z}$:
-$\theta_{t+1}^{z}=\theta_{t}^{z}+\alpha\bigg{(}G_{t}^{Z,\lambda_{z}}-\hat{W}_{z}(s_{t},a_{t})\bigg{)}\phi(s_{t},a_{t}),$ (22)
-where $G_{t}^{Z,\lambda_{z}}$ is specified identically to the TD($\Delta$) return defined in Eq.19, modified for the particular action-value function.
-In these equations, $\alpha$ and $\{\alpha_{z}\}_{z}$ represent positive learning rates.
+$\theta_{n+1}^{\eta}=\theta_{n}^{\eta}+\alpha\bigg{(}G_{n}^{\eta,\lambda}-\hat{Q}_{\eta}(s,a)\bigg{)}\phi(s_{n},a_{n}),$ (24)
+where $G_{n}^{\eta,\lambda}$ represents the TD($\lambda$) return defined in Eq.20. Likewise, the delta estimator approach TD($\lambda_{m},\Delta$) is used to update each $\hat{D}_{m}$:
+$\theta_{n+1}^{m}=\theta_{n}^{m}+\alpha\bigg{(}G_{n}^{M,\lambda_{m}}-\hat{D}_{m}(s_{n},a_{n})\bigg{)}\phi(s_{n},a_{n}),$ (25)
+where $G_{n}^{M,\lambda_{m}}$ is specified identically to the TD($\Delta$) return defined in Eq.22, modified for the particular action-value function.
+In these equations, $\alpha$ and $\{\alpha_{m}\}_{m}$ represent positive learning rates.
 The following theorem demonstrates the equivalence of two SARSA version algorithms in this context.
@@
 Theorem 1.
-If $\alpha_{z}=\alpha$, $\lambda_{z}\gamma_{z}=\lambda_{\gamma}$, $\forall z$, and if we choose the initial conditions in such a way that $\sum_{z=0}^{Z}\theta_{0}^{\gamma}=\theta_{0}^{\gamma}$, then the iterates produced by TD($\lambda$) (Eq.
-21) and TD($\lambda,\Delta$) (Eq.
-22) with linear function approximation satisfy the following conditions:
-$\Sigma_{z=0}^{Z}\theta_{t}^{z}=\theta_{t}^{\gamma}$ (23)
+Inspired by Romoff et al.[8], if $\alpha_{m}=\alpha$, $\lambda_{m}\eta_{m}=\lambda_{\eta}$, $\forall m$, and if we choose the initial conditions in such a way that $\sum_{m=0}^{M}\theta_{0}^{\eta}=\theta_{0}^{\eta}$, then the iterates produced by TD($\lambda$) (Eq.
+24) and TD($\lambda,\Delta$) (Eq.
+25) with linear function approximation satisfy the following conditions:
+$\Sigma_{m=0}^{M}\theta_{n}^{m}=\theta_{n}^{\eta}$ (26)
 The proof is presented in appendix PROOFS.
-Equivalence is attained when $\lambda_{z}\gamma_{z}=\lambda_{\gamma},\forall z$.
-When $\lambda$ approaches 1 and $\gamma_{z}$ is less than $\gamma$, this condition suggests that $\lambda_{z}=\lambda_{\gamma}/\gamma_{z}$ could potentially surpass one, resulting in a risk of divergence in TD($\lambda_{z}$).
-The subsequent theorem demonstrates that the TD($\lambda$) operator, as defined in Eq.18, qualifies as a contraction mapping for the range $1\leq\lambda<\frac{1+\gamma}{2\gamma}$, thereby confirming that $\lambda_{\gamma}<1$[8].
+Equivalence is attained when $\lambda_{m}\eta_{m}=\lambda_{\eta},\forall m$.
+When $\lambda$ approaches 1 and $\eta_{m}$ is less than $\eta$, this condition suggests that $\lambda_{m}=\lambda_{\eta}/\eta_{m}$ could potentially surpass one, resulting in a risk of divergence in TD($\lambda_{m}$).
+The subsequent theorem demonstrates that the TD($\lambda$) operator, as defined in Eq.21, qualifies as a contraction mapping for the range $1\leq\lambda<\frac{1+\eta}{2\eta}$, thereby confirming that $\lambda_{\eta}<1$[8].
 Theorem 2.
-$\forall\lambda\in\left[0,\frac{1+\gamma}{2\gamma}\right]$, the operator $T_{\lambda}$ is defined by the Eq.
-$T_{\lambda}Q=Q+(I-\lambda\gamma P)^{-1}(TQ-Q)$ $,\quad\forall Q\in\mathbb{R}^{|S|\times|A|}$, is well-defined.
-Furthermore, $T_{\lambda}Q$ constitutes a contraction in relation to the max norm, with its contraction coefficient expressed as $\frac{\gamma}{|1-\lambda\gamma|}$.
+Inspired by Romoff et al.[8], $\forall\lambda\in\left[0,\frac{1+\eta}{2\eta}\right]$, the operator $T_{\lambda}$ is defined by the Eq.
+$T_{\lambda}Q=Q+(I-\lambda\eta P)^{-1}(TQ-Q)$ $,\quad\forall Q\in\mathbb{R}^{|S|\times|A|}$, is well-defined.
+Furthermore, $T_{\lambda}Q$ constitutes a contraction in relation to the max norm, with its contraction coefficient expressed as $\frac{\eta}{|1-\lambda\eta|}$.
 The proof is presented in appendix PROOFS.
-The analysis presented in Theorem 1 can be adapted to a new context in which $k_{z}$-step TD ($\Delta$) is employed for each $W_{z}$ component, rather than TD($\lambda$, $\Delta$), a variant of TD learning.
-Theorem 1 demonstrates that with linear function approximation, standard multi-step TD and multi-step TD($\Delta$) can be equivalent when the number of steps ($k_{z}$) is consistent across all time scales (i.e., $k_{z}=k,\forall z$).
+The analysis presented in Theorem 1 can be adapted to a new context in which $k_{m}$-step TD ($\Delta$) is employed for each $D_{m}$ component, rather than TD($\lambda$, $\Delta$), a variant of TD learning.
+Theorem 1 demonstrates that with linear function approximation, standard multi-step TD and multi-step TD($\Delta$) can be equivalent when the number of steps ($k_{m}$) is consistent across all time scales (i.e., $k_{m}=k,\forall m$).
 Although these methods could theoretically be equivalent, such equivalence is not common.
@@
 The effective learning rate varies across different time scales, making it challenging to apply a uniform learning rate in this context.
-In addition to the learning rate, the approach of decomposing the action-value function into various components (specifically, $W_{z}$ components) presents advantages that traditional TD learning, including non-delta estimators or higher $W_{z}$ components, lacks.
-The ability to utilize various k-step returns (or $\lambda$-return) across different time scales provides enhanced control and adaptability in the learning process.
-If $k_{z}<k_{z+1},\forall z\quad(or\quad\gamma_{z}\lambda_{z}<\gamma_{z+1}\lambda_{z+1},\forall z)$, this approach may decrease variance while introducing bias, as shorter time scales, which can be learned more rapidly, may not align perfectly with longer time scales in terms of the values they estimate.
-5.2 Evaluation for Minimizing $k_{z}$ Values in SARSA via Phased Updates
+Besides the learning rate, the method of dissecting the action-value function into distinct components (particularly, $D_{m}$ components) offers benefits that conventional TD learning, including non-delta estimators or elevated $D_{m}$ components, does not possess.
+The capacity to employ diverse k-step returns (or $\lambda$-return) across many time-scales offers improved control and flexibility in the learning process.
+If $k_{m}<k_{m+1},\forall m\quad(or\quad\eta_{m}\lambda_{m}<\eta_{m+1}\lambda_{m+1},\forall m)$, this approach may decrease variance while introducing bias, as shorter time-scales, which can be learned more rapidly, may not align perfectly with longer time scales in terms of the values they estimate.
+5.2 Evaluation for Minimizing $k_{m}$ Values in SARSA via Phased Updates
 To show that how our approach differ from the single estimator case, let’s follow the tabular phased version of k-step TD presented by kearns & Singh[9].
-In SARSA, we aim to estimate the value of a state-action pair (s, a) start from each state $s\in\mathcal{S}$ and with action $a\in\mathcal{A}$, then n trajectories generated $\bigg{\{}S_{0}^{(j)}=s,a_{0},\gamma_{0},...,S_{k}^{(j)},a_{k}^{(j)},\gamma_{k}^{(j)},S_{k+1}^{(j)},...\bigg{\}}_{1\leq j\leq n}$ by following the policy $\pi$ and averaging over trajectories.
-Let $Q_{\gamma,t}(s,a)$ denote the Q-value estimate at phase t for the discount factor $\gamma$.
-For each iteration t, also called phase t, to get the phase-based estimate $Q_{\gamma,t}(s,a)$ for (s, a), we average over the trajectories, resulting in:
-$\hat{Q}_{\gamma,t}(s,a)=\frac{1}{n}\sum_{j=1}^{n}\bigg{(}\sum_{i=0}^{k-1}\gamma^{i}r_{i}^{(j)}+\gamma^{k}\hat{Q}_{\gamma,t-1}(s_{k}^{(j)},a_{k}^{(j)})\bigg{)}$ (24)
-Theorem 3 below from the theorems of kearns & Singh[9] but they show it for state value function while we show an upper bound on the error for action-value function estimation defined by $\Delta_{t}^{\hat{Q}_{\gamma}}=max_{s,a}\bigg{\{}\bigg{|}\hat{Q}_{\gamma,t}(s,a)-Q_{\gamma}(s,a)\bigg{|}\bigg{\}}$
+In SARSA, we aim to estimate the value of a state-action pair (s, a) start from each state $s\in\mathcal{S}$ and with action $a\in\mathcal{A}$, then n trajectories generated $\bigg{\{}S_{0}^{(x)}=s,a_{0},\eta_{0},...,S_{k}^{(x)},a_{k}^{(x)},\eta_{k}^{(x)},S_{k+1}^{(x)},...\bigg{\}}_{1\leq x\leq n}$ by following the policy $\pi$ and averaging over trajectories.
+Let $Q_{\eta,n}(s,a)$ denote the Q-value estimate at phase n for the discount factor $\eta$.
+For each iteration n, also called phase n, to get the phase-based estimate $Q_{\eta,n}(s,a)$ for (s, a), we average over the trajectories, resulting in:
+$\hat{Q}_{\eta,n}(s,a)=\frac{1}{n}\sum_{x=1}^{n}\bigg{(}\sum_{i=0}^{k-1}\gamma^{i}r_{i}^{(x)}+\eta^{k}\hat{Q}_{\eta,n-1}(s_{k}^{(x)},a_{k}^{(x)})\bigg{)}$ (27)
+Theorem 3 below from the theorems of kearns & Singh[9] but they show it for state value function while we show an upper bound on the error for action-value function estimation defined by $\Delta_{n}^{\hat{Q}_{\eta}}=max_{s,a}\bigg{\{}\bigg{|}\hat{Q}_{\eta,n}(s,a)-Q_{\eta}(s,a)\bigg{|}\bigg{\}}$
 Theorem 3.
 Inspired by Kearns & Singh[9], for any $10<\delta<$ 1, let $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$. with probability $1-\delta$,
-$\Delta_{t}^{\hat{Q}_{\gamma}}\leq\epsilon\bigg{(}\underbrace{\frac{1-\gamma^{k}}{1-\gamma}\bigg{)}}_{\text{variance term}}+\underbrace{\gamma^{k}\Delta_{t-1}^{\hat{Q}_{\gamma}}}_{\text{bias term}}$ (25)
+$\Delta_{n}^{\hat{Q}_{\eta}}\leq\epsilon\bigg{(}\underbrace{\frac{1-\eta^{k}}{1-\eta}\bigg{)}}_{\text{variance term}}+\underbrace{\eta^{k}\Delta_{n-1}^{\hat{Q}_{\eta}}}_{\text{bias term}}$ (28)
 The proof is presented in appendix PROOFS.
 In Eq.
-25, A variance term ($\frac{1-\gamma^{k}}{1-\gamma}$) is arising due to sampling error from rewards collected along trajectories.
+28, A variance term ($\frac{1-\eta^{k}}{1-\eta}$) is arising due to sampling error from rewards collected along trajectories.
 In particular, $\epsilon$ bounds the deviation of the empirical average of rewards from the true expected reward.
-Second term ($\gamma^{k}\Delta_{t-1}^{\hat{Q}_{\gamma}}$) in Eq.
+Second term ($\eta^{k}\Delta_{n-1}^{\hat{Q}_{\eta}}$) in Eq.
 28 is a bias term that arises from bootstrapping off the previous phase’s estimates.
 Similarly, we consider a phased SARSA version of multi-step $TD(\Delta)$.
-For each phase t, we update each W as follows:
-$\hat{W}_{z,t}(s,a)=\frac{1}{n}\sum_{j=1}^{n}\bigg{(}\sum_{i=1}^{k-1}(\gamma_{z}^{i}-\gamma_{z-1}^{i})r_{i}^{(j)}+(\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}})Q_{\gamma_{z-1}}(s_{t+k}^{(j)},a_{t+k}^{(j)})+\gamma_{z}^{k_{z}}\hat{W}_{z}(s_{t+k}^{(j)},a_{t+k}^{(j)})$ (26)
-We now determine the phased error’s upper bound.
-The sum of the errors made by each W component $\sum_{z=0}^{Z}\Delta_{t}^{z}$ is known as $TD(\Delta)$, where $\Delta_{t}^{z}=max_{s,a}\bigg{\{}\bigg{|}\hat{W}_{z}(s,a)-W_{z}(s,a)\bigg{|}\bigg{\}}$.
-Key distinctions between SARSA and Q-learning in terms of the error bound: In SARSA, the error is determined by the agent’s actual actions, which means that the Q-values are updated according to the state-action pairs that the agent visits during exploration.
-The error in Q-learning is derived from the greedy action, which is the action that maximizes the Q-value for the next state, under the assumption that the agent always acts optimally.
+For each phase n, we update each D as follows:
+$\hat{D}_{m,n}(s,a)=\frac{1}{n}\sum_{x=1}^{n}\bigg{(}\sum_{i=1}^{k-1}(\eta_{m}^{i}-\eta_{m-1}^{i})r_{i}^{(x)}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k}^{(x)},a_{n+k}^{(x)})+\eta_{m}^{k_{m}}\hat{D}_{m}(s_{n+k}^{(x)},a_{n+k}^{(x)})$ (29)
+We will now ascertain the highest limit of the phased error.
+The cumulative errors of each D component, denoted as $TD(\Delta)$, is expressed as $\sum_{m=0}^{M}\Delta_{n}^{m}$, where $\Delta_{n}^{m}=\max_{s,a}\bigg{\{}\bigg{|}\hat{D}_{m}(s,a)-D_{m}(s,a)\bigg{|}\bigg{\}}$.
+Fundamental differences between SARSA and Q-learning regarding the error bound: In SARSA, the error is ascertained by the agent’s actual actions, indicating that the Q-values are modified based on the state-action pairings encountered by the agent during exploration.
+The inaccuracy in Q-learning arises from the greedy action, defined as the action that maximizes the Q-value for the subsequent state, based on the premise that the agent consistently behaves optimally.
 Theorem 4.
-Assume that $\gamma_{0}\leq\gamma_{1}\leq\gamma_{2}\leq.....\leq\gamma_{z}=\gamma$ and $K_{0}\leq K_{1}\leq...\leq K_{z}=K$, for any $10<\delta<$ 1, let $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$, with probability $1-\delta$,
-$\sum_{z=0}^{Z}\Delta_{t}^{z}\leq\epsilon\bigg{(}\frac{1-\gamma^{k}}{1-\gamma}\bigg{)}+\epsilon\bigg{(}\underbrace{\sum_{z=0}^{Z-1}\frac{\gamma_{z}^{k_{z+1}}-\gamma^{k_{z}}_{z}}{1-\gamma_{z}}\bigg{)}}_{\text{variance reduction}}+\underbrace{\sum_{z=0}^{Z-1}\bigg{(}\gamma^{k_{z}}_{z}-\gamma^{k_{z+1}}_{z}\bigg{)}\sum_{u=0}^{z}\Delta_{t-1}^{u}}_{\text{bias introduction}}+\gamma^{k}\sum_{z=0}^{Z}\Delta_{t-1}^{z}$ (27)
+Inspired by Romoff et al.[8], assume that $\eta_{0}\leq\eta_{1}\leq\eta_{2}\leq.....\leq\eta_{m}=\eta$ and $K_{0}\leq K_{1}\leq...\leq K_{m}=K$, for any $10<\delta<$ 1, let $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$, with probability $1-\delta$,
+$\sum_{m=0}^{M}\Delta_{n}^{m}\leq\epsilon\bigg{(}\frac{1-\eta^{k}}{1-\eta}\bigg{)}+\epsilon\bigg{(}\underbrace{\sum_{m=0}^{M-1}\frac{\eta_{m}^{k_{m+1}}-\eta^{k_{m}}_{m}}{1-\eta_{m}}\bigg{)}}_{\text{variance reduction}}+\underbrace{\sum_{m=0}^{M-1}\bigg{(}\eta^{k_{m}}_{m}-\eta^{k_{m+1}}_{m}\bigg{)}\sum_{q=0}^{m}\Delta_{n-1}^{q}}_{\text{bias introduction}}+\eta^{k}\sum_{m=0}^{M}\Delta_{n-1}^{m}$ (30)
 The proof is presented in appendix PROOFS.
 The author examines the comparison of the limits for phased TD($\lambda$) in theorem 3 with those for phased TD($\Delta$) in theorem 4.
-This comparison demonstrates that phased TD($\Delta$) facilitates variance reduction equivalent to $\epsilon\sum_{z=0}^{Z-1}\frac{\gamma_{z}^{k_{z+1}}-\gamma^{k_{z}}_{z}}{1-\gamma_{z}}\leq 0$ but introduces a potential bias quantified as $\sum_{z=0}^{Z-1}\bigg{(}\gamma^{k_{z}}_{z}-\gamma^{k_{z+1}}_{z}\bigg{)}\sum_{u=0}^{z}\Delta_{t-1}^{u}\geq 0$.
-Utilizing phased TD($\Delta$) to decrease $k_{z}$ synchronizes updates more effectively with recent actions, hence lowering variance from high-discounted return components.
-Shortened $k_{z}$ values accrue bias more swiftly, as SARSA bootstraps from current policy estimations that may vary over phases.
-It is observed that when all $k_{z}$ values are identical, both algorithms yield the same upper bound.
+This comparison demonstrates that phased TD($\Delta$) facilitates variance reduction equivalent to $\epsilon\sum_{m=0}^{M-1}\frac{\eta_{m}^{k_{m+1}}-\eta^{k_{m}}_{m}}{1-\eta_{m}}\leq 0$ but introduces a potential bias quantified as $\sum_{m=0}^{M-1}\bigg{(}\eta^{k_{m}}_{m}-\eta^{k_{m+1}}_{m}\bigg{)}\sum_{q=0}^{m}\Delta_{n-1}^{q}\geq 0$.
+Utilizing phased TD($\Delta$) to decrease $k_{m}$ synchronizes updates more effectively with recent actions, hence lowering variance from high-discounted return components.
+Shortened $k_{m}$ values accrue bias more swiftly, as SARSA bootstraps from current policy estimations that may vary over phases.
+It is observed that when all $k_{m}$ values are identical, both algorithms yield the same upper bound.
 In SARSA, the return estimate encompasses subsequent states and actions dictated by the policy.
-The expected discounted return over T steps closely approximates the infinite-horizon discounted expected return after $\mathrm{T}$, where $\mathrm{T}\approx\frac{1}{1-\gamma}$[31].
-Consequently, we can effectively simplify $k_{z}$ for any $\gamma_{z}$ such that $k_{z}\approx\frac{1}{1-\gamma_{z}}$, thereby adhering to this principle.
-Therefore, with $\mathrm{T}$ samples, similar to TD($\Delta$)[8], SARSA employing TD($\Delta$) utilizes $k_{z}\approx\frac{1}{1-\gamma_{z}}$ to establish a suitable balance between bias and variance over all time scales significantly less than $\mathrm{T}$.
-By establishing $k_{z}$ for each $\gamma_{z}$, where $\gamma_{z}^{\frac{1}{1-\gamma_{z}}}\leq\frac{1}{e}$, the approach guarantees that the effective horizon and variance are maintained within limits without requiring extensive parameter optimization.
-Each increase in $W_{z}$ facilitates the doubling of the effective horizon, hence ensuring logarithmic scaling with the quantity of action-value functions.
+The expected discounted return over T steps closely approximates the infinite-horizon discounted expected return after $\mathrm{T}$, where $\mathrm{T}\approx\frac{1}{1-\eta}$[31].
+Consequently, we can effectively simplify $k_{m}$ for any $\eta_{m}$ such that $k_{m}\approx\frac{1}{1-\eta_{m}}$, thereby adhering to this principle.
+Therefore, with $\mathrm{T}$ samples, similar to TD($\Delta$)[8], SARSA employing TD($\Delta$) utilizes $k_{m}\approx\frac{1}{1-\eta_{m}}$ to establish a suitable balance between bias and variance over all time scales significantly less than $\mathrm{T}$.
+By establishing $k_{m}$ for each $\eta_{m}$, where $\eta_{m}^{\frac{1}{1-\eta_{m}}}\leq\frac{1}{e}$, the approach guarantees that the effective horizon and variance are maintained within limits without requiring extensive parameter optimization.
+Each increase in $D_{m}$ facilitates the doubling of the effective horizon, hence ensuring logarithmic scaling with the quantity of action-value functions.
 SARSA can utilize this to modify additional parameters in alignment with the specified time scales, enhancing the algorithm’s efficiency in calculating long-term rewards without added complexity.
-6 Experiments
-We illustrate the efficacy of SARSA($\Delta$) via tests on multiple RL benchmarks, including the Ring MDP[9] and dense reward Atari environments.
-The findings indicate that SARSA($\Delta$) attains more rapid convergence and superior long-term efficacy relative to conventional SARSA and TD learning.
-All hyperparameter settings, extended details, and the reproducibility checklist for machine learning research[32] are included in the supplemental materials of this paper[8].
+6 Experimental Setup
+We experimented with OpenAI Gym control tasks, and Fig.1 shows the relevant environments.
+The experiments were carried out on a single PC with 16 GB of memory, an Intel Core i7-2600 processor, and a GPU.
+The system ran 64-bit Ubuntu 22.04.5 LTS.
+We utilized Python 3.9 and the PyTorch package for development.
+We also used the OpenAI Gym toolkit, as mentioned by G.
+Brockman et al.
+[brockman2016openai].
+Ten randomly selected seeds were used in each experiment, and the average outcomes are presented.
+(a) FrozenLake-v1
+(b) Taxi-v3
+(c) Blackjack-v1
+(d) CliffWalking-v0
+Figure 1: OpenAI Gym control used for all experiments. In order from the left to right: Frozen Lake, Taxi, Blackjack and Cliff Walking.
+6.1 Tabular
+We utilize Q-value iteration to determine the optimal Q-value function.
+We employ an open-source version of Value Iteration (VI) derived from the code of Peter Henderson and Wei-Di Chang[value-iteration], guided by the algorithmic principles outlined in Bellman[22] and Sutton & Barto[29].
+Although their methodology calculates the optimal value function, we modify it to determine the optimal Q-value function.
+Upon acquiring the optimal Q-values, we employ Temporal Difference (TD) and SARSA($\Delta$) to derive approximation Q-values via sampling.
+After 5000 steps, each learning iteration comes to an end.
+For every setup, we run 200 trials using different random seeds that influence the dynamics of the transition.
+We implement our algorithms using the code that Kostrikov provided[pytorchrl].
+We compare our approach to k-step SARSA, as described in Sutton and Barto’s book, Chapter 7.2[29].
+We evaluate SARSA($\Delta$) on multiple RL benchmarks, such as the Ring MDP[9] and deterministic & stochastic environments, to determine its effectiveness.
+A illustration of the Ring MDP is shown in Fig.
+2, which shows a Markov process with five states arranged in a circle.
+There is a 0.95 chance of moving one state clockwise around the ring and 0.05 chance of staying in the present state at each step.
+While the other states on the ring offer a reward of 0, two adjacent states offer rewards of +1 and -1, respectively.
+Figure 2: Five-state ring MDP as described by Kearns[9]
+All experiments utilize the identical parameters established by Romoff et al.[8], which involve supplying a variable number of gamma values, commencing at 0 and incrementing according to $\eta_{m+1}=\frac{\eta_{m}+1}{2}$ until the requisite maximum $\eta_{m}$ is attained.
+As previously mentioned in Section 5.2, $K_{m}=\frac{1}{(1-\eta_{m})}$ for every m.
+The baseline consists of a singular estimator with $\eta=\eta_{m}$ and $k=k_{m}$.
+Fig.
+3(a) depicts the efficacy of two methodologies: K-step SARSA ($\Delta$) and K-step SARSA TD, employing a discount factor ($\eta$ = 0.996), k = 64 (a parameter associated with the algorithm’s step count), and diverse values of the learning rate $\alpha$.
+The y-axis denotes the absolute error, whereas the x-axis signifies the parameter $\alpha$, ranging from 0.0 to 1.0.
+At $\alpha$=0.0, both methodologies commence at an equivalent error level (about 0.00230).
+As $\alpha$ increases, K-step SARSA ($\Delta$) initially diminishes its inaccuracy marginally before seeing a consistent rise.
+K-step SARSA TD, conversely, exhibits a consistent increase in error with minimal fluctuation.
+Overall course: Relative to K-step SARSA TD, the K-step SARSA ($\Delta$) technique has greater stability and consistently yields superior outcomes (i.e., reduced error) across all values of $\alpha$.
+Both approaches exhibit sensitivity to variations in $\alpha$, although K-step SARSA TD demonstrates greater sensitivity, evidenced by its more pronounced increase in error.
+Fig.
+3(b) demonstrates the effectiveness of the two methodologies, K-step SARSA ($\Delta$) and K-step SARSA TD, in relation to the number of k-steps.
+The y-axis indicates the absolute error, whereas the x-axis represents k-steps, which refers to the number of steps utilized in the algorithms.
+The range of k-steps extends from 0 to roughly 250.
+K-step SARSA ($\Delta$) commences with a reduced error at smaller k-steps, which subsequently increases with bigger k-steps.
+K-step SARSA TD begins with an error level comparable to K-step SARSA ($\Delta$) for low k-steps; however, its error escalates more steeply as k-steps grow.
+The overall pattern indicates that both techniques exhibit escalating inaccuracy as k-steps rise.
+K-step SARSA ($\Delta$) consistently surpasses K-step SARSA TD in all k-step values by exhibiting a reduced error rate.
+The swift error amplification in K-step SARSA TD indicates that the scheduling method employed in K-step SARSA ($\Delta$) mitigates error accumulation.
+Furthermore, in this specific MDP, the error escalates with k due to the bias-variance trade-off associated with k-step returns, aligning with the conclusions of Kearns & Singh[9, 8].
+Comparison of Figs.
+3(a) and 3(b): In the comparison of the two figures, K-step SARSA ($\Delta$) consistently surpasses K-step SARSA TD in both circumstances ($\alpha$ and k-steps).
+Both approaches exhibit sensitivity to elevated parameter values ($\alpha$ or k-steps).
+K-step SARSA ($\Delta$) exhibits enhanced stability and attains reduced errors as a result of its scheduling approach.
+Significantly, errors escalate more rapidly with K-step SARSA TD as k-steps grow in comparison to the increase in $\alpha$.
+(a) Error at different learning rates
+(b) Error at different k values
+Figure 3: Fig.3(a) illustrates the outcomes for $\eta_{M}=0.996$ on the five-state Ring MDP.
+The error signifies the absolute discrepancy between the estimated and actual discounted Q-value function (precomputed via Q-value iteration), averaged across the complete learning trajectory of 5000 timesteps.
+Fig.3(b) illustrates the average absolute error for the optimal learning rate at each k-step return, extending beyond the effective planning horizon of $\eta_{M}$.
+6.2 OpenAI Gym Environments
+The following plots show the results of two algorithms, k-step SARSA($\Delta$) and k-step SARSA TD, on the FrozenLake-v1, Taxi-v3, CliffWalking-v0, and Blackjack environments from OpenAI Gym.
+The y-axis shows the average rewards obtained by the algorithms over episodes, and the x-axis shows the number of episodes needed to train the algorithms.
+Episodes range from 0 to 100,000 episodes.
+The shaded area surrounding each curve shows the standard error over 10 random seeds.
+The algorithms both use k= 32, which is the number of steps.
+The remaining hyperparameter settings are displayed in the table 1.
+Table 1: Hyperparameter configurations and their corresponding values.
+Hyperparameter Value
+seed (10) np.random.randint(40, 45)
+Learning Rate ($\alpha$) [0.1, 0.3, 0.5, 0.7, 1.0]
+Discount Factor ($\eta$) [0.992, 0.95, 0.996, 0.99, 1.0]
+k-Step [1, 4, 8, 16, 32, 128, 256]
+Number of Episodes 100000
+Max Epsilon 0.4
+The red line denotes K-step SARSA ($\Delta$), while the blue line represents K-step SARSA TD in both Fig.
+4(a) and Fig.
+4(b).
+Fig.
+4(a) illustrates the FrozenLake-v1 environment, whereas Fig.
+4(b) depicts the Taxi-v3 environment.
+Fig.
+4(a) demonstrates that k-step SARSA ($\Delta$) continually attains superior average rewards relative to k-SARSA TD as training advances.
+K-step SARSA ($\Delta$) demonstrates considerable enhancement in average rewards over the initial 20,000 episodes, signifying accelerated learning, but k-step SARSA TD exhibits lesser increase in average rewards in comparison to k-step SARSA ($\Delta$) during the same episodes.
+Following 40,000 episodes, the efficacy of k-step SARSA($\Delta$) stabilizes while consistently surpassing that of k-step SARSA TD, indicating enduring reliability.
+k-step SARSA regularly exhibits inferior performance relative to k-step SARSA ($\Delta$).
+Fig.
+4(b) illustrates that K-Step SARSA ($\Delta$) regularly achieves superior rewards (i.e., nearer to -200) during the training phase, signifying enhanced learning efficiency and overall performance.
+Fig.
+4(b) also demonstrates that K-Step SARSA ($\Delta$) stabilizes in the initial phases, indicating robustness and the capacity to converge to an effective policy, whereas K-Step SARSA TD commences at considerably lower rewards (approximately -1800) and exhibits negligible improvement, remaining predominantly flat throughout the training process.
+The K-Step SARSA TD algorithm encounters difficulties in acquiring an effective policy in this setting compared to K-Step SARSA ($\Delta$).
+Experiment result show that K-Step SARSA ($\Delta$) significantly outperforms K-Step SARSA TD in the Taxi-v3 environment, achieving much higher average rewards and faster convergence.
+(a) FrozenLake-v1
+(b) Taxi-v3
+Figure 4: Comparative analysis of K-step SARSA ($\Delta$) versus traditional K-step SARSA TD within the OpenAI Gym environments: FrozenLake-v1 and Taxi-v3.
+The shaded area illustrates the standard error calculated from 10 random seeds.
+(a) CliffWalking-v0
+(b) Blackjack-v1
+Figure 5: Comparative analysis of K-step SARSA ($\Delta$) versus traditional K-step SARSA TD within the OpenAI Gym environments: CliffWalking-v0 and Blackjack-v1.
+The shaded area illustrates the standard error calculated from 10 random seeds.
+The Fig.
+5(a) and 5(b) illustrates the learning effectiveness of two variations of the K-Step SARSA algorithm, namely K-Step SARSA ($\Delta$) and K-Step SARSA TD, across 100,000 episodes in the CliffWalking-v0 and Blackjack-v1 environments.
+Both algorithms are assessed using k=32, a hyperparameter that specifies the number of steps in the updating process.
+The trend of average rewards in fig.
+5(a) demonstrates that both algorithms show an increasing pattern in average rewards as the number of episodes rises, indicating improved policy performance with continued learning.
+Nevertheless, K-Step SARSA ($\Delta$) often outperforms K-Step SARSA TD, achieving higher average rewards over the episodes.
+The red line in fig.
+5(a) denoting K-Step SARSA ($\Delta$) demonstrates a considerable enhancement in the initial episodes (about 20,000 episodes), indicating a swifter convergence to optimal policies relative to K-Step SARSA TD.
+In the 100,000th episode, K-Step SARSA ($\Delta$) attains a reward near -1050, whereas K-Step SARSA TD obtains a slightly lower payout of approximately -1120.
+This demonstrates the enhanced efficacy of K-Step SARSA ($\Delta$) in this task.
+The average reward of approximately +0.15 is dependably achieved across all episodes by the red curve in fig.
+5(b) that represents K-Step SARSA ($\Delta$).
+The blue curve in fig.
+5(b), which represents K-Step SARSA TD, suggests stability; however, it presents an average reward that is significantly lower, at approximately -0.2, across all episodes.
+K-Step SARSA ($\Delta$) markedly surpasses K-Step SARSA TD regarding average rewards.
+The favorable average reward of K-Step SARSA ($\Delta$) demonstrates its superior suitability for policy optimization in the Blackjack-v1 environment, whereas the negative reward for K-Step SARSA TD signifies inferior performance.
+Finally, when comparing the results of all four environments, the findings demonstrate the constant advantage of K-Step SARSA ($\Delta$) over K-Step SARSA TD across all evaluated environments.
+This superiority is seen through elevated average rewards, accelerated convergence, and enhanced stability in learning outcomes.
+Environments characterized by deterministic state transitions, such as CliffWalking-v0 and FrozenLake-v1, demonstrate the clear superiority of K-Step SARSA ($\Delta$), but intricate stochastic environments like Blackjack-v1 and Taxi-v3 further highlight its resilience.
+The performance graph of deterministic environments, such as CliffWalking-v0 and FrozenLake-v1, resembles an exponential growth plateau curve due to predictable state transitions.
+In contrast, the performance graph of complex stochastic environments, like Blackjack-v1 and Taxi-v3, appears as a flat line because of the uncertainty and variability in state transitions.
 7 Conclusion
-SARSA($\Delta$) expands SARSA by including multiple discount factors through TD($\Delta$), resulting in improved performance and stability.
-This approach for a more precise breakdown of future rewards while maintaining the benefits of on-policy learning.
-The Q-Learning version of TD($\Delta$) is a simple extension of this.
-We put it off for later.
+With the help of TD($\Delta$), SARSA($\Delta$) improves SARSA by integrating multiple discount factors, leading to better stability and performance.
+With this approach, we can get a more precise breakdown of action-values while still enjoying the benefits of on-policy learning.
+In addition to the deterministic and stochastic OpenAI Gym environments, we also conducted trials in a tabular synthetic environment, specifically the five-star ring MDP.
+In all investigated environments, our experimental results demonstrate that SARSA($\Delta$), our suggested approach, outperforms k-step SARSA TD in terms of greater average rewards, faster convergence, and more stable learning trajectories.
+The results of this work advance our understanding of multi-step RL algorithms and how they are applied to various sequential decision-making tasks.
+This information can aid in the creation of robust algorithms for unpredictable, real-world environments in the future.
+In essence, the Q-Learning version of TD($\Delta$) extends this technique.
+We decided to take care of it later.
 Acknowledgments
@@
 Data availability
-The data sets produced in this work can be obtained from the corresponding author upon an appropriate request.
 Competing interests
@@
 The proof utilizes induction.
-Initially, it is necessary to establish that the base case is valid at t=0.
+Initially, it is necessary to establish that the base case is valid at n=0.
 This statement is evidently valid given our initial assumption, particularly in the context of zero initialization.
-At a specific time-step t, we assume the validity of the statement, namely, $\sum_{z=0}^{Z}\theta_{t}^{z}=\theta_{t}^{\gamma}$.
-We will now demonstrate that it also holds at the subsequent time step, t+1.
-$\sum_{z=0}^{Z}\theta_{t+1}^{z}$ $=\sum_{z=0}^{Z}\bigg{(}\theta_{t}^{z}+\alpha_{z}\bigg{(}G_{t}^{Z,\lambda_{z}}-\hat{W}_{z}(s_{t},a_{t})\bigg{)}\phi(s_{t},a_{t})\bigg{)}$
+At a specific time-step n, we assume the validity of the statement, namely, $\sum_{m=0}^{M}\theta_{n}^{m}=\theta_{n}^{\eta}$.
+We will now demonstrate that it also holds at the subsequent time step, n+1.
+$\sum_{m=0}^{M}\theta_{n+1}^{m}$ $=\sum_{m=0}^{M}\bigg{(}\theta_{n}^{m}+\alpha_{m}\bigg{(}G_{n}^{M,\lambda_{m}}-\hat{D}_{m}(s_{n},a_{n})\bigg{)}\phi(s_{n},a_{n})\bigg{)}$
 […]
-$=\theta_{t}^{\gamma}+\sum_{z=0}^{Z}\alpha_{z}\bigg{(}\sum_{k=t}^{\infty}(\lambda_{z}\gamma_{z})^{k-t}\delta_{k}^{z}\bigg{)}\phi(s_{t},a_{t})\text{ , Assumption based on induction.}$.
-$=\theta_{t}^{\gamma}+\alpha\sum_{k=t}^{\infty}(\lambda\gamma)^{k-t}\underbrace{\bigg{(}\sum_{z=0}^{Z}\delta_{k}^{z}\bigg{)}}_{*}\phi(s_{t},a_{t})\text{ , Because of $\alpha_{z}=\alpha$, $\lambda_{z}\gamma_{z}=\lambda_{\gamma}$, $\forall z$}$
-To show that $\sum_{z=0}^{Z}\theta_{t+1}^{z}=\theta_{t+1}^{\gamma}$, we must establish that the term (*) = $\sum_{z=0}^{Z}\delta_{k}^{z}$ is equivalent to the conventional TD error $\delta_{k}^{\gamma}$.
-$\sum_{z=0}^{Z}\delta_{k}^{z}$ $=r_{k}+\gamma_{0}\hat{Q}_{\gamma_{0}}(s_{k+1},a_{k+1})-\hat{Q}_{\gamma_{0}}(s_{k},a_{k})+\sum_{z=1}^{Z}\bigg{(}(\gamma_{z}-\gamma_{z-1})\sum_{u=0}^{z-1}\langle\theta_{t}^{u},\phi(s_{k+1},a_{k+1})\rangle+\gamma_{z}\langle\theta_{t}^{z},\phi(s_{k+1},a_{k+1})\rangle-\langle\theta_{t}^{z},\phi(s_{k},a_{k})\rangle\bigg{)}$
-$=r_{k}+\gamma_{0}\hat{Q}_{\gamma_{0}}(s_{k+1},a_{k+1})+\bigg{\langle}\sum_{z=1}^{Z}\bigg{(}\gamma_{z}\sum_{u=0}^{z}\theta_{t}^{u}-\gamma_{z-1}\sum_{u=0}^{z-1}\theta_{t}^{u}\bigg{)},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-\bigg{\langle}\sum_{z=0}^{Z}\theta_{t}^{z},\phi(s_{k},a_{k})\bigg{\rangle}$
-$=r_{k}+\gamma_{0}\hat{Q}_{\gamma_{0}}(s_{k+1},a_{k+1})+\gamma_{z}\bigg{\langle}\sum_{z=0}^{Z}\theta_{t}^{z},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-\gamma_{0}\bigg{\langle}\theta_{t}^{0},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-Q_{\gamma}(s_{k},a_{k})$
-$=r_{k}+\gamma_{0}\hat{Q}_{\gamma_{0}}(s_{k+1},a_{k+1})+\gamma\hat{Q}_{\gamma}(s_{k+1},a_{k+1})-\gamma_{0}\hat{Q}_{\gamma_{0}}(s_{k+1},a_{k+1})-\hat{Q}_{\gamma}(s_{k},a_{k})$
-$=r_{k}+\gamma\hat{Q}_{\gamma}(s_{k+1},a_{k+1})-\hat{Q}_{\gamma}(s_{k},a_{k})$
-$=\delta_{k}^{\gamma}$ (29)
+$=\theta_{n}^{\eta}+\sum_{m=0}^{M}\alpha_{m}\bigg{(}\sum_{k=n}^{\infty}(\lambda_{m}\eta_{m})^{k-n}\delta_{k}^{m}\bigg{)}\phi(s_{n},a_{n})\text{ , Assumption based on induction.}$.
+$=\theta_{n}^{\eta}+\alpha\sum_{k=n}^{\infty}(\lambda\eta)^{k-n}\underbrace{\bigg{(}\sum_{m=0}^{M}\delta_{k}^{m}\bigg{)}}_{*}\phi(s_{n},a_{n})\text{ , Because of $\alpha_{m}=\alpha$, $\lambda_{m}\eta_{m}=\lambda_{\eta}$,$\forall m$}$
+To show that $\sum_{m=0}^{M}\theta_{n+1}^{m}=\theta_{n+1}^{\eta}$, we must establish that the term (*) = $\sum_{m=0}^{M}\delta_{k}^{m}$ is equivalent to the conventional TD error $\delta_{k}^{\eta}$.
+$\sum_{m=0}^{M}\delta_{k}^{m}$ $=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})-\hat{Q}_{\eta_{0}}(s_{k},a_{k})+\sum_{m=1}^{M}\bigg{(}(\eta_{m}-\eta_{m-1})\sum_{q=0}^{m-1}\langle\theta_{n}^{q},\phi(s_{k+1},a_{k+1})\rangle+\eta_{m}\langle\theta_{n}^{m},\phi(s_{k+1},a_{k+1})\rangle-\langle\theta_{n}^{m},\phi(s_{k},a_{k})\rangle\bigg{)}$
+$=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})+\bigg{\langle}\sum_{m=1}^{M}\bigg{(}\eta_{m}\sum_{q=0}^{m}\theta_{n}^{q}-\eta_{m-1}\sum_{q=0}^{m-1}\theta_{n}^{q}\bigg{)},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-\bigg{\langle}\sum_{m=0}^{M}\theta_{n}^{m},\phi(s_{k},a_{k})\bigg{\rangle}$
+$=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})+\eta_{m}\bigg{\langle}\sum_{m=0}^{M}\theta_{n}^{m},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-\eta_{0}\bigg{\langle}\theta_{n}^{0},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-Q_{\eta}(s_{k},a_{k})$
+$=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})+\eta\hat{Q}_{\eta}(s_{k+1},a_{k+1})-\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})-\hat{Q}_{\eta}(s_{k},a_{k})$
+$=r_{k}+\eta\hat{Q}_{\eta}(s_{k+1},a_{k+1})-\hat{Q}_{\eta}(s_{k},a_{k})$
+$=\delta_{k}^{\eta}$ (32)
 Proof: Theorem 2
 The following is the definition of the Bellman operator:
-$\mathrm{T}=r+\gamma P$
-The discount factor, the expected reward function, and the transition probability operator that is induced by the policy $\pi$ are denoted by the symbols $\gamma$, r, and P, respectively.
+$\mathrm{T}=r+\eta P$
+The discount factor, the expected reward function, and the transition probability operator that is induced by the policy $\pi$ are denoted by the symbols $\eta$, r, and P, respectively.
 A geometric weighted sum of $\mathrm{T}$ is the definition of the TD($\lambda$) operator, which can be expressed as follows:
 $\mathrm{T}=(1-\lambda)\sum_{k=0}^{\infty}\lambda^{k}(\mathrm{T})^{k+1}$ (34)
-It may be inferred that $\lambda\in[0,1]$ is necessary for the aforementioned sum to be finite. A corresponding definition of $\mathrm{T}_{\lambda}$ is as follows for every function W:
-$\mathrm{T}_{\lambda}W=W+(I-\lambda\gamma P)^{-1}(\mathrm{T}W-W)$ (32)
-The separated action-value function is denoted by W in this form, and the update is given in terms of the operator $\mathrm{T}_{\lambda}$, which is responsible for applying time-scale separation.
-In the event that $10\leq\lambda\gamma<$ 1, the formula is considered to be well-defined.
-This is because the spectral norm of the operator $\lambda\gamma P$ is less than 1, and thus, the Eq.
-$I-\lambda\gamma P$ is invertible.
+It may be inferred that $\lambda\in[0,1]$ is necessary for the aforementioned sum to be finite. A corresponding definition of $\mathrm{T}_{\lambda}$ is as follows for every function D:
+$\mathrm{T}_{\lambda}D=D+(I-\lambda\eta P)^{-1}(\mathrm{T}D-D)$ (35)
+The separated action-value function is denoted by D in this form, and the update is given in terms of the operator $\mathrm{T}_{\lambda}$, which is responsible for applying time-scale separation.
+In the event that $10\leq\lambda\eta<$ 1, the formula is considered to be well-defined.
+This is because the spectral norm of the operator $\lambda\eta P$ is less than 1, and thus, the Eq.
+$I-\lambda\eta P$ is invertible.
 On the other hand, when the value of $\lambda$ is bigger than one, the equivalence between Eqs.
-31 and 32 is lost.
+34 and 35 is lost.
 This is not a problem because, in fact, we use the TD error for training, which, according to our expectations, corresponds to the definition of $\mathrm{T}_{\lambda}$ that is provided in Eq.
-32.
+35.
 Now, let’s have a look at the contraction property that the operator $\mathrm{T}_{\lambda}$ possesses.
 To begin, the Eq.
-designated as 32 can be rewritten as follows:
-$\mathrm{T}_{\lambda}$ $=(I-\lambda\gamma P)^{-1}(\mathrm{T}W-\lambda\gamma PW)$ (33)
-The contraction property is demonstrated by examining two action-value functions, $W_{1}$ and $W_{2}$.
-The distinction between the updates of $\mathrm{T}_{\lambda}W_{1}$ and $\mathrm{T}_{\lambda}W_{2}$ is:
-$\mathrm{T}_{\lambda}W_{1}-\mathrm{T}_{\lambda}W_{2}$ $=(I-\lambda\gamma P)^{-1}\bigg{(}\mathrm{T}W_{1}-\mathrm{T}W_{2}-\lambda\gamma P(W_{1}-W_{2})\bigg{)}$
-$=(I-\lambda\gamma P)^{-1}\bigg{(}\gamma P(W_{1}-W_{2})-\lambda\gamma P(W_{1}-W_{2})\bigg{)}$
-$=(I-\lambda\gamma P)^{-1}\bigg{(}\gamma(1-\lambda)P(W_{1}-W_{2})\bigg{)}$
+designated as 35 can be rewritten as follows:
+$\mathrm{T}_{\lambda}$ $=(I-\lambda\eta P)^{-1}(\mathrm{T}D-\lambda\eta PD)$ (36)
+The contraction property is demonstrated by examining two action-value functions, $D_{1}$ and $D_{2}$.
+The distinction between the updates of $\mathrm{T}_{\lambda}D_{1}$ and $\mathrm{T}_{\lambda}D_{2}$ is:
+$\mathrm{T}_{\lambda}D_{1}-\mathrm{T}_{\lambda}D_{2}$ $=(I-\lambda\eta P)^{-1}\bigg{(}\mathrm{T}D_{1}-\mathrm{T}D_{2}-\lambda\eta P(D_{1}-D_{2})\bigg{)}$
+$=(I-\lambda\eta P)^{-1}\bigg{(}\eta P(D_{1}-D_{2})-\lambda\eta P(D_{1}-D_{2})\bigg{)}$
+$=(I-\lambda\eta P)^{-1}\bigg{(}\eta(1-\lambda)P(D_{1}-D_{2})\bigg{)}$
 We derive the following:
-$\|\mathrm{T}_{\lambda}W_{1}-\mathrm{T}_{\lambda}W_{2}\|\leq\frac{\gamma|1-\lambda|}{1-\lambda\gamma}\|W_{1}-W_{2}\|\textit{ where P = 1}$
+$\|\mathrm{T}_{\lambda}D_{1}-\mathrm{T}_{\lambda}D_{2}\|\leq\frac{\eta|1-\lambda|}{1-\lambda\eta}\|D_{1}-D_{2}\|\textit{ where P = 1}$
 It is established that $0\leq\lambda\leq 0$ constitutes a contraction. For $\lambda>1$, the following condition must be satisfied:
-$\frac{\gamma(\lambda-1)}{1-\lambda\gamma}<1\Rightarrow\gamma\lambda-\gamma<1-\lambda\gamma$
-$\Rightarrow 2\lambda\gamma<1+\gamma\Rightarrow\lambda<\frac{1+\gamma}{2\gamma}$ (34)
-Consequently, $\mathrm{T}_{\lambda}$ constitutes a contraction if $0\leq\lambda<\frac{1+\gamma}{2\gamma}$. This directly indicates that for $\gamma<1$, $\gamma\lambda<1$.
+$\frac{\eta(\lambda-1)}{1-\lambda\eta}<1\Rightarrow\eta\lambda-\eta<1-\lambda\eta$
+$\Rightarrow 2\lambda\eta<1+\eta\Rightarrow\lambda<\frac{1+\eta}{2\eta}$ (37)
+Consequently, $\mathrm{T}_{\lambda}$ constitutes a contraction if $0\leq\lambda<\frac{1+\eta}{2\eta}$. This directly indicates that for $\eta<1$, $\eta\lambda<1$.
 Proof: Theorem 3
 […]
-$Q^{\pi}(s,a)=\mathbf{E}\bigg{[}\gamma_{0}+\gamma r_{1}+...+\gamma^{k-1}r_{k-1}+\gamma^{k}Q^{\pi}(s_{k},a_{k})\bigg{]}$
+$Q^{\pi}(s,a)=\mathbf{E}\bigg{[}\eta_{0}+\eta r_{1}+...+\eta^{k-1}r_{k-1}+\eta^{k}Q^{\pi}(s_{k},a_{k})\bigg{]}$
 Here the expectations are over a random trajectory under $\pi$; thus $\mathbf{E}[r_{\iota}](\iota\leq k\mathit{-}1)$ denotes the expected value of the $\iota$th reward received, while $\mathbf{E}[Q^{\pi}(s_{k},a_{k})]$ is the expected value of the true value function at the $k$th state-action reached.
-The phased TD(k) update sums the terms $\gamma^{\iota}\bigg{(}\frac{1}{n}\bigg{)}\sum_{i=1}^{n}r_{\iota}^{i}$, whose expectations are exactly the $\gamma^{\iota}\mathbf{E}[r_{\iota}]$ appearing above [9].
+The phased TD(k) update sums the terms $\eta^{\iota}\bigg{(}\frac{1}{n}\bigg{)}\sum_{i=1}^{n}r_{\iota}^{i}$, whose expectations are exactly the $\eta^{\iota}\mathbf{E}[r_{\iota}]$ appearing above [9].
 […]
-$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{j=1}^{n}r_{i}^{(j)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\geq\epsilon\bigg{)}\leq 2e^{\bigg{(}\mathit{-}\frac{2n^{{}^{2}}\epsilon^{2}}{n2^{2}}\bigg{)}}$ (35)
+$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{x=1}^{n}r_{i}^{(x)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\geq\epsilon\bigg{)}\leq 2e^{\bigg{(}\mathit{-}\frac{2n^{{}^{2}}\epsilon^{2}}{n2^{2}}\bigg{)}}$ (38)
 If we assume n and the probability of exceeding an $\epsilon$ value is fixed to be atmost $\delta$, we can solve for the resulting value of $\epsilon$:
@@
 So by Hoeffding, if we have n samples, with probability at least $1\mathit{-}\delta$,
-$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{j=1}^{n}r_{i}^{(j)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\bigg{)}\leq\epsilon=\sqrt{\frac{2log\frac{2}{\delta}}{n}}$ (37)
+$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{x=1}^{n}r_{i}^{(x)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\bigg{)}\leq\epsilon=\sqrt{\frac{2log\frac{2}{\delta}}{n}}$ (40)
 Since we are dealing with k different state-action pairs, we use a union bound. To ensure the probability holds for all k hypotheses, adjust $\delta$ to $\delta/k$:
-$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{j=1}^{n}r_{i}^{(j)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\bigg{)}\leq\epsilon=\sqrt{\frac{2log\frac{2k}{\delta}}{n}}$
+$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{x=1}^{n}r_{i}^{(x)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\bigg{)}\leq\epsilon=\sqrt{\frac{2log\frac{2k}{\delta}}{n}}$
 We can now assume that all $\mathbb{E}[r_{i}]$ terms are estimated to at least $\epsilon$ accuracy. Substituting this back into the definition of the k-step TD update we get
-$\hat{Q}_{t+1}(s,a)\mathit{-}Q(s,a)=\frac{1}{n}\sum_{j=1}^{n}\bigg{(}r_{0}+\gamma r_{1}+...+\gamma^{k-1}r_{k-1}+\gamma^{k}Q_{t}(s_{k},a_{k})\bigg{)}\mathit{-}Q(s,a)$
-$=\sum_{i=0}^{k-1}\gamma^{i}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}r_{i}^{(j)}-\mathbb{E}[r_{i}]\bigg{)}+\gamma^{k}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}Q_{t}(s_{k}^{i},a_{k}^{i})-\mathbb{E}[Q(s_{k},a_{k})]\bigg{)}$ (38)
+$\hat{Q}_{n+1}(s,a)\mathit{-}Q(s,a)=\frac{1}{n}\sum_{x=1}^{n}\bigg{(}r_{0}+\eta r_{1}+...+\eta^{k-1}r_{k-1}+\eta^{k}Q_{n}(s_{k},a_{k})\bigg{)}\mathit{-}Q(s,a)$
+$=\sum_{i=0}^{k-1}\eta^{i}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}r_{i}^{(x)}-\mathbb{E}[r_{i}]\bigg{)}+\eta^{k}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}Q_{n}(s_{k}^{i},a_{k}^{i})-\mathbb{E}[Q(s_{k},a_{k})]\bigg{)}$ (41)
 where in the second line we re-expressed the value in terms of a sum of k rewards. We now upper bounded the difference from $E[r_{i}]$ by $\epsilon$ to get
-$\hat{Q}_{t+1}(s,a)\mathit{-}Q(s,a)\leq\sum_{i=0}^{k-1}\gamma^{l}\epsilon+\gamma^{k}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}Q_{t}(s_{k}^{i},a_{k}^{i})-\mathbb{E}[Q(s_{k},a_{k})]\bigg{)}$
+$\hat{Q}_{n+1}(s,a)\mathit{-}Q(s,a)\leq\sum_{i=0}^{k-1}\eta^{l}\epsilon+\eta^{k}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}Q_{n}(s_{k}^{i},a_{k}^{i})-\mathbb{E}[Q(s_{k},a_{k})]\bigg{)}$
 The variance term arises from the deviation of the empirical average of rewards from the true expected reward:
-$\epsilon\bigg{(}\frac{1-\gamma^{k}}{1-\gamma}\bigg{)}$
-$\leq\epsilon\bigg{(}\frac{1-\gamma^{k}}{1-\gamma}\bigg{)}+\gamma^{k}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}Q_{t}(s_{k}^{i},a_{k}^{i})-\mathbb{E}[Q(s_{k},a_{k})]\bigg{)}$ (39)
+$\epsilon\bigg{(}\frac{1-\eta^{k}}{1-\eta}\bigg{)}$
+$\leq\epsilon\bigg{(}\frac{1-\eta^{k}}{1-\eta}\bigg{)}+\eta^{k}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}Q_{n}(s_{k}^{i},a_{k}^{i})-\mathbb{E}[Q(s_{k},a_{k})]\bigg{)}$ (42)
 The bias term is propagated through bootstrapping:
-$\gamma^{k}\Delta^{\hat{Q}_{\gamma}}_{t-1}$
-and then the second term is bounded by $\Delta^{\hat{Q}_{\gamma}}_{t-1}$ by assumption. Hence, combining these terms, we get the overall bound for SARSA:
-$\Delta^{\hat{Q}_{\gamma}}_{t}\leq\epsilon\bigg{(}\frac{1-\gamma^{k}}{1-\gamma}\bigg{)}+\gamma^{k}\Delta^{\hat{Q}_{\gamma}}_{t-1}$ (40)
+$\eta^{k}\Delta^{\hat{Q}_{\eta}}_{n-1}$
+and then the second term is bounded by $\Delta^{\hat{Q}_{\eta}}_{n-1}$ by assumption. Hence, combining these terms, we get the overall bound for SARSA:
+$\Delta^{\hat{Q}_{\eta}}_{n}\leq\epsilon\bigg{(}\frac{1-\eta^{k}}{1-\eta}\bigg{)}+\eta^{k}\Delta^{\hat{Q}_{\eta}}_{n-1}$ (43)
 Proof: Theorem 4
-Phased $TD(\Delta)$ update rules for $z\geq 1$:
-$\hat{W}_{z,t}(s,a)=\frac{1}{n}\sum_{j-1}^{n}\bigg{(}\sum_{i=1}^{k_{z}-1}(\gamma_{z}^{i}-\gamma_{z-1}^{i})r_{i}^{(j)}+(\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}})\hat{Q}_{\gamma_{z-1}}(s_{k}^{(j)},a_{k}^{(j)})+\gamma_{z}^{k_{z}}\hat{W}_{z}(s_{k}^{(j)},a_{k}^{(j)})\bigg{)}$ (41)
-We know that according to the multi-step update rule 14 for $z\geq 1$:
-$W_{z}(s_{t},a_{t})=\mathbb{E}\bigg{[}\sum_{i=1}^{k_{z}-1}\bigg{(}\gamma_{z}^{i}-\gamma_{z-1}^{i}\bigg{)}r_{i}+\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}Q_{\gamma_{z-1}}(s_{k},a_{k})+\gamma_{z}^{k_{z}}W_{z}(s_{k},a_{k})\bigg{]}$ (42)
-Then, subtracting the two expressions gives for $z\geq 1$:
-$\hat{W}_{z,t}(s,a)-W_{z}(s_{t},a_{t})=\sum_{i=1}^{k_{z}-1}\bigg{(}\gamma_{z}^{i}-\gamma_{z-1}^{i}\bigg{)}\bigg{(}\frac{1}{n}\sum_{j-1}^{n}r_{i}^{(j)}-\mathbb{E}[r_{i}]\bigg{)}+\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\bigg{(}\sum_{u=0}^{z-1}\hat{W}_{u}(s_{k}^{(j)},a_{k}^{(j)})-\mathbb{E}[W_{z}(s_{k},a_{k})]\bigg{)}$
-$+\gamma_{z}^{k_{z}}\bigg{(}W_{z}(s_{k}^{(j)},a_{k}^{(j)})-\mathbb{E}[W_{z}(s_{k},a_{k})]\bigg{)}$ (43)
-Assume that $k_{0}\leq k_{1}\leq...k_{z}=k$, the W estimates share at most $k_{z}=k$ reward terms $\frac{1}{n}\sum_{j-1}^{n}r_{i}^{(j)}$, Using Hoeffding inequality and union bound, we obtain that with probability $1-\delta$, each k empirical average reward $\frac{1}{n}\sum_{j-1}^{n}r_{i}^{(j)}$ deviates from the true expected reward $\mathbb{E}[r_{i}]$ by at most $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$.
-Hence, with probability $1-\delta$, $\forall z\geq 1$, we have:
-$\Delta_{t}^{z}\leq\epsilon\sum_{i=1}^{k_{z}-1}\bigg{(}\gamma_{z}^{i}-\gamma_{z-1}^{i}\bigg{)}+\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\sum_{u=0}^{z-1}\Delta_{t-1}^{u}+\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}$
-$=\epsilon\bigg{(}\frac{1-\gamma_{z}^{k_{z}}}{1-\gamma_{z}}-\frac{1-\gamma_{z-1}^{k_{z}}}{1-\gamma_{z-1}}\bigg{)}+\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\sum_{u=0}^{z-1}\Delta_{t-1}^{u}+\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}$ (44)
-and $\Delta_{t}^{0}\leq\epsilon\frac{1-\gamma_{0}^{k_{0}}}{1-\gamma_{0}}+\gamma_{0}^{k_{0}}\Delta_{t-1}^{0}$
+Phased $TD(\Delta)$ update rules for $m\geq 1$:
+$\hat{D}_{m,n}(s,a)=\frac{1}{n}\sum_{x-1}^{n}\bigg{(}\sum_{i=1}^{k_{m}-1}(\eta_{m}^{i}-\eta_{m-1}^{i})r_{i}^{(x)}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})\hat{Q}_{\eta_{m-1}}(s_{k}^{(x)},a_{k}^{(x)})+\eta_{m}^{k_{m}}\hat{D}_{m}(s_{k}^{(x)},a_{k}^{(x)})\bigg{)}$ (44)
+We know that according to the multi-step update rule 17 for $m\geq 1$:
+$D_{m}(s_{n},a_{n})=\mathbb{E}\bigg{[}\sum_{i=1}^{k_{m}-1}\bigg{(}\eta_{m}^{i}-\eta_{m-1}^{i}\bigg{)}r_{i}+\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}Q_{\eta_{m-1}}(s_{k},a_{k})+\eta_{m}^{k_{m}}D_{m}(s_{k},a_{k})\bigg{]}$ (45)
+Then, subtracting the two expressions gives for $m\geq 1$:
+$\hat{D}_{m,n}(s,a)-D_{m}(s_{n},a_{n})=\sum_{i=1}^{k_{m}-1}\bigg{(}\eta_{m}^{i}-\eta_{m-1}^{i}\bigg{)}\bigg{(}\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}-\mathbb{E}[r_{i}]\bigg{)}+\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\bigg{(}\sum_{q=0}^{m-1}\hat{D}_{q}(s_{k}^{(x)},a_{k}^{(x)})-\mathbb{E}[D_{m}(s_{k},a_{k})]\bigg{)}$
+$+\eta_{m}^{k_{m}}\bigg{(}D_{m}(s_{k}^{(x)},a_{k}^{(x)})-\mathbb{E}[D_{m}(s_{k},a_{k})]\bigg{)}$ (46)
+Assume that $k_{0}\leq k_{1}\leq...k_{m}=k$, the D estimates share at most $k_{m}=k$ reward terms $\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}$, Using Hoeffding inequality and union bound, we obtain that with probability $1-\delta$, each k empirical average reward $\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}$ deviates from the true expected reward $\mathbb{E}[r_{i}]$ by at most $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$.
+Hence, with probability $1-\delta$, $\forall m\geq 1$, we have:
+$\Delta_{n}^{m}\leq\epsilon\sum_{i=1}^{k_{m}-1}\bigg{(}\eta_{m}^{i}-\eta_{m-1}^{i}\bigg{)}+\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\sum_{q=0}^{m-1}\Delta_{n-1}^{q}+\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$
+$=\epsilon\bigg{(}\frac{1-\eta_{m}^{k_{m}}}{1-\eta_{m}}-\frac{1-\eta_{m-1}^{k_{m}}}{1-\eta_{m-1}}\bigg{)}+\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\sum_{q=0}^{m-1}\Delta_{n-1}^{q}+\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$ (47)
+and $\Delta_{n}^{0}\leq\epsilon\frac{1-\eta_{0}^{k_{0}}}{1-\eta_{0}}+\eta_{0}^{k_{0}}\Delta_{n-1}^{0}$
 Summing the two previous inequalities gives:
-$\sum_{z=0}^{Z}\Delta_{t}^{z}\leq\epsilon\frac{1-\gamma_{0}^{k_{0}}}{1-\gamma_{0}}+\epsilon\sum_{z=1}^{Z}\bigg{(}\frac{1-\gamma_{z}^{k_{z}}}{1-\gamma_{z}}-\frac{1-\gamma_{z-1}^{k_{z}}}{1-\gamma_{z-1}}\bigg{)}+\sum_{z=1}^{Z}\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\sum_{u=0}^{z-1}\Delta_{t-1}^{u}+\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}$
-$=\underbrace{\epsilon\frac{1-\gamma_{Z}^{k_{Z}}}{1-\gamma_{Z}}+\epsilon\sum_{z=0}^{Z-1}\frac{\gamma_{z}^{k_{z+1}}-\gamma_{z}^{k_{z}}}{1-\gamma_{z}}}_{\text{(*)variance term}}+\underbrace{\sum_{z=1}^{Z}\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\sum_{u=0}^{z-1}\Delta_{t-1}^{u}+\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}}_{\text{(**)bias term}}$ (45)
+$\sum_{m=0}^{M}\Delta_{n}^{m}\leq\epsilon\frac{1-\eta_{0}^{k_{0}}}{1-\eta_{0}}+\epsilon\sum_{m=1}^{M}\bigg{(}\frac{1-\eta_{m}^{k_{m}}}{1-\eta_{m}}-\frac{1-\eta_{m-1}^{k_{m}}}{1-\eta_{m-1}}\bigg{)}+\sum_{m=1}^{M}\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\sum_{q=0}^{m-1}\Delta_{n-1}^{q}+\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$
+$=\underbrace{\epsilon\frac{1-\eta_{M}^{k_{M}}}{1-\eta_{M}}+\epsilon\sum_{m=0}^{M-1}\frac{\eta_{m}^{k_{m+1}}-\eta_{m}^{k_{m}}}{1-\eta_{m}}}_{\text{(*)variance term}}+\underbrace{\sum_{m=1}^{M}\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\sum_{q=0}^{m-1}\Delta_{n-1}^{q}+\eta_{m}^{k_{m}}\Delta_{n-1}^{m}}_{\text{(**)bias term}}$ (48)
 Let’s focus now further on the bias term (**)
-$\sum_{z=1}^{Z}\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\sum_{u=0}^{z-1}\Delta_{t-1}^{u}+\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}=\sum_{u=0}^{Z-1}\sum_{z=u+1}^{Z}\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\Delta_{t-1}^{u}+\sum_{z=1}^{Z}\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}$
-$=\sum_{u=0}^{Z-1}\Delta_{t-1}^{u}\bigg{(}\sum_{z=u+1}^{Z}\gamma_{z}^{k_{z}}-\sum_{z=u}^{Z-1}\gamma_{z}^{k_{z+1}}\bigg{)}+\sum_{z=1}^{Z}\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}$
-$=\sum_{u=0}^{Z-1}\Delta_{t-1}^{u}\bigg{(}\sum_{z=u+1}^{Z-1}(\gamma_{z}^{k_{z}}-\gamma_{z}^{k_{z+1}})+(\gamma_{Z}^{k_{z}}-\gamma_{u}^{k_{u+1}})\bigg{)}+\sum_{z=1}^{Z}\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}$
-$=\sum_{u=0}^{Z-1}\sum_{z=u+1}^{Z-1}(\gamma_{z}^{k_{z}}-\gamma_{z}^{k_{z+1}})\Delta_{t-1}^{u}+\gamma_{Z}^{k_{z}}\sum_{z=0}^{Z}\Delta_{t-1}^{z}+\sum_{z=0}^{Z-1}(\gamma_{z}^{k_{z}}-\gamma_{z}^{k_{z+1}})\Delta_{t-1}^{z}$
-$=\sum_{u=0}^{Z-1}\sum_{z=u}^{Z-1}(\gamma_{z}^{k_{z}}-\gamma_{z}^{k_{z+1}})\Delta_{t-1}^{u}+\gamma_{Z}^{k_{z}}\sum_{z=0}^{Z}\Delta_{t-1}^{z}$
-$=\sum_{z=0}^{Z-1}(\gamma_{z}^{k_{z}}-\gamma_{z}^{k_{z+1}})\sum_{u=0}^{z}\Delta_{t-1}^{u}+\gamma_{Z}^{k_{z}}\sum_{z=0}^{Z}\Delta_{t-1}^{z}$ (46)
+$\sum_{m=1}^{M}\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\sum_{q=0}^{m-1}\Delta_{n-1}^{q}+\eta_{m}^{k_{m}}\Delta_{n-1}^{m}=\sum_{q=0}^{M-1}\sum_{m=q+1}^{M}\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\Delta_{n-1}^{q}+\sum_{m=1}^{M}\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$
+$=\sum_{q=0}^{M-1}\Delta_{n-1}^{q}\bigg{(}\sum_{m=q+1}^{M}\eta_{m}^{k_{m}}-\sum_{m=q}^{M-1}\eta_{m}^{k_{m+1}}\bigg{)}+\sum_{m=1}^{M}\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$
+$=\sum_{q=0}^{M-1}\Delta_{n-1}^{q}\bigg{(}\sum_{m=q+1}^{M-1}(\eta_{m}^{k_{m}}-\eta_{m}^{k_{m+1}})+(\eta_{M}^{k_{m}}-\eta_{q}^{k_{q+1}})\bigg{)}+\sum_{m=1}^{M}\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$
+$=\sum_{q=0}^{M-1}\sum_{m=q+1}^{M-1}(\eta_{m}^{k_{m}}-\eta_{m}^{k_{m+1}})\Delta_{n-1}^{q}+\eta_{M}^{k_{m}}\sum_{m=0}^{M}\Delta_{n-1}^{m}+\sum_{m=0}^{M-1}(\eta_{m}^{k_{m}}-\eta_{m}^{k_{m+1}})\Delta_{n-1}^{m}$
+$=\sum_{q=0}^{M-1}\sum_{m=q}^{M-1}(\eta_{m}^{k_{m}}-\eta_{m}^{k_{m+1}})\Delta_{n-1}^{q}+\eta_{M}^{k_{m}}\sum_{m=0}^{M}\Delta_{n-1}^{m}$
+$=\sum_{m=0}^{M-1}(\eta_{m}^{k_{m}}-\eta_{m}^{k_{m+1}})\sum_{q=0}^{m}\Delta_{n-1}^{q}+\eta_{M}^{k_{m}}\sum_{m=0}^{M}\Delta_{n-1}^{m}$ (49)
 Finally, we obtain:
-$\sum_{z=0}^{Z}\Delta_{t}^{z}\leq\underbrace{\epsilon\bigg{(}\frac{1-\gamma^{k}}{1-\gamma}\bigg{)}+\epsilon\bigg{(}\sum_{z=0}^{Z-1}\frac{\gamma_{z}^{k_{z+1}}-\gamma^{k_{z}}_{z}}{1-\gamma_{z}}\bigg{)}}_{\text{variance reduction}}+\underbrace{\sum_{z=0}^{Z-1}\bigg{(}\gamma^{k_{z}}_{z}-\gamma^{k_{z+1}}_{z}\bigg{)}\sum_{u=0}^{z}\Delta_{t-1}^{u}+\gamma^{k}\sum_{z=0}^{Z}\Delta_{t-1}^{z}}_{\text{bias introduction}}$ (47)
+$\sum_{m=0}^{M}\Delta_{n}^{m}\leq\underbrace{\epsilon\bigg{(}\frac{1-\eta^{k}}{1-\eta}\bigg{)}+\epsilon\bigg{(}\sum_{m=0}^{M-1}\frac{\eta_{m}^{k_{m+1}}-\eta^{k_{m}}_{m}}{1-\eta_{m}}\bigg{)}}_{\text{variance reduction}}+\underbrace{\sum_{m=0}^{M-1}\bigg{(}\eta^{k_{m}}_{m}-\eta^{k_{m+1}}_{m}\bigg{)}\sum_{q=0}^{m}\Delta_{n-1}^{q}+\eta^{k}\sum_{m=0}^{M}\Delta_{n-1}^{m}}_{\text{bias introduction}}$ (50)

### v4

@@
 Abstract
-In numerous episodic reinforcement learning (RL) settings, SARSA-based methodologies are employed to enhance policies aimed at maximizing returns over long horizons.
-Conventional SARSA algorithms, however, have difficulties in balancing bias and variation due to the reliance on a singular, fixed discount factor.
-This study expands the temporal difference decomposition approach, TD($\Delta$), to the SARSA algorithm, which we designate as SARSA($\Delta$).
-SARSA, a widely utilised on-policy RL method, enhances action-value functions via temporal difference updates.
-TD($\Delta$) facilitates learning over several time-scales by breaking the action-value function into components associated with distinct discount factors.
-This decomposition improves learning efficiency and stability, particularly in problems necessitating long-horizon optimization.
-We illustrate that our methodology mitigates bias in SARSA’s updates while facilitating accelerated convergence in both deterministic and stochastic environments.
-Experimental findings across many benchmark tasks indicate that the proposed SARSA($\Delta$) surpasses conventional TD learning methods in both tabular and deep RL environments.
+In numerous episodic reinforcement learning (RL) environments, SARSA-based methodologies are employed to enhance policies aimed at maximizing returns over long horizons.
+Traditional SARSA algorithms face challenges in achieving an optimal balance between bias and variation, primarily due to their dependence on a single, constant discount factor ($\eta$).
+This investigation enhances the temporal difference decomposition method, TD($\Delta$), by applying it to the SARSA algorithm, now designated as SARSA($\Delta$).
+SARSA is a widely used on-policy RL method that enhances action-value functions via temporal difference updates.
+By splitting the action-value function down into components that are linked to specific discount factors, SARSA($\Delta$) makes learning easier across a range of time scales.
+This analysis makes learning more effective and ensures consistency, particularly in situations where long-horizon improvement is needed.
+The results of this research show that the suggested strategy works to lower bias in SARSA’s updates and speed up convergence in both deterministic and stochastic settings, even in dense reward Atari environments.
+Experimental results from a variety of benchmark settings show that the proposed SARSA($\Delta$) outperforms existing TD learning techniques in both tabular and deep RL environments.
 keywords:
-SARSA, Temporal Difference, Action-Value Functions, On-policy, Finite Horizon
+SARSA, Temporal Difference (TD($\Delta$)), Action-Value Functions, On-policy, Long-Horizon, Time-Scale Decomposition
 1 Introduction
-In reinforcement learning (RL), the objective is to enhance an agent’s actions to maximize cumulative rewards through interaction with an environment.
-Temporal difference (TD) learning methods, like Q-learning and SARSA, have demonstrated efficacy across several tasks by allowing agents to estimate action-value functions that forecast anticipated future rewards.
-Traditionally, these strategies utilise a discount factor $10\leq\eta<$ 1, with values approaching $\eta=0$ to emphasise short-term rewards over long-term ones, therefore diminishing the planning horizon and enhancing learning stability and efficiency[1].
-Discount factors $\eta<1$ frequently yield enhanced results during the early phases of learning, as demonstrated by Prokhorov and Wunsch[2].
-In specific contexts, such as long-horizon tasks[3, 4] where long-term planning is essential—like navigation or decision-making tasks with deferred rewards—this method can create bias, constraining the agent’s capacity to acquire optimal long-horizon policies.
-For instance: SARSA[1] (State-Action-Reward-State-Action) is an on-policy RL algorithm that modifies an action-value function Q(s,a) according to the state-action transitions encountered by the agent.
-The conventional SARSA algorithm performs well in contexts with shorter time horizons; however, in long-horizon tasks, the singular discount factor $\eta$ employed in SARSA results in a compromise between bias and variance.
-Small values of $\eta$ yield myopic policies, whereas larger values induce substantial variance in the estimated action-value functions, rendering the learning process unstable and inefficient.
-In response, other solutions have been suggested, including[4, 2, 5, 6, 7, 8].
-Recent research in TD learning has investigated the segmentation of value functions across multiple time scales, enabling agents to enhance value estimates concurrently at different horizons[8].
-The TD($\Delta$) architecture improves scalability and performance in scenarios involving long-term dependencies by breaking down value functions into components that incorporate different discount factors, known as TD($\Delta$).
-The adaptability of the TD($\Delta$) framework shows considerable promise in addressing the bias-variance trade-offs that are characteristic of TD learning.
-We propose SARSA($\Delta$), an enhancement of the TD($\Delta$) approach integrated into the SARSA algorithm, based on this principle.
+Reinforcement learning (RL) agents encounter significant challenges in the optimization of long-horizon rewards.
+Traditional methods employing singular discount variables encounter inherent challenges: small $\eta$ results in low-variance but myopic policies, whereas large $\eta$ suffers from excessive variance.
+Fixed discounts remain constant despite variations in reward densities[1, 2, 3].
+Temporal difference (TD) learning methods, such as Q-learning and SARSA, have shown effectiveness in various tasks by enabling agents to estimate action-value functions that predict expected future rewards.
+Traditionally, these strategies use a discount factor of $0\leq\eta<1$, with values getting closer to $\eta=0$ to put more weight on short-term rewards than long-term ones.
+This approach shortens the planning horizon and makes learning more stable and efficient[2].
+Prokhorov and Wunsch[4] show that discount factors $\eta<1$ often lead to better results in the early stages of learning.
+In certain scenarios, such as long-horizon tasks[5, 6] where long-term planning is crucial—like navigation or decision-making tasks that involve delayed rewards—this approach can introduce bias, complicating the agent’s ability to determine the most effective long-horizon policies.
+For example, SARSA[2] (State-Action-Reward-State-Action) is an on-policy RL algorithm that updates the action-value function Q(s,a) according to the state-action transitions the agent sees.
+The standard SARSA algorithm works well in environments with shorter time horizons; however, it faces challenges when utilized for tasks with longer horizons.
+The main reason for the observed phenomenon is the use of a single discount factor $\eta$, which requires a trade-off between bias and variance.
+This often leads to instability and inefficiency during the learning process.
+A variety of recommendations exist for addressing this issue [6, 4, 7, 8, 9, 1, 10].
+Through the use of weighted combinations of value functions, Fedus et al.
+[10] presented the concept of hyperbolic discounting.
+However, their formulation is restricted to state-value functions and does not extend to action-value functions.
+Meta-gradient RL[8] is a technique that enables agents to adapt their learning processes dynamically.
+Meta-gradient means finding or improving targets and hyperparameters during training, which could make learning more efficient and adaptable.
+A gradient-based meta-learning algorithm capable of online adaptation to the true value function while engaging with and learning from the environment.
+However, akin to other methodologies, meta-gradient RL primarily emphasizes state-value functions, rather than action-value formulations.
+The work that is most relevant in this case is by Romoff et al.[1], which provides a theoretical basis for temporal difference decomposition using the TD($\Delta$) framework.
+TD($\Delta$) segments the state-value function V(s) into components using different discount factors, also known as TD($\Delta$).
+This lets the agent learn about short-term and long-term returns separately, which can then be combined to improve performance.
+This decomposition framework demonstrates considerable potential in addressing the bias-variance trade-off that is intrinsic to temporal difference learning.
+However, TD($\Delta$) is constrained to state-value functions, which presents a limitation in the application of multi-timescale representations to action-value-based algorithms such as SARSA and Q-learning.
+This article proposes SARSA($\Delta$), an enhancement of the TD($\Delta$) approach integrated into the SARSA algorithm, based on this principle.
 TD($\Delta$) emphasizes the actor-critic framework and conventional TD learning, whereas SARSA($\Delta$) divides the state-action value function into multiple partial estimators corresponding to various discount factors, known as delta estimators.
-These estimators approximate the difference $D_{m}(s,a)=Q_{\eta_{m}}(s,a)-Q_{\eta_{m-1}}(s,a)$ between action-value functions, thereby improving learning efficiency in scenarios where actions affect long-horizon outcomes.
-Similar to TD($\Delta$), SARSA($\Delta$) seeks to optimize a set of delta estimators, where each estimator is associated with a distinct discount factor.
-With smaller discount factors, convergence happens faster because to segmentation, while greater discount factors strengthen this foundation for better long-horizon planning.
-This study introduces a multi-step SARSA($\Delta$) update rule that disaggregates action-value functions, facilitating accelerated convergence and enhanced policy efficacy, especially in intricate deterministic and stochastic environments.
-Furthermore, we illustrate that SARSA($\Delta$) may seamlessly adapt to any multi-step reinforcement learning techniques, including n-step SARSA and eligibility traces.
-Through theoretical analysis and empirical findings, we emphasise the benefits of SARSA($\Delta$) in diverse tasks, including a basic ring MDP utilised by Kearns & Singh[9], demonstrating its superiority over conventional SARSA in scenarios necessitating a balance between short-term action consequences and long-term reward maximisation.
-SARSA($\Delta$) possesses the capability to adapt discount factors across varying time scales, serving as an effective instrument in environments where time scale is critical.
-The impetus for extending TD($\Delta$) to SARSA($\Delta$) lies in the fact that temporal difference (TD($\Delta$)) learning offers a framework for deconstructing value functions across many time scales, hence overcoming the constraints of single-scale learning in long-horizon problems.
-By extending TD($\Delta$) to SARSA($\Delta$), we seek to decompose the action-value function Q(s,a) into elements that emphasise distinct discount factors.
-This decomposition enables SARSA to utilise the advantages of learning across several time scales, hence stabilising the learning process and mitigating the bias-variance tradeoff.
-This addon provides a more detailed and scalable method for action-value function learning in intricate environments.
-The primary contributions of this paper are as follows: (i) We expand TD($\Delta$) to SARSA($\Delta$) by partitioning the action-value function Q(s,a) into delta components over several time-scales.
-(ii) The suggested SARSA($\Delta$) algorithm enhances learning efficiency and stability by independently learning components for various discount factors and integrating them to represent the complete action-value function.
-(iii) Both theoretical and empirical evaluations have shown that this multi-scale breakdown has a number of benefits, particularly in environments where there are long-horizon rewards.
-The ensuing portions of the paper are organized as follows: Section 2 presents a discourse on pertinent literature.
+These estimators give a rough estimate of the difference $D_{m}(s,a)=Q_{\eta_{m}}(s,a)-Q_{\eta_{m-1}}(s,a)$ between action-value functions.
+This makes learning faster in contexts where actions yield long-term effects.
+Like TD($\Delta$), SARSA($\Delta$) aims to enhance a collection of delta estimators, with each estimator linked to a specific discount factor.
+Smaller discount factors lead to quicker convergence via segmentation, whereas larger discount factors enhance this foundation, facilitating improved long-term planning.
+This paper provides a multi-stage SARSA($\Delta$) updating algorithm that separates action-value functions, which allows for faster convergence and improved policy efficacy, especially in complex deterministic and stochastic environments like Atari games.
+Moreover, our research shows that SARSA ($\Delta$) can seamlessly accommodate various multi-step RL methodologies, such as n-step SARSA and eligibility traces.
+This study emphasizes the benefits of SARSA($\Delta$) in various tasks, including a basic ring MDP used by Kearns & Singh[11].
+It outperforms conventional SARSA in scenarios requiring a balance between short-term action consequences and long-term reward maximization.
+SARSA ($\Delta$) may adjust discount factors for multiple time scales, making it valuable where time scale is precious.
+The reason for extending TD($\Delta$) to SARSA($\Delta$) is that temporal difference (TD($\Delta$)) learning provides a way to break down value functions across a range of time scales, which solves the problems that come up when learning on a single scale for long-horizon scenarios.
+We want to break down the action-value function Q(s,a) into components that focus on different discount factors by extending TD($\Delta$) to SARSA($\Delta$).
+This decomposition enables SARSA to leverage the advantages of learning over multiple time scales, which stabilizes the learning process and reduces the bias-variance trade-off.
+This add-on gives you a more detailed and scalable way to learn the action-value function in complex environments.
+Table 1 presents a comparison of the key features of TD($\Delta$)/SARSA($\Delta$) and Standard TD/SARSA-Learning.
+Table 1: Comparison of the essential properties of TD($\Delta$) and SARSA($\Delta$).
+Feature TD($\Delta$)[1] SARSA($\Delta$)-Learning (proposed by us) Standard TD[12] SARSA-Learning[12]
+Value function decomposition Yes (by $\eta$) No No No
+Action-value function decomposition No Yes (by $\eta$) No No
+Handles multiple time-scales Yes Yes No No
+Learning stability Enhanced, particularly long-term Enhanced, particularly long-term Reduced for long-term rewards Reduced for long-term rewards
+Scalability Improved Improved Restricted by variance/bias Restricted by variance/bias
+Recent research on value decomposition[1] has demonstrated potential for state-value functions; however, it identifies three significant shortcomings that this work aims to address: (i) Action-Value Formulation: No current method effectively decomposes Q(s, a) while maintaining on-policy guarantees.
+This study extends TD($\Delta$) to SARSA($\Delta$) by decomposing the action-value function Q(s,a) into delta components across multiple time scales.
+(ii) Theoretical Foundations: Current analyses do not address the bias-variance trade-off concerning action values.
+(iii) The proposed SARSA($\Delta$) algorithm improves learning efficiency and stability through the independent learning of components corresponding to different discount factors, which are subsequently integrated to form the comprehensive action-value function.
+(iv) Both theoretical and empirical evaluations have shown that this multi-scale breakdown has a number of benefits, particularly in environments where there are long-horizon rewards, as demonstrated in Table 1.
+The subsequent portions of the paper are organized as follows: Section 2 presents a discussion on relevant literature.
 Section 3 provides the essential background information.
@@
 2 Related Work
-In reinforcement learning (RL), optimising for long-horizon rewards is tough due to the complexity of learning with an undiscounted return.
-Temporal discounting is frequently employed to streamline this process; nevertheless, it may induce bias.
-Researchers[8] have investigated techniques to separate value functions across various time scales by analysing disparities among value functions with lower discount factors, hence enhancing scalability and performance.
-The study introduces $\gamma$-nets[10], a technique for generalising value function estimate across timescales in RL.
-The authors of the paper[11] present an approach for integrating hyperbolic discounting into RL.
-Traditional RL employs exponential discounting, which does not align with the hyperbolic discounting observed in human and animal behaviors.
-The authors present an agent designed to approximate hyperbolic discount functions by utilizing temporal-difference learning methods.
-Furthermore, the study reveals that learning value functions across multiple time horizons improves performance, especially when implemented as an auxiliary task in value-based RL algorithms such as Rainbow.
-This method demonstrates potential in contexts characterized by uncertainty and risk.
-Expected SARSA[7], a variant of SARSA, utilizes information regarding the stochastic nature of the behavior policy to execute updates with reduced variance, which speeds up learning.
-This study[12] presents temporal-difference (TD) learning methods, which underpin numerous RL algorithms, notably SARSA, by forecasting future outcomes based on the discrepancies between consecutive forecasts rather than the ultimate result.
-The research demonstrates the convergence of TD techniques under specific situations and emphasises their benefits in predictive tasks, particularly when incremental learning is crucial.
-This study[13] presents theoretical convergence results for temporal-difference learning methods utilising function approximation.
-Extending this analysis to SARSA($\Delta$) enables us to assert comparable convergence properties, provided that the delta components are updated in accordance with the SARSA update rule.
-It is noteworthy that none of the previously mentioned works employ short-term estimates for training long-term action-value functions.
-Our method can also be utilised as a generalised action-value function, and its capacity to query smaller time-scales represents an additional advantage.
-Performance improvements primarily arise from the separation of action-value functions across varying time scales through the use of TD($\Delta$) in SARSA.
-Recently, there has been some research conducted on the precise selection of the discount factor[2, 5, 6, 4].
-The authors[6] of this study introduce meta-gradient methods for the adaptive selection of discount factors in RL, emphasising the necessity of balancing short- and long-term rewards, which corresponds with the rationale for dissecting action-value functions across several discount factors in SARSA($\Delta$).
-Lastly, another extensive category of study pertinent to our research indirectly encompasses hierarchical RL and the bias-variance trade-off.
-These studies[14, 15, 16, 17, 18, 19, 20] present the concept of deconstructing value functions into subtasks by hierarchical RL, analogous to the manner in which TD($\Delta$) disaggregates value functions across temporal scales.
-This substantiates the assertion that value function decomposition can enhance learning efficacy and stability.
-This paper[9] functions as a theoretical foundation for comprehending the bias-variance trade-off that arises when decomposing Q-values across various time-scales, and it also establishes constraints on the bias and variance of TD updates.
-This study[21] presents empirical findings demonstrating the efficacy of deep RL techniques utilizing value function approximation in Atari-like environments.
-It also endorses the necessity for techniques that diminish variance in long-horizon tasks.
+Optimizing for long-horizon rewards in reinforcement learning (RL) is difficult owing to the complexities of learning with undiscounted returns.
+Temporal discounting is often used to simplify this process; however, it can introduce bias.
+2.1 Temporal Decomposition Approaches
+In their investigation, Romoff et al.
+[1] explored the differentiation of value functions across multiple time scales by scrutinizing individuals with diminished discount factors.
+The system became more efficient and simpler to expand.
+In RL, Sherstan et al.[13] introduced $\gamma$-nets, a method designed to enhance value function estimates across various timescales.
+Authors’ recent work [10, 14] showed how to add hyperbolic discounting to RL.
+Exponential discounting was used in traditional RL, but this didn’t match the hyperbolic discounting seen in the behavior of human and animals.
+The authors showed an agent that used temporal-difference learning methods to get close to hyperbolic discount functions.
+The research also found that learning value functions across multiple time horizons increases performance, particularly when used as an auxiliary task in value-based RL algorithms like Rainbow.
+This approach has shown potential in risky and uncertain environments.
+Recent research [15, 4, 7, 8, 6, 16, 17] focuses on the exact selection of the discount factor.
+Xu 124 et al.
+[8] proposed meta-gradient methods for choosing discount factors in RL that change over time.
+They stressed the importance of balancing short- and long-term rewards, which is similar to why we break down action-value functions across multiple discount factors in SARSA($\Delta$).
+Lastly, hierarchical RL and the bias-variance trade-off are part of a large body of research that is related to our study in an indirect way.
+These studies [18, 19, 20, 21, 22, 23, 24] introduced the idea of hierarchical RL breaking value functions down into smaller tasks, similar to how TD($\Delta$) broke value functions down across different time scales.
+These results backed up the claim that breaking down value functions can make learning more effective and stable.
+2.2 Multi-Scale RL
+To address issues in RL, especially when the optimum value function is complex and poorly represented by conventional DRL methods, van Seijen et al.
+[24] introduced Hybrid Reward Architectures (HRA).
+The HRA was designed to enhance learning consistency and effectiveness by partitioning the reward function and allocating distinct value functions to each component.
+Efficient multi-horizon learning in off-policy RL aims to improve agents’ ability to make predictions and decisions over a wide range of time scales.
+This is important for navigating through complex and unpredictable environments.
+Ali 141 et al.
+[25, 26] investigated designs that enable agents to learn value estimates over several horizons at the same time.
+This improves their capacity to adjust and flourish in novel environments.
+In the realm of artificial intelligence and robotics, coordinating agents with limited communication and across diverse planning periods is a significant challenge to multi-horizon, multi-agent planning.
+Seiler et al.
+[27] introduced Decentralized Monte Carlo Tree Search (Dec-MCTS) and its variants, which have arisen as formidable instruments for addressing these challenges, allowing agents to plan effectively and adaptively in dynamic, unpredictable environments.
+Benechehab et al.[28] came up with a novel technique in model-based RL (MBRL) called multi-timestep models, aimed at mitigating the problem of escalating prediction errors over extended simulated trajectories.
+By optimizing for several future steps instead of only one-step-ahead forecasts, these models enhance the precision and resilience of long-term predictions, especially in complex or chaotic environments.
+Bonnet et al.
+[29] presented multi-step meta-gradient RL, which uses meta-gradients that have been built up across many stages instead of just one.
+This makes learning algorithms more flexible.
+This method may provide richer and more useful learning signals, but it does come with some significant trade-offs.
+But it also increases computational complexity and variation, which may hinder performance if not handled carefully.
+While deep RL (DRL) can teach agents difficult tasks, it still struggles to improve sampling efficiency and adapt to changing environments.
+Eligibility traces, a well-established RL approach, expedite learning, but parameter dependencies make their integration with deep neural networks difficult.
+Kobayashi[30] introduced several time-scale eligibility traces to equilibrate short-term and long-term credit assignment.
+This technique improves learning speed and policy quality in online DRL environments by substituting the most important adaptively gathered traces.
+2.3 SARSA Variants and Traditional RL Methods
+Expected SARSA [9] is a kind of SARSA that speeds up learning by exploiting information from a stochastic behavior policy to create updates that are less variance.
+Many RL algorithms, like SARSA, employ temporal-difference (TD) learning, which looks at the differences between two predictions made one after the other instead of the final result.
+Sutton[31] came up with this notion for the first time in 1988.
+The research shows that TD approaches may be useful for predicting jobs, especially where incremental learning is crucial, and that they can even operate jointly in certain cases.
+Theoretical convergence for function approximation-based TD learning algorithms was shown by Tsitsiklis and Van Roy[32].
+If you follow the SARSA update rule for delta component updates, this research may be used with SARSA($\Delta$) to get the same convergence properties.
+None of the previously mentioned research discussed short-term estimations to train long-term action-value functions.
+One benefit is that you may ask about shorter time periods and utilize our method as a generic action-value function.
+Separating action-value functions across time scales using TD($\Delta$) in SARSA makes a big difference in performance.
+Theoretical work by Kearns and Singh[11] helped us comprehend the bias-variance trade-off that comes up when we break down Q-values over multiple time scales.
+It also set limits on the bias and variance of TD updates.
+Most of the methods listed above focus on state values, although action-value decomposition has not yet been looked at.
+Our proposed strategy focuses on breaking down $Q(s,a)$ over different time scales, which makes long-term learning more stable.
 3 Background and notation
 Examine a completely observable Markov Decision Process (MDP) [33], characterized by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ signifies the action space, and $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}\rightarrow[0,1]$ represents the transition probabilities that associate state-action pairings with distributions across subsequent states, whereas $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denotes the reward function.
-At each timestep $n$, the agent occupies state $s_{n}$, chooses an action $a_{n}$, obtains a reward $r_{n}=r(s_{n},a_{n})$, and transitions to the subsequent state $s_{n+1}\sim\mathcal{P}(\cdot|s_{n},a_{n})$.
-In a conventional Markov Decision Process (MDP) framework, an agent seeks to maximize the discounted return defined as $Q_{\eta}^{\pi}(s,a)=\mathbb{E}\left[\sum_{n=0}^{\infty}\eta^{n}r_{n+1}\mid s_{n}=s,a_{n}=a\right]$, where $\eta$ denotes the discount factor and $\pi:\mathcal{S}\rightarrow\mathcal{A}\rightarrow[0,1]$ represents the policy adhered to by the agent.
+At each timestep $n$, the agent is state $s_{n}$, selects an action $a_{n}$, receives a reward $r_{n}=r(s_{n},a_{n})$, and transitions to the next state $s_{n+1}\sim\mathcal{P}(\cdot|s_{n},a_{n})$.
+Within a standard MDP framework, an agent tries to get the highest possible discounted return, which is defined as $Q_{\eta}^{\pi}(s,a)=\mathbb{E}\left[\sum_{n=0}^{\infty}\eta^{n}r_{n+1}\mid s_{n}=s,a_{n}=a\right]$, where $\eta$ is the discount factor and $\pi:\mathcal{S}\rightarrow\mathcal{A}\rightarrow[0,1]$ stands for the policy that the agent follows.
 The action-value function $Q_{\eta}^{\pi}(s,a)$ is determined as the fixed point of the Bellman operator $\mathcal{T}Q=r^{\pi}+\eta\mathcal{P}^{\pi}Q$, where $r^{\pi}$ denotes the expected immediate reward and $\mathcal{P}^{\pi}$ represents the transition probability operator associated with the policy $\pi$.
@@
 An on-policy RL technique called SARSA learns the action-value function Q(s, a) for a given policy $\pi$.
-The action-value function represents the expected total reward that comes from state s, taking action a, and following policy $\pi$.
-Then the SARSA update rule is clarified as follows:
+The action-value function represents the expected total reward derived from state s, executing action a, and adhering to policy $\pi$.
+Then the SARSA update rule is expressed as follows:
 $Q(s_{n},a_{n})\leftarrow Q(s_{n},a_{n})+\alpha\left[r_{n}+\eta Q(s_{n+1},a_{n+1})-Q(s_{n},a_{n})\right]$ (1)
@@
 On the other hand, for a complete trajectory, we can employ the discounted sum of one-step TD errors, generally known as the $\lambda$-return [34] or, equivalently, the Generalized Advantage Estimator (GAE) [35].
-The GAE enhances advantage estimates by reconciling the trade-off between variance and bias via parameters $\lambda$ and $\eta$.
-The equation for the Generalized Advantage Estimator is generally expressed as:
+The GAE improves advantage estimates by balancing the trade-off between variance and bias through the parameters $\lambda$ and $\eta$.
+The Generalized Advantage Estimator is typically represented by the following equation:
 $A(s_{n},a_{n})=\sum_{k=0}^{\infty}(\lambda\eta)^{k}\delta_{t+k}^{\eta}$ (2)
-Where $\delta_{n+k}$ is the TD error at time n, computed as follows: $\delta_{n+k}=r_{n+k}+\eta Q(s_{n+k+1},a_{n+k+1})-Q(s_{n+k},a_{n+k})$
+Where $\delta_{n+k}$ is the TD error at time n, computed as follows: $\delta_{n+k}=r_{n+k}+\eta Q(s_{n+k+1},a_{n+k+1})-Q(s_{n+k},a_{n+k})$.
 Loss function for Q-Value estimation utilizing GAE.
@@
 $\mathcal{L}(\theta)=\mathbb{E}\bigg{[}\bigg{(}Q(s_{n},a_{n};\theta)-\bigg{(}Q(s_{n},a_{n})+A(s_{n},a_{n})\bigg{)}\bigg{)}^{2}\bigg{]}$ (3)
-Despite SARSA being on-policy and lacking an explicit policy update, policy selection in SARSA can still be influenced by Eq.4.
-For actor-critic architectures[25, 26, 27], the action-value function is updated according to Eq.3, and a stochastic parameterized policy (actor, $\pi_{\nu}(a|s)$) is learned from this value estimator through the advantage function, with the loss being.
+Despite SARSA being on-policy and lacking an explicit policy update, policy selection in SARSA can still be influenced by Eq.
+4.
+For actor-critic architectures [36, 37, 38], the action-value function is updated according to Eq.
+3, and a stochastic parameterized policy (actor, $\pi_{\nu}(a|s)$) is learned from this value estimator through the advantage function, with the loss being.
 $\mathcal{L}(\nu)=\mathbb{E}\bigg{[}-log\pi(a|s;\nu)A(s,a)\bigg{]}$ (4)
-Proximal Policy Optimization (PPO) [28], an enhancement of actor-critic methodologies, confines policy updates to a designated optimization area, referred to as a trust region, utilizing a clipping objective that evaluates the current parameters, $\nu$, against the prior parameters, $\nu_{old}$:
+Proximal Policy Optimization (PPO) [39] is an improvement on actor-critic methods.
+It limits policy updates to a specific area of optimization called a trust region.
+It does this by using a clipping objective that compares the current parameters, $\nu$, to the old parameters, $\nu_{old}$:
 $\mathcal{L}(\nu)=\mathbb{E}\bigg{[}min\bigg{(}\rho(\nu)A(s,a),\psi(\nu)A(s,a)\bigg{)}\bigg{]}$ (5)
-In on-policy methods like SARSA, the likelihood ratio $\rho(\nu)$ between the current policy and a prior policy is typically employed.
+In on-policy approaches such as SARSA, the probability ratio $\rho(\nu)$ between the current policy and a preceding policy is often used.
 SARSA evaluates actions in accordance with the current policy, concurrently updating the action-value function as actions are performed.
@@
 4 Methodology
-The TD($\Delta$) paradigm boosts regular TD learning by splitting action-value functions over several discount factors $\eta_{0},\eta_{1},...,\eta_{m}$.
-This decomposition facilitates the derivation of the action-value function as a sum of delta estimators, where each estimator indicates the distinction between action-value functions at various discount factors.
-Concentrating on smaller, more manageable time scales facilitates the regulation of learning variance and bias, which constitutes the primary benefit of this approach.
+The TD($\Delta$) paradigm boosts regular TD learning by splitting action-value functions over several discount factors, $\eta_{0},\eta_{1},…\eta_{m}$.
+This breakdown makes it easier to find the action-value function as a sum of delta estimators, where each estimator shows how action-value functions differ at different discount factors.
+The main advantage of this method is that it makes it easier to control learning variance and bias by focusing on smaller, more controllable time scales.
 4.1 TD($\Delta$) Framework
@@
 This section initiates with an overview of the delta estimators ($D_{m}$) pertaining to the action-value functions $Q$ employed in SARSA.
-SARSA is extended to TD($\Delta$), which is referred to as SARSA($\Delta$).
-SARSA($\Delta$) seeks to derive the action-value function $Q(s,a)$, which indicates the anticipated return of following policy $\pi$ from state $s$, executing action $a$, and then persistently adhering to $\pi$ thereafter.
-The objective is to enhance the total reward by optimizing the action-value function over various time-scales.
-The fundamental concept in SARSA($\Delta$) is the segmentation of the action-value function $Q(s,a)$ into multiple delta components $D_{m}(s,a)$, where each component is linked to a distinct discount factor $\eta_{m}$.
+SARSA is expanded to TD($\Delta$), which is called SARSA($\Delta$).
+SARSA($\Delta$) aims to determine the action-value function $Q(s,a)$, representing the expected return of adhering to policy $\pi$ from state $s$, executing action $a$, and subsequently continuing to follow $\pi$.
+The goal is to optimize the total reward by improving the action-value function over various time scales.
+SARSA($\Delta$) fundamentally decomposes the action-value function $Q(s,a)$ into multiple delta components $D_{m}(s,a)$, each associated with a distinct discount factor $\eta_{m}$.
 The relationship between these delta components (i.e., delta function) is expressed in the following manner:
@@
 $Q_{\eta_{m}}(s,a)=\sum_{x=0}^{m}D_{x}(s,a)$ (10)
-The update rule in standard SARSA is determined by the state-action pair in accordance with the policy.
-The action-value function is updated through the TD error, with the action in the subsequent state selected in accordance with the policy:
+The policy dictates how the state-action pair in conventional SARSA determines the update rule.
+The TD error adjusts the action-value function, and the policy determines the action in the following state:
 $Q(s_{n},a_{n})\leftarrow Q(s_{n},a_{n})+\alpha\left[r_{n}+\eta Q(s_{n+1},a_{n+1})-Q(s_{n},a_{n})\right]$ (11)
 In this context, $r_{n}$ signifies the reward acquired from performing action $a_{n}$ in state $s_{n}$, whereas $(s_{n+1},a_{n+1})$ denotes the ensuing state-action pair.
-The SARSA update rule is implemented independently to each delta component $D_{n}(s,a)$.
-The modifications for each component utilize the same architecture as the conventional SARSA approach, but it is tailored to include the delta function at each distinct time-scale.
-The update in Eq.11 for single-step TD SARSA ($\Delta$) can be articulated with numerous time-scales by employing the principle of splitting the update into several discount factors[8] as demonstrated below:
+All of the delta components $D_{n}(s,a)$ are updated separately using the SARSA rule.
+The modifications to each component use the same structure as the standard SARSA method, but they are changed to incorporate the delta function at each time scale.
+The update in Eq.
+11 for single-step TD SARSA ($\Delta$) can be expressed with numerous time-scales by utilizing the principle of splitting the update into several discount factors [1] as demonstrated below:
 $D_{m}(s_{n},a_{n})=\mathbb{E}\left[(\eta_{m}-\eta_{m-1})Q_{\eta_{m-1}}(s_{n+1},a_{n+1})+\eta_{m}D_{m}(s_{n+1},a_{n+1})\right]$ (12)
-where $Q_{\eta_{m}}(s_{n},a_{n})$ represents the action-value function with a discount factor $\eta_{m}$.
-$D_{m}(s_{n},a_{n})$ denotes the delta function for SARSA ($\Delta$), while $\eta_{m}$ represents the discount factor for time-scale m, while $Q_{\eta_{m}}(s_{n+1},a_{n+1})$ is the SARSA update for action-value associated with the subsequent state-action pair.
+In this context, $Q_{\eta_{m}}(s_{n},a_{n})$ denotes the action-value function with a discount factor ($\eta_{m}$).
+$D_{m}(s_{n},a_{n})$ denotes the delta function for SARSA ($\Delta$), while $\eta_{m}$ represents the discount factor for time-scale m, while $Q_{\eta_{m}}(s_{n+1},a_{n+1})$ is the SARSA update for the action-value associated with the subsequent state-action pair.
 The single-step TD SARSA ($\Delta$) mirrors the original TD update, but it is applied to action-value functions Q rather than state-value functions V.
@@
 This denotes the expected value of the reward acquired at time n, in addition to the discounted value of the action-value function in the subsequent state-action pair $(s_{n+1},a_{n+1})$.
-Currently, rather than utilizing a singular discount factor $\eta$, we are employing a sequence of discount factors $\eta_{0},\eta_{1},.....,\eta_{m}$.
+Currently, instead of utilizing a single discount factor $\eta$, we are using a series of discount factors $\eta_{0},\eta_{1},...,\eta_{m}$.
 The delta function $D_{m}(s_{n},a_{n})$ is defined as the difference between action-value functions associated with successive discount factors from Eq.
@@
 $D_{m}(s_{n},a_{n})=Q_{\eta_{m}}(s_{n},a_{n})-Q_{\eta_{m-1}}(s_{n},a_{n})$ (14)
-To expand the delta function $D_{z}$ utilizing its definition, we replace the Bellman Eq.
+To expand the delta function $D_{m}$ utilizing its definition, we replace the Bellman Eq.
 for both $Q_{\eta_{m}}$ and $Q_{\eta_{m-1}}$, thereby expressing $D_{m}$ in relation to these action-value functions.
@@
 $D_{m}(s_{n},a_{n})=\mathbb{E}\left[\eta_{m}Q_{\eta_{m}}(s_{n+1},a_{n+1})-\eta_{m-1}Q_{\eta_{m-1}}(s_{n+1},a_{n+1})\right]$
-Decomposing the terms allows for a more detailed analysis of the expression. Utilizing the recursive relationship from Eq. 14
+Breaking down the terms facilitates a more thorough examination of the expression. Utilizing the recursive relationship from Eq. 14
 $Q_{\eta_{m}}(s_{n+1},a_{n+1})=D_{m}(s_{n+1},a_{n+1})+Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$
@@
 $D_{m}(s_{n},a_{n})=\mathbb{E}\left[(\eta_{m}-\eta_{m-1})Q_{\eta_{m-1}}(s_{n+1},a_{n+1})+\eta_{m}D_{m}(s_{n+1},a_{n+1})\right]$ (15)
-This Eq.
-demonstrates that the update for $D_{m}$ integrates the disparity in discount factors $\eta_{m}-\eta_{m-1}$ with respect to the action-value function $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$ and the bootstrapping from the subsequent step $D_{m}(s_{n+1},a_{n+1})$.
-This denotes a Bellman Eq.
-for $D_{m}$, combining a decay factor $\eta_{m}$ and the reward $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$.
+Eq.
+4.2 shows that the difference in discount factors $\eta_{m}-\eta_{m-1}$ is integrated into the update for $D_{m}$ with regard to the action-value function $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$ and the bootstrapping from the next step $D_{m}(s_{n+1},a_{n+1})$.
+The Eq.
+4.2 denotes a Bellman Eq.
+13 for $D_{m}$, combining a decay factor $\eta_{m}$ and the reward $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$.
 Consequently, it can be utilized to define the expected TD update for $D_{m}$.
 In this expression, $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$ can be expressed as the summation of $D_{m}(s_{n+1},a_{n+1})$ for $m\leq M-1$, indicating that the Bellman Eq.
-for $D_{m}$ is contingent upon the values of all delta functions $D_{m}$ for $m\leq M-1$.
+13 for $D_{m}$ is contingent upon the values of all delta functions $D_{m}$ for $m\leq M-1$.
 This approach treats the delta value function at each time-scale as an autonomous RL issue, with rewards obtained from the action-value function of the immediately preceding time-scale.
-Thus, for a target discounted action-value function $Q_{\eta_{m}}(s,a)$, all delta components can be trained simultaneously through a TD update, utilizing the prior values of each estimator for bootstrapping.
+Consequently, for a target discounted action-value function $Q_{\eta_{m}}(s,a)$, all delta components can be trained concurrently through a TD update, employing the prior values of each estimator for bootstrapping.
 This process necessitates the assumption of a sequence of discount factors, denoted as $\eta_{m}$, which includes both the minimum and maximum values, $\eta_{0}$ and $\eta_{m}$ [1].
@@
 In SARSA($\Delta$), the update Eqs.
-are modified to aggregate rewards across a series of transitions prior to bootstrapping from either the Q-values or delta estimators.
-In multi-step SARSA($\Delta$), the agent aggregates rewards over K steps, leading to a corresponding modification in the update rule for each $D_{m}$.
-$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=1}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{x}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})\right]$ (17)
-In this case, $r_{m+x}$ represents the reward obtained at time step $n+x$.
-$a_{n+k_{m}}$ and the action is chosen greedily as $a=\arg\max_{a}Q(s_{n+k_{m}},a)$, where $k_{m}$ is the number of steps linked to the discount factor $\eta_{m}$.
-This method generalizes Q-learning to multi-step TD learning by aggregating rewards across several stages and utilizing bootstrapping from both the current and prior time-scale action-value functions.
+are adjusted to combine rewards over a sequence of transitions before bootstrapping from either the Q-values or delta estimators.
+In multi-step SARSA($\Delta$), the agent accumulates rewards over K steps, resulting in a necessary adjustment to the update rule for each $D_{m}$.
+$D_{m}(s_{n},a_{n})$ $=\mathbb{E}\biggl{[}\sum_{x=1}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{x}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})$ (16)
+$+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})\biggr{]}$
+In this case, $r_{n+x}$ represents the reward obtained at time step $n+x$.
+The term $a_{n+k_{m}}$ and the action are selected in a greedy manner as $a=\arg\max_{a}Q(s_{n+k_{m}},a)$, with $k_{m}$ representing the number of steps associated with the discount factor $\eta_{m}$.
+This approach extends Q-learning to multi-step temporal difference learning by gathering rewards over multiple stages and employing bootstrapping from both the present and previous time-scale action-value functions.
 In normal multi-step TD learning, the TD error is computed by aggregating rewards over several stages rather than relying on a single future step, followed by bootstrapping from the action-value at the last step.
-For example, utilizing a singular discount factor $\eta$, the multi-step TD update equation is expressed as:
+For instance, employing a single discount factor $\eta$, the multi-step TD update Eq.
+can be articulated as:
 $Q(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k-1}\eta^{x}r_{n+x}+\eta^{k}Q(s_{n+k},a_{n+k})\right]$ (17)
@@
 $D_{m}(s_{n},a_{n})=Q_{\eta_{m}}(s_{n},a_{n})-Q_{\eta_{m-1}}(s_{n},a_{n})$
-We now implement the multi-step variant of the Bellman Eq.
-for both $Q_{\eta_{m}}(s_{n},a_{n})$ and $Q_{\eta_{m-1}}(s_{n},a_{n})$.
+The multi-step variant of the Bellman Eq.
+is now implemented for both $Q_{\eta_{m}}(s_{n},a_{n})$ and $Q_{\eta_{m-1}}(s_{n},a_{n})$.
 The multi-step Bellman Eq.
 for $Q_{\eta_{m}}(s_{n},a_{n})$ is expressed as follows:
-$Q_{\eta_{m}}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k_{m}-1}\eta_{m}^{x}r_{n+x}+\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})\right]$
+$Q_{\eta_{m}}(s_{n},a_{n})=\mathbb{E}\biggl{[}\sum_{x=0}^{k_{m}-1}\eta_{m}^{x}r_{n+x}+\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})\biggr{]}$
 For $Q_{\eta_{m-1}}(s_{n},a_{n})$, the multi-step Bellman Eq. is expressed as follows:
-$Q_{\eta_{m-1}}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k_{m}-1}\eta_{m-1}^{x}r_{n+x}+\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})\right]$
-We will now utilize the delta estimator formulas to subtract these two Eqs.
-$D_{m}(s_{n},a_{n})=\left[\sum_{x=0}^{k_{m}-1}\eta_{m}^{x}r_{n+x}+\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})\right]-\left[\sum_{x=0}^{k_{m}-1}\eta_{m-1}^{x}r_{n+x}+\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})\right]$
-We will now broaden and simplify the expressions.
+$Q_{\eta_{m-1}}(s_{n},a_{n})=\mathbb{E}\biggl{[}\sum_{x=0}^{k_{m}-1}\eta_{m-1}^{x}r_{n+x}+\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})\biggr{]}$
+The formulas for the delta estimator are utilized to perform the subtraction of the two Eqs.
+$\begin{split}D_{m}(s_{n},a_{n})&=\biggl{[}\sum_{x=0}^{k_{m}-1}\eta_{m}^{x}r_{n+x}+\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})\biggr{]}\\ &-\biggl{[}\sum_{x=0}^{k_{m}-1}\eta_{m-1}^{x}r_{n+x}+\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})\biggr{]}\end{split}$
+The expressions are broadened and simplified.
 The immediate reward terms $r_{n+x}$ are present in both Eqs.
@@
 $\sum_{x=0}^{k_{m}-1}\eta_{m}^{x}r_{n+x}-\sum_{x=0}^{k_{m}-1}\eta_{m-1}^{x}r_{n+x}=\sum_{x=0}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}$
-Subsequently, for the bootstrapping terms, we utilize the recursive relationship as outlined:
+Subsequently, for the bootstrapping terms, the recursive relationship is utilized as outlined:
 $Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})=D_{m}(s_{n+k_{m}},a_{n+k_{m}})+Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})$
-As a result, we get:
-$\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})-\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})=(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})$
-We now amalgamate the terms for both the rewards and bootstrapping to produce the final expression:
-$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})\right]$ (19)
-This Eq.
-extends the conventional multi-step TD update to include various multiple discount factors and action-value functions in RL.
+As a result, the following expression is obtained:
+$\begin{split}\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})-\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})&=(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})\\ &+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})\end{split}$
+The definitions for both the rewards and bootstrapping are now integrated to yield the final expression:
+$\begin{split}D_{m}(s_{n},a_{n})&=\mathbb{E}\biggl{[}\sum_{x=0}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})\\ &+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})\biggr{]}\end{split}$ (18)
+Eq.
+18 extends the conventional multi-step TD update to include various multiple discount factors and action-value functions in RL.
 As a result, each $D_{m}$ obtains a share of the rewards from the environment up to time-step $k_{m}-1$.
-Furthermore, each $D_{m}$ utilizes its distinct action-value function in conjunction with the value from the prior time-scale.
+Furthermore, each $D_{m}$ utilizes its distinct action-value function in conjunction with the value from the prior time scale.
 A variation of this algorithm, utilizing k-step bootstrapping as described in [12], is presented in Algorithm 1.
@@
 Set the initial value of $D_{m}(s,a)=0$ for all states, actions, and scales m.
-for episode = 0, 1, 2,… $\triangleright$ Loop for each episode
+[…]
   Initialize state $s_{0}$ and choose an initial action $a_{0}$ in accordance with a policy.
@@
       $G^{0}=\sum_{x=0}^{k_{0}-1}\eta_{0}^{x}r_{n+x}+\eta_{0}^{k_{0}}D_{0}(s_{n+k_{0}},a_{n+k_{0}})$
-              $\triangleright$ Utilizing Eq.10, we substitute $Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})$ by summing the D-components up to $D_{m-1}$ in Eq.4.3.
+[…]
       $G^{m}=\sum_{x=0}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})\sum_{m=0}^{m-1}D_{m-1}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})$
@@
 4.4 SARSA TD($\lambda,\Delta$)
-Eq.20 introduces the $\lambda$-return[23, 1], which combines rewards over multiple steps to create a target for TD($\lambda$) updates. The $\lambda$-return $G_{n}^{\eta,\lambda}$ is defined as:
+Eq.
+19 presents the $\lambda$-return [34, 2], which integrates rewards across several steps to establish a target for TD($\lambda$) updates.
+The $\lambda$-return $G_{n}^{\eta,\lambda}$ is formally defined as follows:
 $G_{n}^{\eta,\lambda}(s_{n},a_{n})=\hat{Q}_{\eta}(s_{n},a_{n})+\sum_{k=0}^{\infty}(\lambda\eta)^{k}\delta_{n+k}^{\eta}$ (19)
-Eq.21 defines the TD($\lambda$) operator, which is used to iteratively apply $\lambda$-returns in updating the value functions.
+Eq.
+20 defines the TD($\lambda$) operator, which is used to iteratively apply $\lambda$-returns in updating the value functions.
 For SARSA, the TD($\lambda$) operator updates the action-value function by summing the $\lambda$-discounted TD errors as follows:
 $\mathrm{T}_{\lambda}Q(s_{n},a_{n})=Q(s_{n},a_{n})+(I-\lambda\eta P)^{-1}(\mathrm{T}Q(s_{n},a_{n})-Q(s_{n},a_{n}))$ (20)
-Here,P is the transition matrix for state-action pairs under the policy $\pi$, and Q is the action-value function.
-Similarly, Eq.22 defines the $\lambda$-return specific to delta estimators in TD($\lambda,\Delta$), denoted as $G_{n}^{m,\lambda_{m}}$ for each delta estimator $D_{m}$:
+P represents the transition matrix for state-action pairs according to the policy $\pi$, while Q denotes the action-value function.
+Similarly, Eq.
+21 defines the $\lambda$-return specific to delta estimators in TD($\lambda,\Delta$), denoted as $G_{n}^{m,\lambda_{m}}$ for each delta estimator $D_{m}$:
 $G_{n}^{m,\lambda_{m}}:=\hat{D}_{m}(s_{n},a_{n})+\sum_{k=0}^{\infty}(\lambda_{m}\eta_{m})^{k}\delta_{n+k}^{m}$ (21)
-where $\delta_{n}^{0}:=\delta_{n}^{\eta_{0}}$ and $\delta_{n}^{m}:=(\eta_{m}-\eta_{m-1})\hat{Q}_{\eta_{m-1}}(s_{n+1},a_{n+1})+\eta_{m}\hat{D}_{m}(s_{n+1},a_{n+1})-\hat{D}_{m}(s_{n+1},a_{n+1})$ are the TD-errors.
+where $\delta_{n}^{0}:=\delta_{n}^{\eta_{0}}$ and $\delta_{n}^{m}:=(\eta_{m}-\eta_{m-1})\hat{Q}_{\eta_{m-1}}(s_{n+1},a_{n+1})+\eta_{m}\hat{D}_{m}(s_{n+1},a_{n+1})\\ -\hat{D}_{m}(s_{n+1},a_{n+1})$ are the TD-errors.
 4.5 SARSA TD($\lambda,\Delta$) with Generalized Advantage Estimation (GAE)
 Generalized Advantage Estimation (GAE) [39] seeks to compute the advantage function by aggregating multi-step TD errors.
-To apply this approach to delta estimators $D_{m}(s_{n},a_{n})$, we compute advantage estimates $A^{\Delta}(s_{n},a_{n})$ for each time-scale.
+To apply this approach to delta estimators $D_{m}(s_{n},a_{n})$, advantage estimates $A^{\Delta}(s_{n},a_{n})$ are computed for each time scale.
 Each advantage estimate $A^{\Delta}(s_{n},a_{n})$ employs a multi-step TD error pertinent to the delta estimator $D_{n}$, expressed as follows:
@@
 where $\delta_{n+k}^{\Delta}:=r_{n}+\eta_{m}\sum_{m=0}^{M}\hat{D}_{m}(s_{n+1},a_{n+1})-\sum_{m=0}^{M}\hat{D}_{m}(s_{n},a_{n})$.
-We utilize $\eta_{m}$ as the discount factor and the sum of all D estimators as a surrogate for $Q_{\eta m}$.
-This objective is applicable to PPO by implementing the policy update from Eq.7 and replacing A with $A^{\Delta}$.
-Additionally, to train each $D_{m}$, we utilize a truncated form of their corresponding $\lambda$-return, as outlined in Eq.22.
+The discount factor $\eta_{m}$ is utilized, and the sum of all D estimators as a surrogate for $Q_{\eta m}$.
+This objective is applicable to PPO by implementing the policy update from Eq.
+7 and replacing A with $A^{\Delta}$.
+Additionally, to train each $D_{m}$, a truncated form of their corresponding $\lambda$-return are used, as outlined in Eq.
+21.
 For more information, see Algorithm 2.
@@
 Initialize policy $\nu$, and values $\theta^{m}\forall m$
-for episode = 0, 1, 2,… $\triangleright$ Loop for each episode
+[…]
   Initialize state $s_{0}$ and choose an initial action $a_{0}$ in accordance with a policy.
@@
      if $n\geq T$ then
-                  $G^{m,\lambda_{m}}\leftarrow\hat{D}_{m}(s_{n-T},a_{n-T})+\sum_{k=0}^{T-1}(\lambda_{m}\eta_{m})^{k}\delta_{n-T+k}^{m}\forall m$triangleright Computing multi-step return $G^{m,\lambda_{m}}$ and TD-error $\delta_{n-T+k}^{m}$ using Eq.22.
+      $G^{m,\lambda_{m}}\leftarrow\hat{D}_{m}(s_{n-T},a_{n-T})+\sum_{k=0}^{T-1}(\lambda_{m}\eta_{m})^{k}\delta_{n-T+k}^{m}\forall m$ $\triangleright$ Computing multi-step return $G^{m,\lambda_{m}}$ and TD-error $\delta_{n-T+k}^{m}$ using Eq.
+21.
      end if
@@
      $\hat{D}_{m}(s_{n-T},a_{n-T})\leftarrow\hat{D}_{m}(s_{n-T},a_{n-T})+\alpha_{m}\left(G^{m,\lambda_{m}}-\hat{D}_{m}(s_{n-T},a_{n-T})\right)$
-              $A^{\Delta}=\sum_{k=0}^{T-1}(\lambda_{m}\eta_{m})^{k}\delta_{n-T+k}^{\Delta}$triangleright Where $A^{\Delta}$ and $\delta_{n-T+k}^{\Delta}$ are computed using Eq.23.
-              $\theta^{m}\leftarrow\theta^{m}+\alpha_{m}\left(G^{m,\lambda_{m}}-\hat{D}_{m}(s_{n-T},a_{n-T})\right)\nabla\hat{D}_{m}(s_{n-T},a_{n-T})$triangleright Update $\theta^{m}$ with TD (Eq.3) using $G^{m,\lambda_{m}}\forall m$.
-              $\mathcal{L}(\nu)=\mathbb{E}\bigg{[}min\bigg{(}\rho(\nu)A^{\Delta}(s,a),clip(\rho(\nu),1-\epsilon,1+\epsilon)A^{\Delta}(s,a)\bigg{)}\bigg{]}$triangleright from Eq.7
-              $\nu\leftarrow\nu+\alpha_{\nu}\nabla_{\nu}\mathcal{L}(\nu)$triangleright Update the policy parameters $\nu$ with PPO (Eq.7) for SARSA using $A^{\Delta}$
+     $A^{\Delta}=\sum_{k=0}^{T-1}(\lambda_{m}\eta_{m})^{k}\delta_{n-T+k}^{\Delta}$ $\triangleright$ Where $A^{\Delta}$ and $\delta_{n-T+k}^{\Delta}$ are computed using Eq. 22.
+     $\theta^{m}\leftarrow\theta^{m}+\alpha_{m}\left(G^{m,\lambda_{m}}-\hat{D}_{m}(s_{n-T},a_{n-T})\right)\nabla\hat{D}_{m}(s_{n-T},a_{n-T})$ $\triangleright$ Update $\theta^{m}$ with TD (Eq.
+3) using $G^{m,\lambda_{m}}\forall m$.
+     $\mathcal{L}(\nu)=\mathbb{E}\bigg{[}min\bigg{(}\rho(\nu)A^{\Delta}(s,a),clip(\rho(\nu),1-\epsilon,1+\epsilon)A^{\Delta}(s,a)\bigg{)}\bigg{]}$ $\triangleright$ from Eq. 7
+     $\nu\leftarrow\nu+\alpha_{\nu}\nabla_{\nu}\mathcal{L}(\nu)$ $\triangleright$ Update the policy parameters $\nu$ with PPO (Eq. 7) for SARSA using $A^{\Delta}$
    end for
@@
 5 Analysis
-Next, We examine our delta estimators through the lens of the bias-variance trade-off.
-In SARSA, bias emerges from dependence on current estimates, affecting updates over short time-scales, particularly with smaller discount factors, such as lower values of $\eta$ in $D_{m}$.
-Updates utilizing smaller discount factors demonstrate reduced variance due to their reliance on current estimates; however, they exhibit increased bias as they fail to integrate substantial information regarding future rewards.
-In contrast, variance escalates for updates associated with extended time-scales, as future reward information adds further variability.
+Subsequently, the delta estimators are analyzed in relation to the bias-variance trade-off.
+In SARSA, bias arises from reliance on current estimates, influencing updates over short time scales, especially with smaller discount factors, like lower values of $\eta$ in $D_{m}$.
+Utilizing smaller discount factors in updates leads to a decrease in variance because they depend on current estimates.
+However, this approach also results in heightened bias, as it does not incorporate significant information about future rewards.
+Conversely, variance increases for updates linked to longer time scales, as future reward information introduces additional variability.
 In TD($\Delta$), action-value components $D_{m}$ with higher $\eta_{m}$ are associated with greater long-term rewards; however, this results in increased variance stemming from the stochastic characteristics of reward sequences.
@@
 This comparison elucidates the essential metrics of our estimator that may indicate potential advantages over the standard $\hat{Q}_{\eta}$ estimator.
-Building on this result and previous research by Kearns and Singh[9], we modify these analyses for SARSA to investigate the effects of bias and variance, along with the TD($\Delta$) decomposition framework, in relation to action-value functions.
+Building on this result and previous research by Kearns and Singh [11], the analyses are adapted for SARSA to investigate the effects of bias and variance, along with the TD($\Delta$) decomposition framework, in relation to action-value functions.
 Our objective is to extend the bias-variance error bound framework to action-value settings using TD($\Delta$)(i.e., Theorem 4), thereby offering a better understanding of how these quantities can be balanced to obtain optimal results [1].
 5.1 SARSA Equivalency Configurations and Enhancements
-In specific situations, we can show that our delta estimators for SARSA align with the conventional action-value estimator when restructured into an action-value function.
-This discussion centers on linear function approximation of the specified form:
+In certain scenarios, we can demonstrate that our delta estimators for SARSA correspond with the traditional action-value estimator when reformulated into an action-value function.
+This discourse focuses on the approximation of the linear function of the specified form:
 $\hat{Q}(s,a)_{\eta}:=\langle\theta^{\eta},\phi(s,a)\rangle\quad\text{and}\quad\hat{D}_{m}(s,a):=\langle\theta^{m},\phi(s,a)\rangle,\forall m$
@@
 $\theta_{n+1}^{\eta}=\theta_{n}^{\eta}+\alpha\bigg{(}G_{n}^{\eta,\lambda}-\hat{Q}_{\eta}(s,a)\bigg{)}\phi(s_{n},a_{n}),$ (23)
-where $G_{n}^{\eta,\lambda}$ represents the TD($\lambda$) return defined in Eq.20. Likewise, the delta estimator approach TD($\lambda_{m},\Delta$) is used to update each $\hat{D}_{m}$:
+Here, $G_{n}^{\eta,\lambda}$ denotes the TD($\lambda$) return as stated in Eq. 19. Likewise, the delta estimator approach TD($\lambda_{m},\Delta$) is used to update each $\hat{D}_{m}$:
 $\theta_{n+1}^{m}=\theta_{n}^{m}+\alpha\bigg{(}G_{n}^{M,\lambda_{m}}-\hat{D}_{m}(s_{n},a_{n})\bigg{)}\phi(s_{n},a_{n}),$ (24)
-where $G_{n}^{M,\lambda_{m}}$ is specified identically to the TD($\Delta$) return defined in Eq.22, modified for the particular action-value function.
+where $G_{n}^{M,\lambda_{m}}$ is specified identically to the TD($\Delta$) return defined in Eq.
+21, modified for the particular action-value function.
 In these equations, $\alpha$ and $\{\alpha_{m}\}_{m}$ represent positive learning rates.
-The following theorem demonstrates the equivalence of two SARSA version algorithms in this context.
-The subsequent theorems resemble those presented by Romoff et al.
-[8]; however, we offer a proof concerning the action-value function in relation to SARSA, whereas they established the proof for the value function.
+Two SARSA version algorithms are shown to be equivalent in this context by the following theorem.
+The subsequent theorems resemble those presented by Romoff et al.[1]; however, a proof is presented for the action-value function in relation to SARSA, whereas Romoff et al.
+established the proof for the value function.
 Theorem 1.
 Inspired by Romoff et al.[1], if $\alpha_{m}=\alpha$, $\lambda_{m}\eta_{m}=\lambda_{\eta}$, $\forall m$, and if we choose the initial conditions in such a way that $\sum_{m=0}^{M}\theta_{0}^{\eta}=\theta_{0}^{\eta}$, then the iterates produced by TD($\lambda$) (Eq.
-24) and TD($\lambda,\Delta$) (Eq.
-25) with linear function approximation satisfy the following conditions:
+23) and TD($\lambda,\Delta$) (Eq.
+24) with linear function approximation satisfy the following conditions:
 $\Sigma_{m=0}^{M}\theta_{n}^{m}=\theta_{n}^{\eta}$ (25)
-The proof is presented in appendix PROOFS.
+The proof is presented in appendix A.
 Equivalence is attained when $\lambda_{m}\eta_{m}=\lambda_{\eta},\forall m$.
 When $\lambda$ approaches 1 and $\eta_{m}$ is less than $\eta$, this condition suggests that $\lambda_{m}=\lambda_{\eta}/\eta_{m}$ could potentially surpass one, resulting in a risk of divergence in TD($\lambda_{m}$).
-The subsequent theorem demonstrates that the TD($\lambda$) operator, as defined in Eq.21, qualifies as a contraction mapping for the range $1\leq\lambda<\frac{1+\eta}{2\eta}$, thereby confirming that $\lambda_{\eta}<1$[8].
+The subsequent theorem demonstrates that the TD($\lambda$) operator, as defined in Eq.
+20, qualifies as a contraction mapping for the range $1\leq\lambda<\frac{1+\eta}{2\eta}$, thereby confirming that $\lambda_{\eta}<1$ [1].
 Theorem 2.
 Inspired by Romoff et al.[1], $\forall\lambda\in\left[0,\frac{1+\eta}{2\eta}\right]$, the operator $T_{\lambda}$ is defined by the Eq.
-$T_{\lambda}Q=Q+(I-\lambda\eta P)^{-1}(TQ-Q)$ $,\quad\forall Q\in\mathbb{R}^{|S|\times|A|}$, is well-defined.
+$T_{\lambda}Q=Q+(I-\lambda\eta P)^{-1}(TQ-Q)$ $\quad\forall Q\in\mathbb{R}^{|S|\times|A|},$ is well-defined.
 Furthermore, $T_{\lambda}Q$ constitutes a contraction in relation to the max norm, with its contraction coefficient expressed as $\frac{\eta}{|1-\lambda\eta|}$.
-The proof is presented in appendix PROOFS.
+The proof is presented in appendix A.
 The analysis presented in Theorem 1 can be adapted to a new context in which $k_{m}$-step TD ($\Delta$) is employed for each $D_{m}$ component, rather than TD($\lambda$, $\Delta$), a variant of TD learning.
@@
 Although these methods could theoretically be equivalent, such equivalence is not common.
-Generally, for equivalence to be maintained, the learning rate must remain consistent across all time scales.
-This represents a significant limitation, as shorter time scales, which involve fewer future steps, can be learned more rapidly than longer time scales that require consideration of additional steps ahead.
+To maintain the preservation of equivalence, it is crucial for the learning rate to remain consistent across all time scales.
+This underscores a significant limitation, as shorter time scales, which involve fewer future steps, can be acquired more rapidly than longer time scales that require the assessment of additional steps beforehand.
 Furthermore, in practical applications, particularly involving nonlinear function approximation such as deep neural networks, adaptive optimizers are frequently employed [40, 39].
-These optimizers modify the learning rate in accordance with the complexity of the learning task, which is contingent upon the characteristics of the delta estimator and its target.
-The effective learning rate varies across different time scales, making it challenging to apply a uniform learning rate in this context.
-Besides the learning rate, the method of dissecting the action-value function into distinct components (particularly, $D_{m}$ components) offers benefits that conventional TD learning, including non-delta estimators or elevated $D_{m}$ components, does not possess.
-The capacity to employ diverse k-step returns (or $\lambda$-return) across many time-scales offers improved control and flexibility in the learning process.
-If $k_{m}<k_{m+1},\forall m\quad(or\quad\eta_{m}\lambda_{m}<\eta_{m+1}\lambda_{m+1},\forall m)$, this approach may decrease variance while introducing bias, as shorter time-scales, which can be learned more rapidly, may not align perfectly with longer time scales in terms of the values they estimate.
+The optimizers modify the learning rate accordance with the complexity of the learning task, which depends on the attributes of the delta estimator and its target.
+There is no universally applicable learning rate in this setting since the effective learning rate changes with time scale.
+In addition to the learning rate, breaking the action-value function into separate components (especially the $D_{m}$ components) has advantages that standard TD learning, such as non-delta estimators, does not have.
+The ability to utilize various k-step returns (or $\lambda$-return) across multiple time scales enhances control and adaptability in the learning process.
+If $k_{m}<k_{m+1},\forall m\quad(or\quad\eta_{m}\lambda_{m}<\eta_{m+1}\lambda_{m+1},\forall m)$, this approach has the potential to reduce variance while simultaneously introducing bias, since shorter time scales, which can be learned more quickly, may not correspond perfectly with longer time scales regarding the values they estimate.
 5.2 Evaluation for Minimizing $k_{m}$ Values in SARSA via Phased Updates
-To show that how our approach differ from the single estimator case, let’s follow the tabular phased version of k-step TD presented by kearns & Singh[9].
-In SARSA, we aim to estimate the value of a state-action pair (s, a) start from each state $s\in\mathcal{S}$ and with action $a\in\mathcal{A}$, then n trajectories generated $\bigg{\{}S_{0}^{(x)}=s,a_{0},\eta_{0},...,S_{k}^{(x)},a_{k}^{(x)},\eta_{k}^{(x)},S_{k+1}^{(x)},...\bigg{\}}_{1\leq x\leq n}$ by following the policy $\pi$ and averaging over trajectories.
+To show how our approach differs from the single estimator case, let’s follow the tabular phased version of k-step TD presented by Kearns and Singh[11].
+In SARSA, the goal is to evaluate the value associated with a state-action pair (s, a) starting from each state $s\in\mathcal{S}$ and with action $a\in\mathcal{A}$, then n trajectories are generated $\bigg{\{}S_{0}^{(x)}=s,a_{0},\eta_{0},...,S_{k}^{(x)},a_{k}^{(x)},\eta_{k}^{(x)},S_{k+1}^{(x)},...\bigg{\}}_{1\leq x\leq n}$ by following the policy $\pi$ and averaging over trajectories.
 Let $Q_{\eta,n}(s,a)$ denote the Q-value estimate at phase n for the discount factor $\eta$.
-For each iteration n, also called phase n, to get the phase-based estimate $Q_{\eta,n}(s,a)$ for (s, a), we average over the trajectories, resulting in:
+For each iteration n, also called phase n, obtain the phase-based estimate $Q_{\eta,n}(s,a)$ for (s, a); we average over the trajectories, resulting in:
 $\hat{Q}_{\eta,n}(s,a)=\frac{1}{n}\sum_{x=1}^{n}\bigg{(}\sum_{i=0}^{k-1}\gamma^{i}r_{i}^{(x)}+\eta^{k}\hat{Q}_{\eta,n-1}(s_{k}^{(x)},a_{k}^{(x)})\bigg{)}$ (26)
-Theorem 3 below from the theorems of kearns & Singh[9] but they show it for state value function while we show an upper bound on the error for action-value function estimation defined by $\Delta_{n}^{\hat{Q}_{\eta}}=max_{s,a}\bigg{\{}\bigg{|}\hat{Q}_{\eta,n}(s,a)-Q_{\eta}(s,a)\bigg{|}\bigg{\}}$
+Theorem 3 draws upon the findings of Kearns and Singh[11], who demonstrated the result concerning the state-value function.
+In contrast, the subsequent theorem introduces an upper limit on the error associated with action-value function estimation, expressed as $\Delta_{n}^{\hat{Q}_{\eta}}=max_{s,a}\bigg{\{}\bigg{|}\hat{Q}_{\eta,n}(s,a)-Q_{\eta}(s,a)\bigg{|}\bigg{\}}$
 Theorem 3.
-Inspired by Kearns & Singh[9], for any $10<\delta<$ 1, let $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$. with probability $1-\delta$,
+Inspired by Kearns and Singh[11], for any $0<\delta<1$, let $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$. with probability $1-\delta$,
 $\Delta_{n}^{\hat{Q}_{\eta}}\leq\epsilon\bigg{(}\underbrace{\frac{1-\eta^{k}}{1-\eta}\bigg{)}}_{\text{variance term}}+\underbrace{\eta^{k}\Delta_{n-1}^{\hat{Q}_{\eta}}}_{\text{bias term}}$ (27)
-The proof is presented in appendix PROOFS.
+The proof is presented in appendix A.
 In Eq.
-28, A variance term ($\frac{1-\eta^{k}}{1-\eta}$) is arising due to sampling error from rewards collected along trajectories.
+27, a variance term ($\frac{1-\eta^{k}}{1-\eta}$) is arising due to sampling error from rewards collected along trajectories.
 In particular, $\epsilon$ bounds the deviation of the empirical average of rewards from the true expected reward.
-Second term ($\eta^{k}\Delta_{n-1}^{\hat{Q}_{\eta}}$) in Eq.
+The second term ($\eta^{k}\Delta_{n-1}^{\hat{Q}_{\eta}}$) in Eq.
 27 is a bias term that arises from bootstrapping off the previous phase’s estimates.
-Similarly, we consider a phased SARSA version of multi-step $TD(\Delta)$.
-For each phase n, we update each D as follows:
-$\hat{D}_{m,n}(s,a)=\frac{1}{n}\sum_{x=1}^{n}\bigg{(}\sum_{i=1}^{k-1}(\eta_{m}^{i}-\eta_{m-1}^{i})r_{i}^{(x)}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k}^{(x)},a_{n+k}^{(x)})+\eta_{m}^{k_{m}}\hat{D}_{m}(s_{n+k}^{(x)},a_{n+k}^{(x)})$ (29)
-We will now ascertain the highest limit of the phased error.
+Similarly, a phased SARSA variant of the multi-step $TD(\Delta)$ method is considered. For each phase n, each D component is updated as follows:
+$\hat{D}_{m,n}(s,a)=\frac{1}{n}\sum_{x=1}^{n}\biggl{(}\sum_{i=1}^{k-1}(\eta_{m}^{i}-\eta_{m-1}^{i})r_{i}^{(x)}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k}^{(x)},a_{n+k}^{(x)})+\eta_{m}^{k_{m}}\hat{D}_{m}(s_{n+k}^{(x)},a_{n+k}^{(x)})\biggr{)}$ (28)
+The maximum threshold for the phased error has now been established.
 The cumulative errors of each D component, denoted as $TD(\Delta)$, is expressed as $\sum_{m=0}^{M}\Delta_{n}^{m}$, where $\Delta_{n}^{m}=\max_{s,a}\bigg{\{}\bigg{|}\hat{D}_{m}(s,a)-D_{m}(s,a)\bigg{|}\bigg{\}}$.
-Fundamental differences between SARSA and Q-learning regarding the error bound: In SARSA, the error is ascertained by the agent’s actual actions, indicating that the Q-values are modified based on the state-action pairings encountered by the agent during exploration.
-The inaccuracy in Q-learning arises from the greedy action, defined as the action that maximizes the Q-value for the subsequent state, based on the premise that the agent consistently behaves optimally.
+Essential differences between SARSA and Q-learning regarding the error bound: In SARSA, the error is defined by the actions taken by the agent, indicating that the Q-values are modified based on the state-action pairings encountered by the agent during exploration.
+The inaccuracy in Q-learning emerges from the greedy action, which is defined as the action that maximizes the Q-value for the future state, assuming that the agent consistently operates in an optimal manner.
 Theorem 4.
-Inspired by Romoff et al.[8], assume that $\eta_{0}\leq\eta_{1}\leq\eta_{2}\leq.....\leq\eta_{m}=\eta$ and $K_{0}\leq K_{1}\leq...\leq K_{m}=K$, for any $10<\delta<$ 1, let $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$, with probability $1-\delta$,
+Inspired by Romoff et al.[1], assume that $\eta_{0}\leq\eta_{1}\leq\eta_{2}\leq.....\leq\eta_{m}=\eta$ and $K_{0}\leq K_{1}\leq...\leq K_{m}=K$, for any $0<\delta<1$, let $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$, with probability $1-\delta$,
 $\sum_{m=0}^{M}\Delta_{n}^{m}\leq\epsilon\bigg{(}\frac{1-\eta^{k}}{1-\eta}\bigg{)}+\epsilon\bigg{(}\underbrace{\sum_{m=0}^{M-1}\frac{\eta_{m}^{k_{m+1}}-\eta^{k_{m}}_{m}}{1-\eta_{m}}\bigg{)}}_{\text{variance reduction}}+\underbrace{\sum_{m=0}^{M-1}\bigg{(}\eta^{k_{m}}_{m}-\eta^{k_{m+1}}_{m}\bigg{)}\sum_{q=0}^{m}\Delta_{n-1}^{q}}_{\text{bias introduction}}+\eta^{k}\sum_{m=0}^{M}\Delta_{n-1}^{m}$ (29)
-The proof is presented in appendix PROOFS.
-The author examines the comparison of the limits for phased TD($\lambda$) in theorem 3 with those for phased TD($\Delta$) in theorem 4.
+The proof is presented in appendix A.
+The authors examine the comparison of the limits for phased TD($\lambda$) in theorem 3 with those for phased TD($\Delta$) in theorem 4.
 This comparison demonstrates that phased TD($\Delta$) facilitates variance reduction equivalent to $\epsilon\sum_{m=0}^{M-1}\frac{\eta_{m}^{k_{m+1}}-\eta^{k_{m}}_{m}}{1-\eta_{m}}\leq 0$ but introduces a potential bias quantified as $\sum_{m=0}^{M-1}\bigg{(}\eta^{k_{m}}_{m}-\eta^{k_{m+1}}_{m}\bigg{)}\sum_{q=0}^{m}\Delta_{n-1}^{q}\geq 0$.
@@
 It is observed that when all $k_{m}$ values are identical, both algorithms yield the same upper bound.
-In SARSA, the return estimate encompasses subsequent states and actions dictated by the policy.
-The expected discounted return over T steps closely approximates the infinite-horizon discounted expected return after $\mathrm{T}$, where $\mathrm{T}\approx\frac{1}{1-\eta}$[31].
-Consequently, we can effectively simplify $k_{m}$ for any $\eta_{m}$ such that $k_{m}\approx\frac{1}{1-\eta_{m}}$, thereby adhering to this principle.
+In SARSA, the return estimate includes subsequent states and actions determined by the policy.
+The anticipated discounted return over T steps closely approximates the infinite-horizon discounted expected return after $\mathrm{T}$, where $\mathrm{T}\approx\frac{1}{1-\eta}$ [41].
+Consequently, $k_{m}$ can be effectively simplified for any $\eta_{m}$ such that $k_{m}\approx\frac{1}{1-\eta_{m}}$, thereby adhering to this principle.
 Therefore, with $\mathrm{T}$ samples, similar to TD($\Delta$) [1], SARSA employing TD($\Delta$) utilizes $k_{m}\approx\frac{1}{1-\eta_{m}}$ to establish a suitable balance between bias and variance over all time scales significantly less than $\mathrm{T}$.
@@
 SARSA can utilize this to modify additional parameters in alignment with the specified time scales, enhancing the algorithm’s efficiency in calculating long-term rewards without added complexity.
-6 Experimental Setup
-We experimented with OpenAI Gym control tasks, and Fig.1 shows the relevant environments.
+6 Experiments and Results
+Experiments were conducted in three distinct environments to evaluate the effectiveness of the proposed method: the Tabular Ring MDP environment detailed in Section 6.1, the deterministic and stochastic OpenAI Gym environments examined in Section 6.2, and the Atari environments from OpenAI Gym described in Section 6.3.
+This section initiates with a detailed overview of the experimental setup.
+Experimental Setup: Experiments were conducted using OpenAI Gym control tasks [42], and Figs.1, 3, and 6 show the relevant environments.
 The experiments were carried out on a single PC with 16 GB of memory, an Intel Core i7-2600 processor, and a GPU.
-The system ran 64-bit Ubuntu 22.04.5 LTS.
-We utilized Python 3.9 and the PyTorch package for development.
-We also used the OpenAI Gym toolkit, as mentioned by G.
-Brockman et al.
-[brockman2016openai].
-Ten randomly selected seeds were used in each experiment, and the average outcomes are presented.
-(a) FrozenLake-v1
-(b) Taxi-v3
-(c) Blackjack-v1
-(d) CliffWalking-v0
-Figure 1: OpenAI Gym control used for all experiments. In order from the left to right: Frozen Lake, Taxi, Blackjack and Cliff Walking.
+The machine ran on 64-bit Ubuntu 22.04.5 LTS.
+Python 3.9 was used during development, along with the PyTorch package.
+In each experiment, ten randomly selected seeds were utilized, and the average results are provided.
+The remaining hyperparameter settings are displayed in the table 4 of appendix B.
 6.1 Tabular
-We utilize Q-value iteration to determine the optimal Q-value function.
-We employ an open-source version of Value Iteration (VI) derived from the code of Peter Henderson and Wei-Di Chang[value-iteration], guided by the algorithmic principles outlined in Bellman[22] and Sutton & Barto[29].
-Although their methodology calculates the optimal value function, we modify it to determine the optimal Q-value function.
-Upon acquiring the optimal Q-values, we employ Temporal Difference (TD) and SARSA($\Delta$) to derive approximation Q-values via sampling.
+Q-value iteration is utilized to determine the optimal Q-value function.
+An open-source implementation of Value Iteration (VI), based on the code by Peter Henderson[43], is employed in accordance with the algorithmic principles outlined in Bellman[33] and Sutton et al.[12].
+The methodology calculates the optimal value function; however, this approach is adjusted to ascertain the optimal Q-value function.
+Upon acquiring the optimal Q-values, Temporal Difference (TD) and SARSA($\Delta$) methods are applied to derive approximation Q-values via sampling.
 After 5000 steps, each learning iteration comes to an end.
-For every setup, we run 200 trials using different random seeds that influence the dynamics of the transition.
-We implement our algorithms using the code that Kostrikov provided[pytorchrl].
-We compare our approach to k-step SARSA, as described in Sutton and Barto’s book, Chapter 7.2[29].
-We evaluate SARSA($\Delta$) on multiple RL benchmarks, such as the Ring MDP[9] and deterministic & stochastic environments, to determine its effectiveness.
-A illustration of the Ring MDP is shown in Fig.
-2, which shows a Markov process with five states arranged in a circle.
-There is a 0.95 chance of moving one state clockwise around the ring and 0.05 chance of staying in the present state at each step.
-While the other states on the ring offer a reward of 0, two adjacent states offer rewards of +1 and -1, respectively.
-Figure 2: Five-state ring MDP as described by Kearns[9]
+For every setup, 200 trials are executed using different random seeds that influence the dynamics of the transition.
+The algorithms are implemented using the codebase provided by Kostrikov[44].
+The proposed method is contrasted with k-step SARSA, as described in Sutton et al.[12, Chapter 7.2].
+SARSA ($\Delta$) is evaluated against different RL benchmarks, including the Ring MDP [11] and both deterministic and stochastic environments, to determine its efficacy.
+Fig.
+1 illustrates the Ring MDP, which consists of a Markov process featuring five states arranged in a circular form.
+The probability of transitioning one state clockwise around the ring is 0.95, while the likelihood of remaining in the current state at each step is 0.05.
+In contrast to the other states on the ring that provide a reward of 0, two neighboring states present rewards of +1 and -1, respectively.
+Figure 1: Five-state ring MDP as described by Kearns and Singh[11]
 All experiments utilize the identical parameters established by Romoff et al.[1], which involve supplying a variable number of gamma values, commencing at 0 and incrementing according to $\eta_{m+1}=\frac{\eta_{m}+1}{2}$ until the requisite maximum $\eta_{m}$ is attained.
@@
 Fig.
-3(a) depicts the efficacy of two methodologies: K-step SARSA ($\Delta$) and K-step SARSA TD, employing a discount factor ($\eta$ = 0.996), k = 64 (a parameter associated with the algorithm’s step count), and diverse values of the learning rate $\alpha$.
-The y-axis denotes the absolute error, whereas the x-axis signifies the parameter $\alpha$, ranging from 0.0 to 1.0.
-At $\alpha$=0.0, both methodologies commence at an equivalent error level (about 0.00230).
-As $\alpha$ increases, K-step SARSA ($\Delta$) initially diminishes its inaccuracy marginally before seeing a consistent rise.
-K-step SARSA TD, conversely, exhibits a consistent increase in error with minimal fluctuation.
-Overall course: Relative to K-step SARSA TD, the K-step SARSA ($\Delta$) technique has greater stability and consistently yields superior outcomes (i.e., reduced error) across all values of $\alpha$.
-Both approaches exhibit sensitivity to variations in $\alpha$, although K-step SARSA TD demonstrates greater sensitivity, evidenced by its more pronounced increase in error.
+2a illustrates the effectiveness of two methodologies: K-step SARSA ($\Delta$) and K-step SARSA TD, utilizing a discount factor ($\eta$ = 0.996), k = 64 (a parameter related to the algorithm’s step count), and various values of the learning rate $\alpha$.
+The y-axis represents the absolute error, while the x-axis indicates the parameter $\alpha$, which ranges from 0.0 to 1.0.
+At $\alpha$=0.0, both methods start with about the same amount of error (around 0.00230).
+As $\alpha$ increases, K-step SARSA ($\Delta$) exhibits a slight initial reduction in inaccuracy, followed by a consistent rise thereafter.
+In contrast, K-step SARSA TD shows a constant rise in error with low variability.
+Overall, the K-step SARSA ($\Delta$) approach outperforms K-step SARSA TD in terms of stability and accuracy across all $\alpha$ values.
+Both techniques are sensitive to fluctuations in $\alpha$, but K-step SARSA TD has a more significant rise in error.
 Fig.
-3(b) demonstrates the effectiveness of the two methodologies, K-step SARSA ($\Delta$) and K-step SARSA TD, in relation to the number of k-steps.
-The y-axis indicates the absolute error, whereas the x-axis represents k-steps, which refers to the number of steps utilized in the algorithms.
-The range of k-steps extends from 0 to roughly 250.
-K-step SARSA ($\Delta$) commences with a reduced error at smaller k-steps, which subsequently increases with bigger k-steps.
-K-step SARSA TD begins with an error level comparable to K-step SARSA ($\Delta$) for low k-steps; however, its error escalates more steeply as k-steps grow.
-The overall pattern indicates that both techniques exhibit escalating inaccuracy as k-steps rise.
-K-step SARSA ($\Delta$) consistently surpasses K-step SARSA TD in all k-step values by exhibiting a reduced error rate.
-The swift error amplification in K-step SARSA TD indicates that the scheduling method employed in K-step SARSA ($\Delta$) mitigates error accumulation.
-Furthermore, in this specific MDP, the error escalates with k due to the bias-variance trade-off associated with k-step returns, aligning with the conclusions of Kearns & Singh[9, 8].
-Comparison of Figs.
-3(a) and 3(b): In the comparison of the two figures, K-step SARSA ($\Delta$) consistently surpasses K-step SARSA TD in both circumstances ($\alpha$ and k-steps).
-Both approaches exhibit sensitivity to elevated parameter values ($\alpha$ or k-steps).
-K-step SARSA ($\Delta$) exhibits enhanced stability and attains reduced errors as a result of its scheduling approach.
-Significantly, errors escalate more rapidly with K-step SARSA TD as k-steps grow in comparison to the increase in $\alpha$.
-(a) Error at different learning rates
-(b) Error at different k values
-Figure 3: Fig.3(a) illustrates the outcomes for $\eta_{M}=0.996$ on the five-state Ring MDP.
+2b compares the efficacy of two approaches, K-step SARSA ($\Delta$) and K-step SARSA TD, based on the number of k-steps.
+The y-axis represents the absolute error, whereas the x-axis depicts k-steps, which relates to the number of steps utilized in the algorithms.
+The range of k-steps spans from 0 to approximately 250.
+K-step SARSA ($\Delta$) begins with a lower error at smaller k-steps, which then escalates with larger k-steps.
+K-step SARSA TD starts with an error level similar to that of K-step SARSA ($\Delta$) for lower k-steps; however, its error increases more sharply as k-steps increase.
+The overall pattern indicates that both techniques exhibit a rise in inaccuracy as the number of k-steps increases.
+K-step SARSA ($\Delta$) consistently demonstrates superior performance compared to K-step SARSA TD across all k-step values, evidenced by a lower error rate.
+The rapid error amplification noted in K-step SARSA TD indicates that the scheduling method employed in K-step SARSA ($\Delta$) successfully mitigates error accumulation.
+Additionally, in this particular MDP, the error increases with k as a result of the bias-variance trade-off linked to k-step returns, consistent with the findings of Romoff et al.[1], Kearns and Singh[11].
+Analysis of Figs.
+2a and 2b: In comparing the two figures, K-step SARSA ($\Delta$) consistently outperforms K-step SARSA TD in both scenarios ($\alpha$ and k-steps).
+Both approaches demonstrate sensitivity to increased parameter values ($\alpha$ or k-steps).
+K-step SARSA ($\Delta$) demonstrates improved stability and achieves lower error rates due to its scheduling methodology.
+Interestingly, compared to the growth in $\alpha$, errors climb faster with K-step SARSA TD as the number of k-steps increases.
+(a) Error at different learning rates.
+(b) Error at different k values.
+Figure 2: Fig.
+2a illustrates the outcomes for $\eta_{M}=0.996$ on the five-state Ring MDP.
 The error signifies the absolute discrepancy between the estimated and actual discounted Q-value function (precomputed via Q-value iteration), averaged across the complete learning trajectory of 5000 timesteps.
-Fig.3(b) illustrates the average absolute error for the optimal learning rate at each k-step return, extending beyond the effective planning horizon of $\eta_{M}$.
+Fig.
+2b illustrates the average absolute error for the optimal learning rate at each k-step return, extending beyond the effective planning horizon of $\eta_{M}$.
 6.2 OpenAI Gym Environments
-The following plots show the results of two algorithms, k-step SARSA($\Delta$) and k-step SARSA TD, on the FrozenLake-v1, Taxi-v3, CliffWalking-v0, and Blackjack environments from OpenAI Gym.
-The y-axis shows the average rewards obtained by the algorithms over episodes, and the x-axis shows the number of episodes needed to train the algorithms.
+(a) FrozenLake-v1.
+(b) Taxi-v3.
+(c) Blackjack-v1.
+(d) CliffWalking-v0.
+Figure 3: OpenAI Gym control used for all experiments. In order from left to right: Frozen Lake, Taxi, Blackjack, and Cliff Walking.
+Two algorithms, k-step SARSA($\Delta$) and k-step SARSA TD, were tested in the FrozenLake-v1, Taxi-v3, CliffWalking-v0, and Blackjack environments of OpenAI Gym.
+The results are shown in the figures below.
+The y-axis displays how many awards the algorithms got on average across episodes, while the x-axis shows how many episodes it took to train the algorithms.
 Episodes range from 0 to 100,000 episodes.
-The shaded area surrounding each curve shows the standard error over 10 random seeds.
-The algorithms both use k= 32, which is the number of steps.
-The remaining hyperparameter settings are displayed in the table 1.
-Table 1: Hyperparameter configurations and their corresponding values.
-Hyperparameter Value
-seed (10) np.random.randint(40, 45)
-Learning Rate ($\alpha$) [0.1, 0.3, 0.5, 0.7, 1.0]
-Discount Factor ($\eta$) [0.992, 0.95, 0.996, 0.99, 1.0]
-k-Step [1, 4, 8, 16, 32, 128, 256]
-Number of Episodes 100000
-Max Epsilon 0.4
+The shaded area around each curve displays the standard error for 10 random seeds.
+Both algorithms utilize k = 32, which is the number of steps.
 The red line denotes K-step SARSA ($\Delta$), while the blue line represents K-step SARSA TD in both Fig.
-4(a) and Fig.
-4(b).
+4a and Fig.
+4b.
 Fig.
-4(a) illustrates the FrozenLake-v1 environment, whereas Fig.
-4(b) depicts the Taxi-v3 environment.
+4a presents the FrozenLake-v1 environment, while Fig.
+4b showcases the Taxi-v3 environment.
 Fig.
-4(a) demonstrates that k-step SARSA ($\Delta$) continually attains superior average rewards relative to k-SARSA TD as training advances.
-K-step SARSA ($\Delta$) demonstrates considerable enhancement in average rewards over the initial 20,000 episodes, signifying accelerated learning, but k-step SARSA TD exhibits lesser increase in average rewards in comparison to k-step SARSA ($\Delta$) during the same episodes.
-Following 40,000 episodes, the efficacy of k-step SARSA($\Delta$) stabilizes while consistently surpassing that of k-step SARSA TD, indicating enduring reliability.
-k-step SARSA regularly exhibits inferior performance relative to k-step SARSA ($\Delta$).
+4a illustrates that k-step SARSA ($\Delta$) consistently achieves higher average rewards compared to k-SARSA TD as training progresses.
+Over the first 20,000 episodes, K-step SARSA ($\Delta$) shows fast learning with a significant rise in average rewards.
+However, k-step SARSA TD displays a smaller increase in average rewards when compared to k-step SARSA ($\Delta$) during the same episodes.
+After 40,000 episodes, k-step SARSA ($\Delta$) performance levels off and regularly beats k-step SARSA TD, showing that it is reliable over the long-term.
+k-step SARSA consistently performs worse than k-step SARSA ($\Delta$).
 Fig.
-4(b) illustrates that K-Step SARSA ($\Delta$) regularly achieves superior rewards (i.e., nearer to -200) during the training phase, signifying enhanced learning efficiency and overall performance.
+4b illustrates that K-Step SARSA ($\Delta$) regularly achieves superior rewards (i.e., nearer to -200) during the training phase, signifying enhanced learning efficiency and overall performance.
 Fig.
-4(b) also demonstrates that K-Step SARSA ($\Delta$) stabilizes in the initial phases, indicating robustness and the capacity to converge to an effective policy, whereas K-Step SARSA TD commences at considerably lower rewards (approximately -1800) and exhibits negligible improvement, remaining predominantly flat throughout the training process.
-The K-Step SARSA TD algorithm encounters difficulties in acquiring an effective policy in this setting compared to K-Step SARSA ($\Delta$).
-Experiment result show that K-Step SARSA ($\Delta$) significantly outperforms K-Step SARSA TD in the Taxi-v3 environment, achieving much higher average rewards and faster convergence.
-(a) FrozenLake-v1
-(b) Taxi-v3
+4b also demonstrates that K-Step SARSA ($\Delta$) stabilizes in the initial phases, indicating robustness and the capacity to converge to an effective policy, whereas K-Step SARSA TD commences at considerably lower rewards (approximately -1800) and exhibits negligible improvement, remaining predominantly flat throughout the training process.
+The K-Step SARSA TD algorithm encounters difficulties in formulating an effective policy in this context relative to K-Step SARSA ($\Delta$).
+The results of the experiment indicate that K-Step SARSA ($\Delta$) demonstrates a significant advantage over K-Step SARSA TD in the Taxi-v3 environment, attaining considerably higher average rewards and achieving faster convergence.
+(a) FrozenLake-v1.
+(b) Taxi-v3.
 Figure 4: Comparative analysis of K-step SARSA ($\Delta$) versus traditional K-step SARSA TD within the OpenAI Gym environments: FrozenLake-v1 and Taxi-v3.
 The shaded area illustrates the standard error calculated from 10 random seeds.
-(a) CliffWalking-v0
-(b) Blackjack-v1
+(a) CliffWalking-v0.
+(b) Blackjack-v1.
 Figure 5: Comparative analysis of K-step SARSA ($\Delta$) versus traditional K-step SARSA TD within the OpenAI Gym environments: CliffWalking-v0 and Blackjack-v1.
 The shaded area illustrates the standard error calculated from 10 random seeds.
-The Fig.
-5(a) and 5(b) illustrates the learning effectiveness of two variations of the K-Step SARSA algorithm, namely K-Step SARSA ($\Delta$) and K-Step SARSA TD, across 100,000 episodes in the CliffWalking-v0 and Blackjack-v1 environments.
-Both algorithms are assessed using k=32, a hyperparameter that specifies the number of steps in the updating process.
-The trend of average rewards in fig.
-5(a) demonstrates that both algorithms show an increasing pattern in average rewards as the number of episodes rises, indicating improved policy performance with continued learning.
-Nevertheless, K-Step SARSA ($\Delta$) often outperforms K-Step SARSA TD, achieving higher average rewards over the episodes.
-The red line in fig.
-5(a) denoting K-Step SARSA ($\Delta$) demonstrates a considerable enhancement in the initial episodes (about 20,000 episodes), indicating a swifter convergence to optimal policies relative to K-Step SARSA TD.
-In the 100,000th episode, K-Step SARSA ($\Delta$) attains a reward near -1050, whereas K-Step SARSA TD obtains a slightly lower payout of approximately -1120.
-This demonstrates the enhanced efficacy of K-Step SARSA ($\Delta$) in this task.
+Figs.
+5a and 5b demonstrate the learning effectiveness of two variations of the K-Step SARSA algorithm: K-Step SARSA ($\Delta$) and K-Step SARSA TD.
+This analysis is based on 100,000 episodes conducted in the CliffWalking-v0 and Blackjack-v1 environments.
+The evaluation of both algorithms is conducted with k=32, which is a hyperparameter indicating the number of steps involved in the updating process.
+The overall pattern of average rewards shown in Fig.
+5a indicates that both algorithms exhibit an upward trajectory in average rewards as the number of episodes increases, reflecting enhanced policy performance through ongoing learning.
+Nonetheless, K-Step SARSA ($\Delta$) routinely exhibits better performance than K-Step SARSA TD, leading to increased average rewards across the episodes.
+In comparison to K-Step SARSA TD, the red line in Fig.
+5a representing K-Step SARSA ($\Delta$) shows a significant improvement in the first episodes (about 20,000 episodes), suggesting a faster convergence to optimum policies.
+K-Step SARSA ($\Delta$) receives a reward close to -1050 in the 100,000th episode, whereas K-Step SARSA TD receives a somewhat smaller reward of around -1120.
+This illustrates how K-Step SARSA ($\Delta$) performs better on this task.
 The average reward of approximately +0.15 is dependably achieved across all episodes by the red curve in Fig.
-5(b) that represents K-Step SARSA ($\Delta$).
+5b that represents K-Step SARSA ($\Delta$).
 The blue curve in Fig.
-5(b), which represents K-Step SARSA TD, suggests stability; however, it presents an average reward that is significantly lower, at approximately -0.2, across all episodes.
+5b, which represents K-Step SARSA TD, suggests stability; however, it presents an average reward that is significantly lower, at approximately -0.2, across all episodes.
 K-Step SARSA ($\Delta$) markedly surpasses K-Step SARSA TD regarding average rewards.
 The favorable average reward of K-Step SARSA ($\Delta$) demonstrates its superior suitability for policy optimization in the Blackjack-v1 environment, whereas the negative reward for K-Step SARSA TD signifies inferior performance.
-Finally, when comparing the results of all four environments, the findings demonstrate the constant advantage of K-Step SARSA ($\Delta$) over K-Step SARSA TD across all evaluated environments.
-This superiority is seen through elevated average rewards, accelerated convergence, and enhanced stability in learning outcomes.
-Environments characterized by deterministic state transitions, such as CliffWalking-v0 and FrozenLake-v1, demonstrate the clear superiority of K-Step SARSA ($\Delta$), but intricate stochastic environments like Blackjack-v1 and Taxi-v3 further highlight its resilience.
+In conclusion, the comparison of results across all four environments reveals a consistent advantage of K-Step SARSA ($\Delta$) over K-Step SARSA TD in each case.
+This advantage is demonstrated by increased average rewards, quicker convergence, and enhanced stability in learning results.
+Environments defined by deterministic state transitions, such as CliffWalking-v0 and FrozenLake-v1, clearly showcase the advantages of K-Step SARSA ($\Delta$).
+In contrast, more complex stochastic environments like Blackjack-v1 and Taxi-v3 further emphasize its robustness.
 The performance graph of deterministic environments, such as CliffWalking-v0 and FrozenLake-v1, resembles an exponential growth plateau curve due to predictable state transitions.
-In contrast, the performance graph of complex stochastic environments, like Blackjack-v1 and Taxi-v3, appears as a flat line because of the uncertainty and variability in state transitions.
-7 Conclusion
-With the help of TD($\Delta$), SARSA($\Delta$) improves SARSA by integrating multiple discount factors, leading to better stability and performance.
-With this approach, we can get a more precise breakdown of action-values while still enjoying the benefits of on-policy learning.
-In addition to the deterministic and stochastic OpenAI Gym environments, we also conducted trials in a tabular synthetic environment, specifically the five-star ring MDP.
-In all investigated environments, our experimental results demonstrate that SARSA($\Delta$), our suggested approach, outperforms k-step SARSA TD in terms of greater average rewards, faster convergence, and more stable learning trajectories.
-The results of this work advance our understanding of multi-step RL algorithms and how they are applied to various sequential decision-making tasks.
-This information can aid in the creation of robust algorithms for unpredictable, real-world environments in the future.
-In essence, the Q-Learning version of TD($\Delta$) extends this technique.
+On the contrary, the performance graph for sophisticated stochastic environments, such as Blackjack-v1 and Taxi-v3, resembles a flat line due to the uncertainty and unpredictability in state transitions.
+Table 2 presents the average rewards along with confidence intervals (CI) for each algorithm within their respective environments.
+Kruskal statistical tests [45] were conducted at a significance level of $\alpha=0.05$ to compare K-step SARSA TD and K-Step SARSA ($\Delta$) baseline models.
+Table 2: This table presents a comparative analysis of K-Step SARSA TD and K-Step SARSA ($\Delta$) across multiple OpenAI Gym Toy Text environments.
+Algorithm Environments
+Blackjack-v1 CliffWalking-v0 FrozenLake-v1 Taxi-v3
+[…]
+[…]
+6.3 OpenAI Gym Atari Environments
+To assess the efficacy of the proposed PPO-TD($\lambda$, SARSA($\Delta$)) method, experiments were conducted in two Atari environments: SeaquestNoFrameskip-v4 and QbertNoFrameskip-v4 [3], and the results were compared with those obtained from the standard PPO algorithm [39].
+The results shown in Figs.
+7a and  7b show that in both environments, the proposed technique achieves better stability and better long-term rewards.
+The experiments tracked average rewards during training episodes ranging from 10,000 to 100,000.
+The experiments employ the code cited in Kostrikov[44].
+(a) SeaquestNoFrameskip-v4
+(b) QbertNoFrameskip-v4
+Figure 6: The OpenAI Gym Atari control framework is utilized for conducting long-horizon experiments. Arranged from left to right: (a) SeaquestNoFrameskip-v4 and (b) QbertNoFrameskip-v4.
+In SeaquestNoFrameskip-v4, PPO-TD($\lambda$, SARSA($\Delta$)) demonstrates superior performance compared to PPO-Standard at four out of five evaluation points, as illustrated in Fig.
+7a.
+Following 100K episodes, the mean reward for PPO-TD($\lambda$, SARSA($\Delta$)) is around 14.35, while PPO-Standard attains 14.25, reflecting a 0.7% enhancement.
+PPO-Standard has a pronounced peak after 60K episodes, achieving an average reward of around 15.25, after which it declines and plateaus.
+In contrast, PPO-TD($\lambda$, SARSA($\Delta$)) exhibits more refined learning dynamics and sustains stable performance following its peak at 40K episodes.
+The performance difference in QbertNoFrameskip-v4 is more substantial, as illustrated in Fig.
+7b.
+At 100K episodes, PPO-TD($\lambda$, SARSA($\Delta$)) attains an average reward of approximately 275, whereas PPO-Standard achieves about 210, leading to an improvement of approximately 31%.
+It is noteworthy that PPO-TD($\lambda$, SARSA($\Delta$)) starts with lower rewards, around 155 at 10K episodes, in contrast to 210 for PPO-Standard.
+However, it exhibits a steady progress, finally surpassing the baseline after 60K episodes, and reaching a high of around 290 episodes after 80K episodes.
+The wider shaded region for the proposed technique in the earlier phases indicates more variance, which diminishes as training progresses—signifying improved convergence and learning efficiency with time.
+In conclusion, these results highlight the advantages of integrating PPO and SARSA($\Delta$) with TD($\lambda$, $\Delta$).
+The suggested method enhances the stability of action-value estimates across various time scales and fosters superior long-term performance in dense-reward environments, especially in intricate games such as Qbert.
+(a) SeaquestNoFrameskip-v4
+(b) QbertNoFrameskip-v4
+Figure 7: Comparison of performance between PPO-Standard and PPO-TD($\lambda$, SARSA($\Delta$)) across two Atari games: (a) SeaquestNoFrameskip-v4 and (b) QbertNoFrameskip-v4.
+The x-axis denotes the quantity of training episodes (measured in thousands), whereas the y-axis illustrates the average rewards obtained.
+The shaded areas represent the standard deviation calculated from 10 runs with different random seeds.
+Table 3 shows the average rewards for each algorithm in their own environments, along with confidence intervals (CI).
+We used Kruskal statistical tests [45] at a significant level ($\alpha=0.05$) to compare the PPO-TD($\lambda$, SARSA($\Delta$))/PPO baseline models.
+Table 3: Comparative analysis of PPO and PPO-TD($\lambda$, SARSA($\Delta$)) across various OpenAI Gym Atari environments.
+Algorithm Atari Environments
+Seaquest Qbert
+[…]
+[…]
+7 Discussion and Conclusions
+By combining SARSA with TD($\Delta$) and various discount factors, SARSA($\Delta$) improves SARSA’s stability and performance.
+This method maintains the benefits of on-policy learning while enabling a more precise breakdown of action-values.
+To evaluate the effectiveness of SARSA($\Delta$), experiments were performed in both deterministic and stochastic environments from OpenAI Gym, as well as in a synthetic tabular environment—specifically, the five-star ring MDP.
+Furthermore, enhancements in performance within dense-reward environments of the Atari suite were demonstrated through a PPO-based variant of SARSA($\Delta$).
+The experimental results across all examined environments indicate that the proposed method, SARSA($\Delta$), surpasses k-step SARSA TD by achieving higher average rewards, quicker convergence, and more consistent learning trajectories.
+The PPO-based variant of SARSA ($\Delta$) also demonstrates superior performance compared to standard PPO algorithms in dense-reward Atari environments, as evidenced by the experimental results outlined in Section 6.3.
+This research contributes to our knowledge of multi-step RL algorithms and its applicability to diverse sequential decision-making tasks.
+Limitations: The proposed SARSA($\Delta$) method has four limitations that require attention: (i) Memory Overhead: When using a large number of $\Delta$ components, the memory needs increase since the component-level updates scale linearly with the number of discount factors.
+This technique requires an extra O(M) in storage compared to conventional SARSA.
+(ii) Stochastic Environments: Best results may not be attained through rigid scheduling of $\eta_{m}$ in scenarios characterized by high unpredictability of rewards.
+In certain cases, adaptive scheduling of $\eta_{m}$ can boost both the stability and the effectiveness of learning.
+(iii) Partial Observability Markov Decision Process (POMDP): Current formulations are inapplicable to environments characterized by partial observability, as they presuppose full observability of the Markov Decision Process (MDP).
+(iv) Extension to off-policy: There are numerous issues associated with adapting SARSA($\Delta$) to off-policy algorithms such as Q-learning.
+Off-policy learning has the potential to significantly increase variance in long-horizon components, which may subsequently lead to instability during the training process.
+Future Directions: (i) Incorporation with Off-Policy Q-Learning: The adaptability of the SARSA($\Delta$) idea might be enhanced by applying it to off-policy algorithms, similar to the Q-learning version of TD($\Delta$).
+(ii) Applications to Multi-Agent Systems: In a multi-agent system, the agents coordinate on aligned subgoals by sharing $D_{m}$ components.
+(iii) Adaptive Scheduling via Meta-Gradients: Performance in various environments may be improved by using meta-gradient approaches for dynamically adjusting $\eta_{m}$.
+(iv) Extension to Partially Observable: Adapting SARSA($\Delta$) for environments characterized by partial observability is an effective approach to enhance its relevance in real-world applications.
 We decided to take care of it later.
 Acknowledgments
+This research is funded by the Innovation Teams of Ordinary Universities in Guangdong Province under Grants (2021KCXTD038, 2023KCXTD022), the Key Laboratory of Ordinary Universities in Guangdong Province (2022KSYS003), Key Discipline Research Ability Improvement Project of Guangdong Province (2021ZDJS043, 2022ZDJS068), the Natural Science Foundation of Guangdong Province (2022A1515010990), the Research Fund of the Department of Education of Guangdong Province Grants (2022KTSCX079, 2023ZDZX1013, 2022ZDZX3012, 2022ZDZX3011, 2023ZDZX2038), Chaozhou Engineering Technology Research Center (z25025) and Hanshan Normal University project (XY202105).
+We express our gratitude to the editors and referees for their invaluable suggestions and remarks.
 Author contributions statement
+M.H.
+participated in all stages, from Conceptualization design to writing—original draft preparation.
+G.Z., and X.D.
+funding acquisition.
+X.C.
+and Y.Z.
+supervision and project administration.
+W.H., L.M., S.Q., Z.Z., and P.W.
+investigation.
+Z.U.
+and N.J.
+validation and formal analysis.
 Data availability
+The corresponding authors may be contacted to request data related to this paper.
 Competing interests
@@
 Appendix
-PROOFS
+Appendix A PROOFS
 Proof: Theorem 1
-The proof utilizes induction.
-Initially, it is necessary to establish that the base case is valid at n=0.
-This statement is evidently valid given our initial assumption, particularly in the context of zero initialization.
-At a specific time-step n, we assume the validity of the statement, namely, $\sum_{m=0}^{M}\theta_{n}^{m}=\theta_{n}^{\eta}$.
-We will now demonstrate that it also holds at the subsequent time step, n+1.
-$\sum_{m=0}^{M}\theta_{n+1}^{m}$ $=\sum_{m=0}^{M}\bigg{(}\theta_{n}^{m}+\alpha_{m}\bigg{(}G_{n}^{M,\lambda_{m}}-\hat{D}_{m}(s_{n},a_{n})\bigg{)}\phi(s_{n},a_{n})\bigg{)}$
 […]
-$=\theta_{n}^{\eta}+\sum_{m=0}^{M}\alpha_{m}\bigg{(}\sum_{k=n}^{\infty}(\lambda_{m}\eta_{m})^{k-n}\delta_{k}^{m}\bigg{)}\phi(s_{n},a_{n})\text{ , Assumption based on induction.}$.
-$=\theta_{n}^{\eta}+\alpha\sum_{k=n}^{\infty}(\lambda\eta)^{k-n}\underbrace{\bigg{(}\sum_{m=0}^{M}\delta_{k}^{m}\bigg{)}}_{*}\phi(s_{n},a_{n})\text{ , Because of $\alpha_{m}=\alpha$, $\lambda_{m}\eta_{m}=\lambda_{\eta}$,$\forall m$}$
-To show that $\sum_{m=0}^{M}\theta_{n+1}^{m}=\theta_{n+1}^{\eta}$, we must establish that the term (*) = $\sum_{m=0}^{M}\delta_{k}^{m}$ is equivalent to the conventional TD error $\delta_{k}^{\eta}$.
-$\sum_{m=0}^{M}\delta_{k}^{m}$ $=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})-\hat{Q}_{\eta_{0}}(s_{k},a_{k})+\sum_{m=1}^{M}\bigg{(}(\eta_{m}-\eta_{m-1})\sum_{q=0}^{m-1}\langle\theta_{n}^{q},\phi(s_{k+1},a_{k+1})\rangle+\eta_{m}\langle\theta_{n}^{m},\phi(s_{k+1},a_{k+1})\rangle-\langle\theta_{n}^{m},\phi(s_{k},a_{k})\rangle\bigg{)}$
-$=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})+\bigg{\langle}\sum_{m=1}^{M}\bigg{(}\eta_{m}\sum_{q=0}^{m}\theta_{n}^{q}-\eta_{m-1}\sum_{q=0}^{m-1}\theta_{n}^{q}\bigg{)},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-\bigg{\langle}\sum_{m=0}^{M}\theta_{n}^{m},\phi(s_{k},a_{k})\bigg{\rangle}$
-$=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})+\eta_{m}\bigg{\langle}\sum_{m=0}^{M}\theta_{n}^{m},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-\eta_{0}\bigg{\langle}\theta_{n}^{0},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-Q_{\eta}(s_{k},a_{k})$
-$=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})+\eta\hat{Q}_{\eta}(s_{k+1},a_{k+1})-\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})-\hat{Q}_{\eta}(s_{k},a_{k})$
-$=r_{k}+\eta\hat{Q}_{\eta}(s_{k+1},a_{k+1})-\hat{Q}_{\eta}(s_{k},a_{k})$
-$=\delta_{k}^{\eta}$ (32)
 Proof: Theorem 2
@@
 The separated action-value function is denoted by D in this form, and the update is given in terms of the operator $\mathrm{T}_{\lambda}$, which is responsible for applying time-scale separation.
-In the event that $10\leq\lambda\eta<$ 1, the formula is considered to be well-defined.
-This is because the spectral norm of the operator $\lambda\eta P$ is less than 1, and thus, the Eq.
+In the event that $0\leq\lambda\eta<1$, the formula is considered to be well-defined.
+This is because the spectral norm of the operator $\lambda\eta P$ is less than 1, and thus, Eq.
 $I-\lambda\eta P$ is invertible.
 On the other hand, when the value of $\lambda$ is bigger than one, the equivalence between Eqs.
-34 and 35 is lost.
-This is not a problem because, in fact, we use the TD error for training, which, according to our expectations, corresponds to the definition of $\mathrm{T}_{\lambda}$ that is provided in Eq.
-35.
+31 and 32 is lost.
+This is not a problem because, in fact, training is based on the TD error, which, according to expectations, corresponds to the definition of $\mathrm{T}_{\lambda}$ that is provided in Eq.
+32.
 Now, let’s have a look at the contraction property that the operator $\mathrm{T}_{\lambda}$ possesses.
 To begin, the Eq.
-designated as 35 can be rewritten as follows:
+designated as 32 can be rewritten as follows:
 $\mathrm{T}_{\lambda}=(I-\lambda\eta P)^{-1}(\mathrm{T}D-\lambda\eta PD)$ (33)
@@
 $=(I-\lambda\eta P)^{-1}\bigg{(}\eta(1-\lambda)P(D_{1}-D_{2})\bigg{)}$
-We derive the following:
+The following is derived:
 $\|\mathrm{T}_{\lambda}D_{1}-\mathrm{T}_{\lambda}D_{2}\|\leq\frac{\eta|1-\lambda|}{1-\lambda\eta}\|D_{1}-D_{2}\|\textit{ where P = 1}$
@@
 Proof: Theorem 3
+In SARSA, the $\mathit{Q-value}$ function, Q(s, a), is estimated using samples of state-action pairs and their corresponding rewards.
+Let $Q_{n}(s,a)$ denote the estimate of the $\mathit{Q-value}$ at time n.
+For $\mathit{n}$ samples, by Hoeffding’s inequality provides a guarantee for a variable that is bounded between [1,+1][-1,+1], that, the probability that the sample mean deviates from the true mean by more than $\epsilon$ is expressed as follows:
+$Q^{\pi}(s,a)=\mathbf{E}\bigg{[}\eta_{0}+\eta r_{1}+...+\eta^{k-1}r_{k-1}+\eta^{k}Q^{\pi}(s_{k},a_{k})\bigg{]}$
 […]
-$Q^{\pi}(s,a)=\mathbf{E}\bigg{[}\eta_{0}+\eta r_{1}+...+\eta^{k-1}r_{k-1}+\eta^{k}Q^{\pi}(s_{k},a_{k})\bigg{]}$
-Here the expectations are over a random trajectory under $\pi$; thus $\mathbf{E}[r_{\iota}](\iota\leq k\mathit{-}1)$ denotes the expected value of the $\iota$th reward received, while $\mathbf{E}[Q^{\pi}(s_{k},a_{k})]$ is the expected value of the true value function at the $k$th state-action reached.
-The phased TD(k) update sums the terms $\eta^{\iota}\bigg{(}\frac{1}{n}\bigg{)}\sum_{i=1}^{n}r_{\iota}^{i}$, whose expectations are exactly the $\eta^{\iota}\mathbf{E}[r_{\iota}]$ appearing above [9].
-[…]
+$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{x=1}^{n}r_{i}^{(x)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\geq\epsilon\bigg{)}\leq 2e^{\bigg{(}\mathit{-}\frac{2n^{{}^{2}}\epsilon^{2}}{\sum_{x=1}^{n}(b\mathit{-}a)^{2}}\bigg{)}}\textit{where a=-1 and b=+1}$
 $\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{x=1}^{n}r_{i}^{(x)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\geq\epsilon\bigg{)}\leq 2e^{\bigg{(}\mathit{-}\frac{2n^{{}^{2}}\epsilon^{2}}{n2^{2}}\bigg{)}}$ (35)
-If we assume n and the probability of exceeding an $\epsilon$ value is fixed to be atmost $\delta$, we can solve for the resulting value of $\epsilon$:
+Assuming that n and the probability of surpassing an $\epsilon$ value is fixed to be atmost $\delta$, we can derive the corresponding value of $\epsilon$:
 $2e^{\bigg{(}\mathit{-}\frac{2n^{{}^{2}}\epsilon^{2}}{n2^{2}}\bigg{)}}=\delta$
@@
 $\epsilon=\sqrt{\frac{2log\frac{2}{\delta}}{n}}$ (36)
-So by Hoeffding, if we have n samples, with probability at least $1\mathit{-}\delta$,
+So by Hoeffding’s inequality, for n samples, the following holds with probability at least $1\mathit{-}\delta$,
 $\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{x=1}^{n}r_{i}^{(x)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\bigg{)}\leq\epsilon=\sqrt{\frac{2log\frac{2}{\delta}}{n}}$ (37)
-Since we are dealing with k different state-action pairs, we use a union bound. To ensure the probability holds for all k hypotheses, adjust $\delta$ to $\delta/k$:
+Since the analysis involves k different state-action pairs, a union bound is applied. To ensure the probability holds for all k hypotheses, adjust $\delta$ to $\delta/k$:
 $\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{x=1}^{n}r_{i}^{(x)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\bigg{)}\leq\epsilon=\sqrt{\frac{2log\frac{2k}{\delta}}{n}}$
-We can now assume that all $\mathbb{E}[r_{i}]$ terms are estimated to at least $\epsilon$ accuracy. Substituting this back into the definition of the k-step TD update we get
+It is now assumed that all $\mathbb{E}[r_{i}]$ terms are estimated to at least $\epsilon$ accuracy. Substituting this back into the definition of the k-step TD update we get
 $\hat{Q}_{n+1}(s,a)\mathit{-}Q(s,a)=\frac{1}{n}\sum_{x=1}^{n}\bigg{(}r_{0}+\eta r_{1}+...+\eta^{k-1}r_{k-1}+\eta^{k}Q_{n}(s_{k},a_{k})\bigg{)}\mathit{-}Q(s,a)$
@@
 $\eta^{k}\Delta^{\hat{Q}_{\eta}}_{n-1}$
-and then the second term is bounded by $\Delta^{\hat{Q}_{\eta}}_{n-1}$ by assumption. Hence, combining these terms, we get the overall bound for SARSA:
+and then the second term is bounded by $\Delta^{\hat{Q}_{\eta}}_{n-1}$ by assumption. Therefore, the combination of these terms results in the comprehensive bound for SARSA:
 $\Delta^{\hat{Q}_{\eta}}_{n}\leq\epsilon\bigg{(}\frac{1-\eta^{k}}{1-\eta}\bigg{)}+\eta^{k}\Delta^{\hat{Q}_{\eta}}_{n-1}$ (40)
@@
 Phased $TD(\Delta)$ update rules for $m\geq 1$:
-$\hat{D}_{m,n}(s,a)=\frac{1}{n}\sum_{x-1}^{n}\bigg{(}\sum_{i=1}^{k_{m}-1}(\eta_{m}^{i}-\eta_{m-1}^{i})r_{i}^{(x)}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})\hat{Q}_{\eta_{m-1}}(s_{k}^{(x)},a_{k}^{(x)})+\eta_{m}^{k_{m}}\hat{D}_{m}(s_{k}^{(x)},a_{k}^{(x)})\bigg{)}$ (44)
-We know that according to the multi-step update rule 17 for $m\geq 1$:
+$\begin{split}\hat{D}_{m,n}(s,a)=\frac{1}{n}\sum_{x-1}^{n}\biggl{(}\sum_{i=1}^{k_{m}-1}(\eta_{m}^{i}-\eta_{m-1}^{i})r_{i}^{(x)}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})\hat{Q}_{\eta_{m-1}}(s_{k}^{(x)},a_{k}^{(x)})+\eta_{m}^{k_{m}}\hat{D}_{m}(s_{k}^{(x)},a_{k}^{(x)})\biggr{)}\end{split}$ (41)
+According to the multi-step update rule 16 for $m\geq 1$:
 $D_{m}(s_{n},a_{n})=\mathbb{E}\bigg{[}\sum_{i=1}^{k_{m}-1}\bigg{(}\eta_{m}^{i}-\eta_{m-1}^{i}\bigg{)}r_{i}+\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}Q_{\eta_{m-1}}(s_{k},a_{k})+\eta_{m}^{k_{m}}D_{m}(s_{k},a_{k})\bigg{]}$ (42)
 Then, subtracting the two expressions gives for $m\geq 1$:
-$\hat{D}_{m,n}(s,a)-D_{m}(s_{n},a_{n})=\sum_{i=1}^{k_{m}-1}\bigg{(}\eta_{m}^{i}-\eta_{m-1}^{i}\bigg{)}\bigg{(}\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}-\mathbb{E}[r_{i}]\bigg{)}+\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\bigg{(}\sum_{q=0}^{m-1}\hat{D}_{q}(s_{k}^{(x)},a_{k}^{(x)})-\mathbb{E}[D_{m}(s_{k},a_{k})]\bigg{)}$
-$+\eta_{m}^{k_{m}}\bigg{(}D_{m}(s_{k}^{(x)},a_{k}^{(x)})-\mathbb{E}[D_{m}(s_{k},a_{k})]\bigg{)}$ (46)
-Assume that $k_{0}\leq k_{1}\leq...k_{m}=k$, the D estimates share at most $k_{m}=k$ reward terms $\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}$, Using Hoeffding inequality and union bound, we obtain that with probability $1-\delta$, each k empirical average reward $\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}$ deviates from the true expected reward $\mathbb{E}[r_{i}]$ by at most $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$.
-Hence, with probability $1-\delta$, $\forall m\geq 1$, we have:
+$\begin{split}\hat{D}_{m,n}(s,a)-D_{m}(s_{n},a_{n})&=\sum_{i=1}^{k_{m}-1}\bigg{(}\eta_{m}^{i}-\eta_{m-1}^{i}\bigg{)}\bigg{(}\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}-\mathbb{E}[r_{i}]\bigg{)}+\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\\ &\bigg{(}\sum_{q=0}^{m-1}\hat{D}_{q}(s_{k}^{(x)},a_{k}^{(x)})-\mathbb{E}[D_{m}(s_{k},a_{k})]\bigg{)}+\eta_{m}^{k_{m}}\bigg{(}D_{m}(s_{k}^{(x)},a_{k}^{(x)})-\mathbb{E}[D_{m}(s_{k},a_{k})]\bigg{)}\end{split}$
+Assume that $k_{0}\leq k_{1}\leq...k_{m}=k$, the D estimates share at most $k_{m}=k$ reward terms $\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}$, Using Hoeffding inequality and union bound, it follows that, with probability $1-\delta$, each k empirical average reward $\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}$ deviates from the true expected reward $\mathbb{E}[r_{i}]$ by at most $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$.
+Hence, with probability $1-\delta$, $\forall m\geq 1$, the following holds:
 $\Delta_{n}^{m}\leq\epsilon\sum_{i=1}^{k_{m}-1}\bigg{(}\eta_{m}^{i}-\eta_{m-1}^{i}\bigg{)}+\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\sum_{q=0}^{m-1}\Delta_{n-1}^{q}+\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$
@@
 $=\sum_{m=0}^{M-1}(\eta_{m}^{k_{m}}-\eta_{m}^{k_{m+1}})\sum_{q=0}^{m}\Delta_{n-1}^{q}+\eta_{M}^{k_{m}}\sum_{m=0}^{M}\Delta_{n-1}^{m}$ (45)
-Finally, we obtain:
+Finally, the following is obtained:
 $\sum_{m=0}^{M}\Delta_{n}^{m}\leq\underbrace{\epsilon\bigg{(}\frac{1-\eta^{k}}{1-\eta}\bigg{)}+\epsilon\bigg{(}\sum_{m=0}^{M-1}\frac{\eta_{m}^{k_{m+1}}-\eta^{k_{m}}_{m}}{1-\eta_{m}}\bigg{)}}_{\text{variance reduction}}+\underbrace{\sum_{m=0}^{M-1}\bigg{(}\eta^{k_{m}}_{m}-\eta^{k_{m+1}}_{m}\bigg{)}\sum_{q=0}^{m}\Delta_{n-1}^{q}+\eta^{k}\sum_{m=0}^{M}\Delta_{n-1}^{m}}_{\text{bias introduction}}$ (46)
+Appendix B Hyperparameters
+The experiments primarily utilize the hyperparameters recommended by Romoff 771 et al.[1], Schulman et al.[39] and Kostrikov[44].
+The experiments utilize the following randomly generated seeds: 125125, 513, 90135, 81212, 3523401, 15709, 17, 0, 8412, and 1153780.
+Additional hyperparameter configurations are detailed in Table 4.
+Table 4: Hyperparameter configurations and their corresponding values.
+Hyperparameter Value
+seed (10) np.random.randint(40, 45)
+Learning Rate ($\alpha$) [0.1, 0.3, 0.5, 0.7, 1.0]
+Discount Factor ($\eta$) [0.992, 0.95, 0.996, 0.99, 1.0]
+k-Step [1, 4, 8, 16, 32, 128, 256]
+Number of Episodes 100000
+Max Epsilon 0.4
+Clipping parameter 0.1
+Number of actors 8
+Horizon(T) 129
+Number of Epochs 3
+Lambda ($\lambda$) 0.95